### v1

# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

Abstract
Large Language Models (LLMs) struggle with hallucinations and outdated knowledge due to their reliance on static training data. Retrieval-Augmented Generation (RAG) mitigates these issues by integrating external dynamic information enhancing factual and updated grounding. Recent advances in multimodal learning have led to the development of Multimodal RAG, incorporating multiple modalities such as text, images, audio, and video to enhance the generated outputs. However, cross-modal alignment and reasoning introduce unique challenges to Multimodal RAG, distinguishing it from traditional unimodal RAG. This survey offers a structured and comprehensive analysis of Multimodal RAG systems, covering datasets, metrics, benchmarks, evaluation, methodologies, and innovations in retrieval, fusion, augmentation, and generation. We precisely review training strategies, robustness enhancements, and loss functions, while also exploring the diverse Multimodal RAG scenarios. Furthermore, we discuss open challenges and future research directions to support advancements in this evolving field. This survey lays the foundation for developing more capable and reliable AI systems that effectively leverage multimodal dynamic external knowledge bases. Resources are available at https://github.com/llm-lab-org/Multimodal-RAG-Survey.
Ask in Any Modality:
A Comprehensive Survey on Multimodal Retrieval-Augmented Generation
[…]
**footnotetext: These authors contributed equally.
1 Introduction & Background
In recent years, significant breakthroughs have been achieved in language models, driven primarily by the advent of transformers Vaswani et al. (2017), enhanced computational capabilities, and the availability of large-scale training data Naveed et al. (2024). The emergence of foundational Large Language Models (LLMs) Ouyang et al. (2022); Grattafiori et al. (2024); Touvron et al. (2023); Qwen et al. (2025); Anil et al. (2023) has revolutionized natural language processing (NLP), demonstrating unprecedented capabilities in a wide range of tasks including instruction following Qin et al. (2024), sophisticated reasoning Wei et al. (2024), In-context Learning Brown et al. (2020), and multilingual machine translation Zhu et al. (2024a). These advancements have elevated the performance of various NLP tasks, opening new avenues for research and application. Despite their remarkable achievements, LLMs face significant challenges, including hallucination, outdated internal knowledge, and a lack of verifiable reasoning Huang et al. (2024a); Xu et al. (2024b). Their reliance on parametric memory restricts their ability to access up-to-date knowledge, making them less effective for knowledge-intensive tasks compared to task-specific architectures. Moreover, providing provenance for their decisions and updating their world knowledge remain critical open problems Lewis et al. (2020).
Retrieval-Augmented Generation (RAG)
RAG Lewis et al. (2020) has emerged as a promising solution to these limitations by enabling LLMs to retrieve and incorporate external knowledge, improving factual accuracy and reducing hallucinations Shuster et al. (2021); Ding et al. (2024a). By dynamically accessing vast external knowledge repositories, RAG systems enhance knowledge-intensive tasks while ensuring responses remain grounded in verifiable sources Gao et al. (2023). In practice, RAG systems operate through a retriever-generator pipeline. The retriever leverages embedding models Chen et al. (2024b); Rau et al. (2024) to identify relevant passages from external knowledge bases and optionally applies re-ranking techniques to improve retrieval precision Dong et al. (2024a). These retrieved passages are then passed to the generator, incorporating this external context to produce informed responses. Recent advancements in RAG frameworks Asai et al. (2023); An et al. (2024); Lee et al. (2024); Liu et al. (2024c) have introduced iterative inference processes, enabling more accurate and reliable answers by refining both retrieval and generation stages. However, traditional RAG architectures are primarily designed for textual information, limiting their ability to address multimodal challenges that require integrating diverse data formats.
Multimodal Learning
Parallel to these developments, significant advances in multimodal learning have reshaped artificial intelligence by enabling systems to integrate and analyze heterogeneous data sources for a holistic representation of information. The introduction of CLIP (Contrastive Language-Image Pre-training) Radford et al. (2021) marked a pivotal moment in connecting visual and textual information through contrastive learning, inspiring numerous subsequent models and applications Alayrac et al. (2024); Wang et al. (2023); Pramanick et al. (2023). These breakthroughs have driven progress in various domains, including sentiment analysis Das and Singh (2023) and cutting-edge biomedical research Hemker et al. (2024), demonstrating the value of multimodal approaches. By enabling systems to process and understand diverse data types such as text, images, audio, and video, multimodal learning has become essential to advancing artificial general intelligence (AGI) Song et al. (2025).
Multimodal RAG
The extension of LLMs to multimodal LLMs (MLLMs) has further expanded their capabilities, allowing them to process, reason, and generate outputs across diverse modalities Liu et al. (2023a); Team et al. (2024); Li et al. (2023b). For instance, GPT-4 OpenAI et al. (2024) demonstrates human-level performance in various benchmarks by accepting inputs in both text and images, marking a significant milestone in multimodal perception and interaction. Building on this foundation, multimodal RAG systems extend traditional RAG frameworks by incorporating multimodal knowledge sources, such as images and audio, to provide enriched context for generation Hu et al. (2023); Chen et al. (2022a). This integration enhances the precision of generated outputs while leveraging multimodal cues to improve the reasoning capabilities of MLLMs. The general flow of the Multimodal RAG pipeline is illustrated in Figure 1. However, these multimodal systems also present unique challenges, including determining which modalities to retrieve, effectively fusing diverse data types, and addressing the complexities of cross-modal relevance Zhao et al. (2023).
Task Formulation
We present a mathematical formulation of the general task for multimodal RAG systems. These systems generate a multimodal response, denoted as $r$, in response to a query $q$, which is typically in textual format.
Let $D=\{d_{1},d_{2},...,d_{n}\}$ be a corpus consisting of $n$ multimodal documents. Each document $d_{i}$ in $D$ is associated with a modality $M_{d_{i}}$ and is processed by a modality-specific encoder $Enc_{M_{d_{i}}}$:
$z_{i}=Enc_{M_{d_{i}}}(d_{i})$ (1)
The set of all encoded representations is denoted by:
$Z=\{z_{1},z_{2},...,z_{n}\}$ (2)
Modality-specific encoders map different modalities into a shared semantic space for cross-modal alignment. A retrieval model $R$ assesses the relevance of each encoded document representation $z$ with respect to the query $q$, represented as $R(q,z)$. To construct the retrieval-augmented multimodal context, the retrieval model selects the most relevant documents based on a modality-specific threshold:
$X=\{d_{i}\mid s(e_{q},z_{i})\geq\tau_{M_{d_{i}}}\}$ (3)
where $\tau_{M_{d_{i}}}$ is a relevancy threshold for the modality of $M_{d_{i}}$, $e_{q}$ is the encoded representation of $q$ in the shared semantic space, and $s$ is a scoring function that measures the relevance between the encoded query and document representations. The generative model $G$ produces the final multimodal response by conditioning on both the user query $q$ and the retrieved documents $X$ as context:
$r=G(q,X)$ (4)
Related Works
As the field of multimodal RAGs is newly introduced and evolving rapidly, especially in recent years, there is a pressing need for a comprehensive survey that explores the current innovations and frontiers of these systems. While more than ten surveys have been published on RAG-related topics such as Agentic RAG Singh et al. (2025), none provide a detailed and comprehensive overview of advancements in multimodal RAGs. The only related survey to date Zhao et al. (2023) categorizes multimodal RAGs by grouping relevant papers based on their applications and modalities. However, our survey provides a more detailed and different innovation-driven perspective, offering a detailed taxonomy and exploring emerging trends and challenges in depth. Furthermore, significant advancements have been made in the field since its publication, and interest in this topic has grown substantially within the research community. In this survey, we review over 100 papers on multimodal RAGs published in recent years, primarily from the ACL Anthology and other repositories such as the ACM Digital Library.
Contributions
In this work, (i) we provide a comprehensive review of the multimodal RAG field, covering task formulation, datasets, benchmarks, tasks and domain-specific applications, evaluation, and key innovations across retrieval, fusion, augmentation, generation, training strategies, and loss functions. (ii) We introduce a precise structured taxonomy (Figure 2) that categorizes state-of-the-art models based on their primary contributions, highlighting methodological advancements and emerging frontiers. (iii) To support further research, we make resources, including datasets, benchmarks, and key innovations, publicly available. (iv) We identify current research trends and knowledge gaps, providing insights and recommendations to guide future advancements in this evolving field.
Figure 1: Overview of the multimodal retrieval-augmented generation (RAG) pipeline, highlighting the advancements and techniques employed at each stage. The flow begins with query preprocessing, where user queries are refined and then encoded into a shared embedding space alongside a multimodal database. Retrieval strategies, such as modality-centric retrieval, similarity search, and re-ranking, enhance document selection, while fusion mechanisms align and integrate data from multiple modalities using score fusion or attention-based methods. Augmentation techniques such as iterative retrieval with feedback mechanisms, further refine the retrieved documents for the multimodal LLM. The generation stage incorporates innovations like Chain-of-Thought reasoning and source attribution for better outputs, with loss functions combining alignment loss and generation loss to optimize both retrieval and generation components. Noise management techniques are also applied to improve training stability and robustness.
2 Datasets and Benchmarks
Multimodal RAG research employs diverse datasets and benchmarks to evaluate retrieval, integration, and generation across heterogeneous sources. Image–text tasks, including captioning and retrieval, commonly use MS-COCO Lin et al. (2014), Flickr30K Young et al. (2014), and LAION-400M Schuhmann et al. (2021), while visual question answering with external knowledge is supported by OK-VQA Marino et al. (2019) and WebQA Chang et al. (2022). For complex multimodal reasoning, MultimodalQA Talmor et al. (2021) integrates text, images, and tables, whereas video-text tasks leverage ActivityNet Caba Heilbron et al. (2015) and YouCook2 Zhou et al. (2018). In the medical domain, MIMIC-CXR Johnson et al. (2019) and CheXpert Irvin et al. (2019) facilitate tasks such as medical report generation. It is noteworthy that a number of these datasets are unimodal (e.g., solely text-based or image-based). Unimodal datasets are frequently employed to represent a specific modality and are subsequently integrated with complementary datasets from other modalities. This modular approach allows each dataset to contribute its domain-specific strengths, thereby enhancing the overall performance of the multimodal retrieval and generation processes.
Benchmarks assess multimodal RAG systems on visual reasoning, external knowledge integration, and dynamic retrieval. The $M^{2}RAG$ Ma et al. (2024c) benchmark provides a unified evaluation framework that combines fine-grained text-modal and multimodal metrics to jointly assess both the quality of generated language and the effective integration of visual elements Vision-focused evaluations, including MRAG-Bench Hu et al. (2024c), VQAv2 Goyal et al. (2017a) and VisDoMBench Suri et al. (2024), test models on complex visual tasks. Dyn-VQA Li et al. (2024b), MMBench Liu et al. (2025), and ScienceQA Saikh et al. (2022) evaluate dynamic retrieval and multi-hop reasoning across textual, visual, and diagrammatic inputs. knowledge-intensive benchmarks, such as TriviaQA Joshi et al. (2017) and Natural Questions Kwiatkowski et al. (2019), together with document-oriented evaluations such as OmniDocBench Ouyang et al. (2024), measure integration of unstructured and structured data. Advanced retrieval benchmarks such as RAG-Check Mortaheb et al. (2025a) evaluate retrieval relevance and system reliability, while specialized assessments like Counterfactual VQA Niu et al. (2021) test robustness against adversarial inputs. Additionally, OCR impact studies such as OHRBench Zhang et al. (2024d) examine the cascading effects of errors on RAG systems. Additional details about datasets, benchmarks, and their categorization are presented in Table 1 and Table 2 in Appendix B.
3 Evaluation
Assessing multimodal RAG models presents significant challenges due to their diverse modalities, intricate architectures, and the diverse evaluation dimensions involved. Given the broad spectrum of modalities and system components, evaluation requires a combination of metrics drawn from vision-language models (VLMs), generative AI, and retrieval systems. These metrics enable the assessment of various aspects of multimodal RAG, such as text and image generation or information retrieval, either independently or in combination. By analyzing prior works, we identified approximately 60 distinct evaluation metrics used in multimodal RAG studies. More details, including the formulas for the RAG evaluation metrics, can be found in Appendix C. In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
Retrieval Evaluation
Evaluates how well a system retrieves relevant information from a database or knowledge source. Accuracy, recall, and precision are used to evaluate retrieval and downstream task performance, the F1 score also being used to assess both Recall and Precision simultaneously. A more commonly used metric than recall is Recall@K, which measures the proportion of relevant items appearing within the top K retrieved results. Another important metric Mean Reciprocal Rank (MRR), which is utilized by Adjali et al. (2024); Nguyen et al. (2024).
Modality Evaluation
These evaluations focus on criteria that rely on modalities. Text and image are the most commonly used modalities in multimodal RAG. This category of evaluations considers various aspects of these modalities, including the alignment between generated text and images, the fluency of the text, the quality of image captions, and other related factors. In text similarity and overlap metrics, Exact Match (EM), BLEU Papineni et al. (2002), ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005) are used. MultiRAGen Shohan et al. (2024) uses Multilingual ROUGE for multilingual settings. For image captioning evaluation, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al. (2015) is a widely used metric for evaluating image captioning quality. It quantifies the alignment between a generated caption and a set of human-written reference captions by leveraging the term frequency-inverse document frequency (TF-IDF) and cosine similarity. Yasunaga et al. (2023); Zhao et al. (2024); Luo et al. (2024a); Yuan et al. (2024); Sharifymoghaddam et al. (2024); Hu et al. (2023); Rao et al. (2024); Xu et al. (2024a); Kim et al. (2024); Zhang et al. (2024c) use CIDEr for evaluation. SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al. (2016) is a metric that emphasizes semantic aspects, in contrast to CIDEr. SPIDEr Liu et al. (2017), utilized in Zhang et al. (2024c), is a linear combination of both SPICE and CIDEr. In semantic alignment evaluations, BERTScore Zhang et al. (2020), employed in Sun et al. (2024b); Shohan et al. (2024), measures semantic similarity by comparing BERT-generated embeddings. Chen et al. (2022a); Zhi Lim et al. (2024); Ma et al. (2024c) evaluate Fluency by employing BERTScore. CLIP Score Hessel et al. (2021), used in Sharifymoghaddam et al. (2024); Zhang et al. (2024c), measures image-text similarity using the CLIP model Radford et al. (2021). FID (Fréchet Inception Distance) Heusel et al. (2017) quantifies image quality by computing the distributional difference between the generated and original image features. Yasunaga et al. (2023); Zhao et al. (2024); Sharifymoghaddam et al. (2024); Zhang et al. (2024c) use FID in their works. Zhao et al. (2024) use Kernel Inception Distance (KID) Bińkowski et al. (2018) for evaluation. KID is another metric used to evaluate the quality of generated images. It is similar to FID but offers an unbiased estimator. Zhi Lim et al. (2024) employs the Inception Score (IS) to assess the diversity and quality of generated images based on classification probabilities obtained from the Inception network. For the evaluation of audio synthesis, Zhang et al. (2024c) assesses overall sound quality (OVL) and relevance to the text (REL) with the assistance of human annotators. Additionally, this study employs the Fréchet Audio Distance (FAD) Kilgour et al. (2019) as an evaluation metric. FAD serves a similar purpose to the FID but is specifically designed for the audio domain.
FLOPs, execution time, response time, and average retrieval time per query are utilized to assess efficiency and computational performance Nguyen et al. (2024); Strand et al. (2024); Dang (2024); Zhou (2024). Zhong et al. (2024) employs Spearman’s Rank Correlation (SRC), which calculates the Pearson correlation between the rank values of two variables. Several evaluation metrics are domain- and problem-specific. Geodesic distance measurement quantifies the model’s accuracy by calculating the geodesic distance between predicted and actual geographical coordinates Zhou et al. (2024e). Lahiri and Hu (2024) utilizes Clinical Relevance (CR) for evaluation.
4 Key Innovations and Methodologies
4.1 Retrieval Strategy
Efficient Search and Similarity Retrieval
Modern multimodal RAG systems first encode diverse modalities (text, images, tables) into a unified embedding space to enable direct cross-modal retrieval. The foundational CLIP-Radford et al. (2021) or BLIP-based Li et al. (2022a) approaches leverage contrastive learning, which aligns text and images into a shared latent space via dual-tower encoders. Recent advancements extend this paradigm with novel multimodal retrieval architectures and training strategies Zhou et al. (2024c); Wei et al. (2025); Zhang et al. (2024j). Multimodal RAG systems rely on efficient search strategies to retrieve relevant external knowledge from large-scale multimodal memory. Since dual-encoder models map multimodal inputs into a shared high-dimensional embedding space, maximum inner product search (MIPS) is commonly used for fast and direct similarity comparisons. Systems such as MuRAG Chen et al. (2022a) and RA-CM3 Yasunaga et al. (2023) employ approximate MIPS to efficiently retrieve the top-k candidates by maximizing the inner product between the query vector and a large collection of image–text embeddings. Large-scale implementations leverage distributed MIPS techniques, such as TPU-KNN Chern et al. (2022), to enable high-speed retrieval across extensive datasets. Additionally, ScaNN (Scalable Nearest Neighbors) Guo et al. (2020), MAXSIM score Chan and Ng (2008); Cho et al. (2024), and approximate KNN methods Caffagni et al. (2024) have been adopted for efficient image-text similarity computation.
Recent advancements in MIPS optimization focus on reducing retrieval latency and improving accuracy. These include adaptive quantization Li et al. (2024a); Huang et al. (2024b), hybrid sparse-dense representations Nguyen et al. (2024); Ma et al. (2024a); Zhang et al. (2024a), and learned index structures Zhai et al. (2023); Basnet et al. (2024). These techniques help optimize retrieval performance by balancing efficiency and precision in multimodal RAG systems.
Modality-Based Retrieval
Modality-aware retrieval techniques optimize retrieval efficiency by leveraging the unique characteristics of each modality. Text-centric retrieval remains foundational in multimodal RAG systems. Traditional methods like BM25 Robertson and Zaragoza (2009), along with dense retrievers such as BGE-1.5 Xiao et al. (2023) and BGE-M3 Chen et al. (2024a) dominate text-based evidence retrieval Chen et al. (2022b); Suri et al. (2024); Nan et al. (2024). Novel approaches also address the need for fine-grained semantic matching and domain specificity: for instance, ColBERT Khattab and Zaharia (2020) and PreFLMR Lin et al. (2024b) employ token-level interaction mechanisms that preserve nuanced textual details to improve precision for multimodal queries, while RAFT Zhang et al. (2024i) and CRAG Yan et al. (2024) enhance retrieval performance by ensuring precise citation of text spans, thereby optimizing domain-specific question answering within multimodal RAG pipelines. Vision-centric retrieval focuses on directly leveraging image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al. (2023). Systems such as EchoSight Yan and Xie (2024) and ImgRet Shohan et al. (2024) retrieve visually similar content by using reference images as queries. Composed Image Retrieval (CMI) models Feng et al. (2023); Zhao et al. (2024); Jang et al. (2024) enhance retrieval by integrating multiple image features into a unified query representation. Similarly, Pic2word Saito et al. (2023) maps visual content to textual descriptions, enabling zero-shot image retrieval. Video-centric retrieval extends vision-based techniques by incorporating temporal dynamics and large video-language models (LVLMs), driven by novel frameworks like iRAG Arefeen et al. (2024), which introduces incremental retrieval for sequential video understanding, and MV-Adapter Jin et al. (2024), optimizing multimodal transfer learning for video-text retrieval. Recent breakthroughs focus on long-context processing: Video-RAG Luo et al. (2024b) leverages visually aligned auxiliary texts (OCR/ASR) to enhance retrieval without proprietary models, while VideoRAG Ren et al. (2025) employs dual-channel architectures with graph-based knowledge grounding for extreme-length videos. For temporal reasoning, CTCH Shen et al. (2024) uses contrastive transformer hashing to model long-term dependencies, and RTime Du et al. (2024) introduces reversed-video hard negatives to rigorously benchmark temporal causality. Meanwhile, for complex video understanding, OmAgent Zhang et al. (2024e) adopts a divide-and-conquer framework, while DRVideo Ma et al. (2024d) addresses long-video understanding by a document-based retrieval approach.
Document Retrieval and Layout Understanding
Recent research has advanced beyond traditional uni-modal retrieval methods toward models that directly process entire document pages. These approaches integrate textual, visual, and structural elements like tables, images, font styles, and page layouts to enhance retrieval performance in complex documents. ColPali Faysse et al. (2024) pioneers end-to-end document image retrieval by embedding page patches with a vision-language backbone, bypassing OCR entirely. Models like ColQwen2 Wang et al. (2024b); Khattab and Zaharia (2020) and M3DocVQA Cho et al. (2024) extend this paradigm with dynamic resolution handling and holistic multi-page reasoning. Newer frameworks refine efficiency and layout understanding: ViTLP Mao et al. (2024) and DocLLM Wang et al. (2024a) pre-train generative models to align spatial layouts with text, while CREAM Zhang et al. (2024b) employs coarse-to-fine retrieval with multimodal efficient tuning to balance accuracy and computation costs. Finally, mPLUG-DocOwl 1.5 Hu et al. (2024a) and 2 Hu et al. (2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies. Together, these approaches highlight a shift toward holistic, layout-sensitive retrieval that leverages visual semantics and generative pre-training.
Re-ranking and Selection Strategies
Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also refining and prioritizing retrieved candidates. Re-ranking and selection strategies enhance retrieval quality by optimizing example selection, refining relevance scoring, and applying filtering mechanisms. Optimized example selection techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection strategies Luo et al. (2024a). Su et al. (2024a) refine multimodal inputs using probabilistic control keywords to improve credibility, while RULE Xia et al. (2024b) calibrates retrieved context selection through statistical methods like the Bonferroni correction to mitigate factuality risks. Furthermore, selective fusion mechanisms dynamically refine the retrieval outcomes Yuan et al. (2023), and key-frame extraction techniques ensure diversity in video-based retrieval through clustering-based frame selection Dong et al. (2024b). Several methods employ advanced scoring mechanisms to improve retrieval relevance Mortaheb et al. (2025b, a). Multimodal similarity measures, including negative log-likelihood (NLL), structural similarity index measure (SSIM) Wang et al. (2020), normalized cross-correlation (NCC), and BERTScore Zhang et al. (2020), aid in re-ranking retrieved documents based on textual and visual similarity. Cross-encoders trained on sequence classification tasks refine relevance scoring within retrieval pipelines Zhi Lim et al. (2024), while hierarchical post-processing techniques integrate retrieval, passage-level, and answer confidence scores for improved ranking Zhang et al. (2024h). LDRE Yang et al. (2024) employs semantic ensemble methods to adaptively weigh multiple caption features, whereas RAGTrans Cheng et al. (2024) and OMG-QA Nan et al. (2024) integrate traditional ranking functions like BM25 Robertson and Zaragoza (2009). Advanced re-ranking techniques also incorporate Q-Former-based image-to-text correspondence for fine-grained retrieval Yan and Xie (2024); Xu et al. (2024a). Advanced filtering methods are crucial for enhancing multi-modal RAG performance by ensuring that only high-quality, relevant data is processed. Hard negative mining techniques, as seen in GME Zhang et al. (2024j) and MM-Embed Lin et al. (2024a), use modality-aware sampling and synthesized negatives from imbalanced data to mitigate modality bias and refine contrastive learning. Similarly, consensus-based filtering approaches, exemplified by MuRAR Zhu et al. (2024d) and ColPali Faysse et al. (2024), employ source attribution and multi-vector mapping to filter out low-similarity candidates and reduce hallucinations in enterprise QA. Moreover, dynamic modality filtering methods—including RAFT Zhang et al. (2024i), Img2Loc Zhou et al. (2024e), and MAIN-RAG Chang et al. (2024)—train retrievers to disregard confusing data, thereby enhancing the discriminative capacity and overall robustness of multi-modal retrieval systems.
4.2 Fusion Mechanisms
Score Fusion and Alignment
Recent multimodal RAG models emphasize aligning embeddings or fusing scores across different modalities to enable adaptive retrieval. Zhi Lim et al. (2024) transform text, tables, and images into a unified textual representation using a cross-encoder trained for relevance scoring. To address multimodal LLM input constraints, Sharifymoghaddam et al. (2024) introduce interleaved image-text pairs, merging multiple few-shot images vertically for compatibility with models like LLaVA Liu et al. (2023a) while aligning modalities via CLIP score fusion Hessel et al. (2021) and BLIP feature fusionLi et al. (2022a). For fine-grained alignment, REVEAL Hu et al. (2023) injects retrieval scores into attention layers and minimizes L2-Norm distances between query and knowledge embeddings, and MA-LMM He et al. (2024) aligns video-text embeddings via a BLIP-inspired Query Transformer (Q-Former)Li et al. (2022a). In contrast, LLM-RA Jian et al. (2024) concatenates text and visual embeddings into joint queries to reduce retrieval noise. RA-BLIP Ding et al. (2024b) employs a 3-layer BERT-based adaptive fusion module to unify visual-textual semantics. CLIP-based alignment strategies are widely adopted; Riedler and Langer (2024) embed images and questions into a shared CLIP space, while VISA Ma et al. (2024b) uses the Document Screenshot Embedding (DSE) model as the retrieval component of the RAG system to align and fuse textual queries with visual document representations. DSE directly encodes both textual queries and document screenshots into a shared embedding space, allowing efficient retrieval based on cosine similarity. Xue et al. (2024) map object-predicate pairs into a shared semantic space using a prototype-based embedding network (PENET) Zheng et al. (2023), aligning visual features with textual prototypes. Re-IMAGEN Chen et al. (2022b) refines text-to-image synthesis with interleaved classifier-free guidance during diffusion sampling, balancing creativity and entity fidelity. To improve multimodal alignment, MegaPairs Zhou et al. (2024a) integrates CLIP-based and MLLM-based retrieval through score fusion. Meanwhile, Wiki-LLaVA Caffagni et al. (2024) employs CLIP encoders for hierarchical document retrieval. VISRAG Yu et al. (2024) refines alignment via position-weighted mean pooling on VLM hidden states, prioritizing later tokens for relevance. RAG-Driver Yuan et al. (2024) aligns visual-language embeddings via visual instruction tuning and an MLP projector.
Attention-Based Mechanisms
Attention-based fusion dynamically weights cross-modal interactions for task-specific reasoning. RAMM Yuan et al. (2023) employs a dual-stream co-attention transformer, combining self-attention (intra-modal) and cross-attention (inter-modal) to fuse retrieved biomedical images/texts with input data. Similarly, EMERGE Zhu et al. (2024b) integrates Electronic Health Records (EHR) data with a cross-attention adaptive network, while MORE Cui et al. (2024) uses cross-attention to weigh multimodal results for commonsense reasoning. For user-centric tasks, RAGTrans Cheng et al. (2024) integrates user-aware attention to fuse social media features, and AlzheimerRAG Lahiri and Hu (2024) applies cross-modal attention to synthesize text, images, and tables. Video-text alignment is advanced by MV-Adapter Jin et al. (2024), which leverages a Cross Modality Tying (CMT) to align embeddings, and M2-RAAP Dong et al. (2024b), enhances video-text fusion using an Auxiliary-Caption-Guided (ACG) strategy. It employs Frame-Caption Re-weighting (FCR) to boost frames with high inter-modal consistency and Text-Caption Re-weighting (TCR) to adjust weights based on intra-modal similarity. A Mutual-guided (Mug) Alignment Head then filters misaligned features via a dot-product similarity and frame-to-token attention, producing refined frame-specific text representations. Generative refinement is demonstrated in Xu et al. (2024a), where gated cross-attention conditions text generation on visual features. Similarly, C3Net Zhang et al. (2024c) enhances alignment between queries and candidates within CLIP’s contrastive space. Mu-RAG Chen et al. (2022a) uses intermediate fusion with cross-attention for open QA, balancing modality-specific and joint representations. Kim et al. (2024) leverage cross-modal memory retrieval with pre-trained CLIP ViT-L/14 to map video-text pairs into a shared space, enabling dense captioning through the attention-based fusion of retrieved memories.
Unified Frameworks and Projections
Architectures unifying multimodal inputs through frameworks or projections address scalability. Su et al. (2024a) employ hierarchical cross-chains and late fusion for healthcare data, while IRAMIG Liu et al. (2024b) iteratively consolidates multimodal results into knowledge representations. For document understanding, M3DocRAG Cho et al. (2024) flattens multi-page documents into a unified embedding tensor, and PDF-MVQA Ding et al. (2024c) combines Region-of-Interest (RoI)-based (VisualBERT) and patch-based (CLIP) VLPMs Long et al. (2022). DQU-CIR Wen et al. (2024) exemplifies raw-data unification by converting images into text captions for complex queries and overlaying text onto images for simple queries. The model then fuses embeddings using MLP-learned weights. SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input into unimodal text representation for subsequent processing. UFineBench Zuo et al. (2024) uses a shared granularity decoder for ultra-fine text-person retrieval. Nguyen et al. (2024) introduce Dense2Sparse (D2S) projection, converting dense embeddings from models like BLIP/ALBEF Li et al. (2022a) into sparse lexical vectors via layer normalization and probabilistic expansion control, optimizing storage and interpretability.
4.3 Augmentation Techniques
Basic RAG systems typically follow a single retrieval step and pass retrieved content directly to the generation phase without further refinement, which can lead to inefficiencies and suboptimal outputs. To address this, augmentation techniques have been introduced to enhance the utility of retrieved data before generation, ensuring that the model can better interpret, structure, and integrate multimodal information Gao et al. (2023).
Context Enrichment
This focuses on enhancing the relevance of retrieved knowledge for the generation phase by refining external knowledge or retrieved data. General approaches such as EMERGE Zhu et al. (2024b) integrate entity relationships and semantic descriptions to provide richer context. MiRAG Adjali et al. (2024) improves retrieval effectiveness by expanding initial queries through entity retrieval and query reformulation, enhancing subsequent stages for the visual question-answering task. Wiki-LLaVA Caffagni et al. (2024) incorporates text chunks and image tokens into multimodal inputs to enrich prompts for generation. Video-RAG Luo et al. (2024b) enhances long-video understanding through Query Decoupling, which reformulates user queries into structured retrieval requests to extract additional auxiliary multimodal context. Img2Loc Zhou et al. (2024e) enhances accuracy by including both the most similar and the most dissimilar points from the database in the prompt, allowing the model to rule out implausible locations for its predictions. Finally, Xue et al. (2024) introduce semantic-enhanced prompts, aligning structured image data with queries for contextual accuracy.
Adaptive and Iterative Retrieval
For more complex queries, dynamic retrieval mechanisms have proven effective. SKURG Yang et al. (2023) employs adaptive retrieval steps to dynamically adjust the number of hops based on the query complexity. IRAMIG Liu et al. (2024b) iteratively refines multimodal queries by incorporating feedback from previous iterations and context created by the fusion of retrieved items from text, images, and audio modalities. Similarly, OMG-QA Nan et al. (2024) introduces a multi-round retrieval strategy, where each retrieval step incorporates episodic memory to refine subsequent queries. An evaluation module assesses retrieval effectiveness at each step, guiding the refinement of subsequent retrieval efforts through feedback. Taking a different approach, SAM-RAG Zhai (2024) further improves relevance by dynamically selecting the optimal number of retrieved contexts for each query. MMed-RAG Xia et al. (2024a) uses context filtering to discard low-relevance results. OmniSearch Li et al. (2024b) introduces a self-adaptive retrieval agent that dynamically decomposes complex multimodal questions into sub-question chains and plans retrieval actions in real-time based on retrieved content. mR2AG Zhang et al. (2024g) introduces reflection mechanisms to enhance retrieval effectiveness. Retrieval-Reflection dynamically determines when external knowledge is necessary, while Relevance-Reflection filters retrieved content, retaining only task-critical evidence. Similarly, RAGAR Khaliq et al. (2024) iteratively refines retrieved data by leveraging prior responses and multimodal analysis, ensuring contextual consistency.
4.4 Generation Techniques
Recent advancements in multimodal RAG generation focus on robustness, cross-modal coherence, and task-specific adaptability. These innovations can be broadly categorized into these sections:
In-Context Learning
In-context learning (ICL) with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining. Models such as RMR Tan et al. (2024), Sharifymoghaddam et al. (2024), and RA-CM3 Yasunaga et al. (2023), extend this paradigm to multimodal RAG settings. RAG-Driver Yuan et al. (2024) refines ICL by retrieving relevant driving experiences from a memory database, ensuring scenario-specific contextual alignment. MSIER Luo et al. (2024a) further enhances example selection through a Multimodal Supervised In-Context Examples Retrieval framework, leveraging a foundation MLLM scorer to evaluate both textual and visual relevance. Meanwhile, Raven Rao et al. (2024) introduces Fusion-in-Context Learning, a novel approach that enriches ICL by incorporating a more diverse set of in-context examples, leading to better performance than standard ICL.
Reasoning
Structured reasoning techniques, such as chain-of-thought (CoT), decompose complex reasoning into smaller sequential steps, enhancing coherence and robustness in multimodal RAG systems. RAGAR Khaliq et al. (2024) introduces Chain of RAG and Tree of RAG to iteratively refine fact-checking queries and explore branching reasoning paths for more robust evidence generation. VisDoM Suri et al. (2024) integrates CoT with evidence curation to ensure logical and contextual accuracy, while SAM-RAG Zhai (2024) employs reasoning chains alongside multi-stage answer verification to enhance the relevance, utility, and support of generated responses. Meanwhile, LDRE Yang et al. (2024) leverages LLMs for divergent compositional reasoning, generating refined captions by incorporating dense captions and modification text.
Instruction Tuning
Several works have fine-tuned generation components for specific applications. RA-BLIP Ding et al. (2024b), uses Q-Former architecture from InstructBLIP Dai et al. (2023) to extract visual features based on question instructions. RAGPT Lang et al. (2025) uses a context-aware prompter, which captures contextual knowledge from relevant instances and generates dynamic prompts. mR2AG Zhang et al. (2024g) uses instruction tuning with the mR2AG-IT dataset to train MLLMs to adaptively invoke retrieval, identify relevant evidence, and generate accurate answers for knowledge-based VQA tasks. RagVL Chen et al. (2024d) employs instruction tuning to enhance the ranking capability of MLLMs, serving them as a re-ranker for filtering the top-k retrieved images. Jang et al. (2024) propose a Visual Delta Generator to discriminate differences between image pairs and produce a textual response. MMed-RAG Xia et al. (2024a) uses preference fine-tuning to help the model learn when to rely on retrieved information and when to trust its internal knowledge. Additionally, several methods have been proposed to improve generation quality by constructing datasets from the previous mistakes of LLMs and fine-tuning them accordingly. MegaPairs Zhou et al. (2024a) synthesizes large-scale multimodal instruction-tuning datasets for this purpose. Similarly, Surf Sun et al. (2024a) creates a dataset of positive and negative pairs derived from the prior errors of LVLMs and subsequently fine-tunes the LVLM using this dataset. Additionally, Rule Xia et al. (2024b) refines Med-LVLM by applying direct preference optimization on curated preference data, specifically focusing on instances where overreliance on retrieved contexts resulted in errors.
Source Attribution and Evidence Transparency
Ensuring source attribution in multimodal RAG systems is a key focus of recent research. MuRAR Zhu et al. (2024d) integrates multimodal data, fetched by a source-based retriever, to refine LLM’s initial response, ensuring coherence and informativeness. VISA Ma et al. (2024b) uses large vision-language models (VLMs) to generate answers with visual source attribution by identifying and highlighting supporting evidence in retrieved document screenshots. Similarly, OMG-QA Nan et al. (2024) enhances transparency by prompting the LLM to explicitly cite evidence in generated responses.
4.5 Training Strategies
Training multimodal RAG models involves a multi-stage process to effectively handle cross-modal interactions Chen et al. (2022a). Pretraining establishes the foundation using large paired datasets to learn cross-modal relationships while fine-tuning adapts models to specific tasks by leveraging cross-modal attention Ye et al. (2019). For instance, REVEAL Hu et al. (2023) integrates multiple training objectives. Its pretraining phase optimizes Prefix Language Modeling Loss ($L_{\text{PrefixLM}}$), where the model predicts text continuations from a prefix and an image. Supporting losses include Contrastive Loss ($L_{\text{contra}}$) for aligning queries with pseudo-ground-truth knowledge, Disentangled Regularization Loss ($L_{\text{decor}}$) to improve embedding expressiveness, and Alignment Regularization Loss ($L_{\text{align}}$) to align query and knowledge embeddings. During fine-tuning, a cross-entropy objective trains the model for tasks like VQA and image captioning.
Alignment
Contrastive learning improves representation quality by pulling positive pairs closer and pushing negative pairs apart in the embedding space. A common objective is the InfoNCE loss van den Oord et al. (2019), which maximizes the mutual information between positive pairs while minimizing similarity to negatives. Several multimodal RAG models, such as VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a) and SAM-RAG Zhai (2024) utilize InfoNCE loss to improve retrieval-augmented generation. Furthermore, EchoSight Yan and Xie (2024) enhances retrieval accuracy by selecting visually similar yet contextually distinct negatives, while HACL Jiang et al. (2024) improves generation by introducing hallucinative captions as distractors. Similarly, UniRaG Zhi Lim et al. (2024) strengthens retrieval by incorporating hard negative documents, helping the model distinguish relevant contexts from noise.
The eCLIP loss Kumar and Marttinen (2024) extends contrastive training by integrating expert-annotated data and an auxiliary Mean Squared Error (MSE) loss to refine embedding quality. Mixup strategies further augment training by generating synthetic positive pairs, improving generalization in contrastive learning Kumar and Marttinen (2024). Dense2Sparse Nguyen et al. (2024) incorporates two unidirectional losses: an image-to-caption loss $\ell(I\to C)$ and a caption-to-image loss $\ell(C\to I)$. It enforces sparsity through L1 regularization, optimizing retrieval precision by balancing dense and sparse representations. Details of formulas for widely used RAG loss functions can be found in Appendix D.
Robustness
Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024). Managing noisy retrieval inputs is critical for maintaining model performance. MORE Cui et al. (2024) injects irrelevant results during training to enhance focus on relevant inputs. AlzheimerRAG Lahiri and Hu (2024) uses progressive knowledge distillation to reduce noise while maintaining multimodal alignment. RAGTrans Cheng et al. (2024) leverages hypergraph-based knowledge aggregation to refine multimodal representations, ensuring more effective propagation of relevant information. RA-BLIP Ding et al. (2024b) introduces the Adaptive Selection Knowledge Generation (ASKG) strategy, which leverages the implicit capabilities of LLMs to filter relevant knowledge for generation through a denoising-enhanced loss term, eliminating the need for fine-tuning. This approach achieves strong performance compared to baselines while significantly reducing computational overhead by minimizing trainable parameters. RagVL Chen et al. (2024d) improves robustness through noise-injected training by adding hard negative samples at the data level and applying Gaussian noise with loss reweighting at the token level, enhancing the model’s resilience to multimodal noise. Finally, RA-CM3 Yasunaga et al. (2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2em, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0.4em, inner ysep=0.6em, line width=0.8pt, text width=8.5em, ver/.style= fill=gray!10, rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=8em , leaf/.style= text opacity=1, inner sep=2pt, fill opacity=.5, fill=hidden-blue!90, text=black, text width=44.5em font=, inner xsep=0.4em, inner ysep=0.6em, draw, , , [Multimodal RAG, ver [ Retrieval
Strategies (§4.1) [ Efficient-Search and
Similarity Retrieval
 (§4.1) [ Maximum Inner Product Search (MIPS) [ TPU-KNN Chern et al. (2022), ScaNN Guo et al. (2020), MAXSIM score Chan and Ng (2008), OmniSearch Li et al. (2024b), MuRAG Chen et al. (2022a), RetrievalAttention Liu et al. (2024a), RA-CM3 Yasunaga et al. (2023), RAVEN Rao et al. (2024) FactMM-RAG Sun et al. (2024b) , leaf, text width=44.5em ] ] [ Multimodal
Encoders [ CLIP Radford et al. (2021), BLIP Li et al. (2022a), MARVEL Zhou et al. (2024c), ALIGN Jia et al. (2021), FLAVA Singh et al. (2022), UniVL-DR Liu et al. (2023b), UniIR Wei et al. (2025), GME Zhang et al. (2024j), VISTA Zhou et al. (2024b), ColPali Faysse et al. (2024), InternVideo Wang et al. (2022), Ovis Lu et al. (2024) Mi-RAG Adjali et al. (2024) , leaf, text width=44.5em ] ] ] [ Modality-Centric
Retrieval (§4.1) [ Text-Centric [ Contriever Izacard et al. (2022), GTE Li et al. (2023c), Re-Imagen Chen et al. (2022b), BM25 Robertson and Zaragoza (2009), BGE-1.5 Xiao et al. (2023), BGE-M3 Chen et al. (2024a), CapRet Shohan et al. (2024), OMG-QA Nan et al. (2024), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al. (2024b), RAFT Zhang et al. (2024i), CRAG Yan et al. (2024), M2-RAG Ma et al. (2024c) , leaf, text width=44.5em ] ] [ Vision-Centric [ VQA4CIR Feng et al. (2023), Unifashion Zhao et al. (2024), Jang et al. Jang et al. (2024), Pic2word Saito et al. (2023), eClip Kumar and Marttinen (2024), RAMM Yuan et al. (2023), Joshi et al. Joshi et al. (2024), VISA Ma et al. (2024b), ImgRet Shohan et al. (2024), EchoSight Yan and Xie (2024), Xue et al. Xue et al. (2024) , leaf, text width=44.5em ] ] [ Video-Centric [ iRAG Arefeen et al. (2024), VideoRAG Ren et al. (2025), VideoRAG Jeong et al. (2025), MV-Adapter Jin et al. (2024), OmAgent Zhang et al. (2024f), CM2 Kim et al. (2024), Video-RAG Luo et al. (2024b), CTCH Shen et al. (2024), RTime Du et al. (2024), VideoMAE Tong et al. (2022), DrVideo Ma et al. (2024d) , leaf, text width=44.5em ] ] [ Document Retrieval [ ColPali Faysse et al. (2024), ColQwen2 Wang et al. (2024b), M3DocVQA Cho et al. (2024), ViTLP Mao et al. (2024), DocLLM Wang et al. (2024a), CREAM Zhang et al. (2024b), mPLUG-DocOwl 1.5 Hu et al. (2024a), mPLUG-DocOwl 2 Hu et al. (2024b), VisDom Suri et al. (2024), DSE Ma et al. (2024a) , leaf, text width=44.5em ] ] ] [ Re-ranking
Strategies (§4.1) [ Optimized Example
Selection [ MSIER Luo et al. (2024a), Hybrid RAG Su et al. (2024b), RULE Xia et al. (2024b), RAMM Yuan et al. (2023), M2RAAP Dong et al. (2024b) , leaf, text width=44.5em ] ] [ Relevance Score
Evaluation [ RAG-Check Mortaheb et al. (2025a, b), UniRaG Zhi Lim et al. (2024), MR2AG Zhang et al. (2024h), LDRE Yang et al. (2024), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al. (2024), OMG-QA Nan et al. (2024), EchoSight Yan and Xie (2024), EgoInstructor Xu et al. (2024a) , leaf, text width=44.5em ] ] [ Filtering
Mechanisms [ MAIN-RAG Chang et al. (2024), MM-Embed Lin et al. (2024a), GME Zhang et al. (2024j), Img2Loc Zhou (2024), MuRAR Zhu et al. (2024d) RAFT Zhang et al. (2024i) , leaf, text width=44.5em ] ] ] ] [ Fusion
Mechanisms (§4.2) [ Score Fusion
and Alignment
 (§4.2) [ M3 Cai et al. (2025)  Zhi Lim et al. (2024),  Sharifymoghaddam et al. (2024), REVEAL Hu et al. (2023), RAG-Driver Yuan et al. (2024), LLM-RA Jian et al. (2024),  Riedler and Langer (2024), VISA Ma et al. (2024b), MA-LMM He et al. (2024),  Xue et al. (2024), RA-BLIP Ding et al. (2024b), Re-IMAGEN Chen et al. (2022b), MegaPairs Zhou et al. (2024a), Wiki-LLaVA Caffagni et al. (2024), VISRAG Yu et al. (2024) , leaf, text width=55.0em ] ] [ Attention-Based
Mechanisms (§4.2) [ RAMM Yuan et al. (2023), EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), RAGTrans Cheng et al. (2024), AlzheimerRAG Lahiri and Hu (2024), MV-Adapter Jin et al. (2024), C3Net Zhang et al. (2024c),  Xu et al. (2024a),  Kim et al. (2024), M2-RAAP Dong et al. (2024b), Mu-RAG Chen et al. (2022a), Ou et al. Ou et al. (2025), CADMR Khalafaoui et al. (2024) , leaf, text width=55.0em ] ] [ Unified Frameworks
and Projections
 (§4.2) [ Hybrid-RAG Su et al. (2024b), Dense2Sparse Nguyen et al. (2024), IRAMIG Liu et al. (2024b), M3DocRAG Cho et al. (2024), DQU-CIR Wen et al. (2024), PDF-MVQA Ding et al. (2024c), SAM-RAG Zhai (2024), UFineBench Zuo et al. (2024),  Li et al. (2022a), , leaf, text width=55.0em ] ] ] [ Augmentation
Techniques (§4.3) [ Context-Enrichment (§4.3) [ EMERGE Zhu et al. (2024b), MiRAG Adjali et al. (2024), Wiki-LLaVA Caffagni et al. (2024), Video-RAG Luo et al. (2024b), Img2Loc Zhou et al. (2024e),  Xue et al. (2024) , leaf, text width=55.0em ] ] [ Adaptive and
Iterative
Retrieval (§4.3) [ SKURG Yang et al. (2023), IRAMIG Liu et al. (2024b), OMG-QA Nan et al. (2024), SAM-RAG Zhai (2024), MMed-RAG Xia et al. (2024a), OmniSearch Li et al. (2024b), mR2AG Zhang et al. (2024g), RAGAR Khaliq et al. (2024) , leaf, text width=55.0em ] ] ] [ Generation
Techniques (§4.4) [ In-Context
Learning (§4.4) [ RMR Tan et al. (2024),  Sharifymoghaddam et al. (2024), RA-CM3 Yasunaga et al. (2023), RAG-Driver Yuan et al. (2024), MSIER Luo et al. (2024a), Raven Rao et al. (2024) , leaf, text width=55.0em ] ] [ Reasoning (§4.4) [ RAGAR Khaliq et al. (2024), VisDoM Suri et al. (2024), SAM-RAG Zhai (2024), LDRE Yang et al. (2024) , leaf, text width=55.0em ] ] [ Instruction
Tuning (§4.4) [ RA-BLIP Ding et al. (2024b), RAGPT Lang et al. (2025), mR2AG Zhang et al. (2024g), RagVL Chen et al. (2024d), Jang et al. (2024), MMed-RAG Xia et al. (2024a), MegaPairs Zhou et al. (2024a), Surf Sun et al. (2024a), Rule Xia et al. (2024b) , leaf, text width=55.0em ] ] [ Source
Attribution (§4.4) [ MuRAR Zhu et al. (2024d), VISA Ma et al. (2024b), OMG-QA Nan et al. (2024) , leaf, text width=55.0em ] ] ] [ Training Strategies (§4.5) [ Alignment (§4.5) [ VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a), SAM-RAG Zhai (2024), EchoSight Yan and Xie (2024), HACL Jiang et al. (2024),  Zhi Lim et al. (2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al. (2024) , leaf, text width=55.0em ] ] [ Robustness (§4.5) [  Buettner and Kovashka (2024), MORE Cui et al. (2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al. (2024), RA-BLIP Ding et al. (2024b), RagVL Chen et al. (2024d), RA-CM3 Yasunaga et al. (2023) , leaf, text width=55.0em ] ] ] ] ]
Figure 2: Taxonomy of recent advances and enhancements in multimodal retrieval-augmented generation research.
5 Tasks Addressed by Multimodal RAGs
Multimodal RAG systems extend RAG beyond unimodal settings to tasks requiring cross-modal integration, enhancing performance across modalities such as text, images, and audio. In content generation, these models enhance image captioning Zhi Lim et al. (2024); Hu et al. (2023); Rao et al. (2024) and text-to-image synthesis Yasunaga et al. (2023); Chen et al. (2022b) by retrieving relevant contextual information. They also improve coherence in visual storytelling and ensure factual alignment in multimodal summarization Tonmoy et al. (2024). In knowledge-intensive applications, multimodal RAG supports open-domain and knowledge-seeking question answering Chen et al. (2024d); Ding et al. (2024b); Yuan et al. (2023), video-based QA Luo et al. (2024b), and fact verification Khaliq et al. (2024), grounding responses in retrieved knowledge and thereby mitigating hallucinations. Cross-modal retrieval further advances zero-shot image–text retrieval Yang et al. (2024); Dong et al. (2024b). Additionally, the recent incorporation of chain-of-thought reasoning Zhai (2024); Khaliq et al. (2024) has enhanced its ability to support complex problem solving and logical inference. Finally, integrating multimodal RAG into interactive agents and AI assistants such as Gemini Team et al. (2024) enables natural language-driven visual search, document understanding, and multimodal reasoning. The taxonomy of application domains can be seen in Figure 3. The following sections explore domain-specific adaptations of these techniques in greater depth.
Healthcare and Medicine
Multimodal RAG enhances clinical decision-making through integrated analysis of medical imaging, electronic health records, and biomedical literature. Systems like MMED-RAG Xia et al. (2024a) address diagnostic uncertainty in medical visual question answering by aligning radiology images with contextual patient data. RULE Xia et al. (2024b) mitigates hallucinations in automated report generation through dynamic retrieval of clinically similar cases. AsthmaBot Bahaj and Ghogho (2024) introduces a multimodal RAG-based approach for supporting asthma patients across multiple languages, enabling structured, language-specific semantic searches. Predictive frameworks such as Realm Zhu et al. (2024c) demonstrate robust risk assessment by fusing heterogeneous patient data streams, while Hybrid RAG Su et al. (2024a) advances privacy-preserving architectures for federated clinical data integration. FactMM-RAG Sun et al. (2024b) automates radiology report drafting by retrieving biomarker correlations from medical ontologies, exemplifying the paradigm’s capacity to operationalize expert knowledge at scale.
Software Engineering
Code generation systems leverage multimodal RAG to synthesize context-aware solutions from technical documentation and version histories. DocPrompting Zhou et al. (2023) improves semantic coherence in code completion by retrieving API specifications and debugging patterns. Commit message generation models like RACE Shi et al. (2022) contextualize code diffs against historical repository activity, while CEDAR Nashid et al. (2023) optimizes few-shot learning through retrieval-based prompt engineering. REDCODER Parvez et al. (2021) enhances code summarization via semantic search across open-source repositories, preserving syntactic conventions across programming paradigms.
Fashion and E-Commerce
Cross-modal alignment drives advancements in product discovery and design automation. UniFashion Zhao et al. (2024) enables style-aware retrieval by jointly embedding garment images and textual descriptors, while Dang (2024) reduces search friction through multimodal query expansion. LLM4DESIGN Chen et al. (2024c) demonstrates architectural design automation by retrieving compliance constraints and environmental impact assessments, underscoring RAG’s adaptability to creative domains.
Entertainment and Social Computing
Multimedia analytics benefit from RAG’s capacity to correlate heterogeneous signals. SoccerRAG Strand et al. (2024) derives tactical insights by linking match footage with player statistics. MMRA Zhong et al. (2024) predicts content virality through joint modeling of visual aesthetics and linguistic engagement patterns.
Emerging Applications
Autonomous systems adopt multimodal RAG for explainable decision-making, as seen in RAG-Driver’s Yuan et al. (2024) real-time retrieval of traffic scenarios during navigation. ENWAR Nazar et al. (2024) enhances wireless network resilience through multi-sensor fusion, while Riedler and Langer (2024) streamline equipment maintenance by retrieving schematics during fault diagnosis. Geospatial systems such as Img2Loc Zhou et al. (2024e) advance image geolocalization through cross-modal landmark correlation.
6 Open Problems and Future Directions
Despite rapid advancements in multimodal RAG systems, fundamental challenges remain in achieving robust, efficient, and human-like reasoning across modalities.
Generalization, Explainability, and Robustness
Multimodal RAG systems often struggle with domain-specific training and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al. (2020). Explainability remains a key challenge, as these systems typically fail to attribute answers to precise sources. Current methods often cite entire documents or large visual regions as source attribution rather than identifying the exact part of an image, speech, or other modality that led to the answer Ma et al. (2024b); Hu et al. (2023). In addition, the interplay between modalities affects the quality of outcomes produced by these models; for example, answers derived solely from text sources may differ in quality compared to those requiring a combination of text and image inputs Baltrusaitis et al. (2019). They are also vulnerable to adversarial perturbations, such as misleading images influencing textual outputs, and their performance can degrade when relying on low-quality or outdated sources Chen et al. (2022b). While the trustworthiness of unimodal RAGs has been studied Zhou et al. (2024d), enhancing the robustness of multimodal RAGs remains an open challenge and a promising research direction.
Reasoning, Alignment, and Retrieval Enhancement
Multimodal RAGs struggle with compositional reasoning, where information from different modalities must be logically integrated to generate coherent and contextually rich outputs. While cross-modal techniques such as Multimodal-CoT have been proposed Zhang et al. (2023), further innovations are needed to improve the coherence and contextual relevance of multimodal outputs. Enhancing modality alignment and retrieval strategies, particularly for entity-aware retrieval, is essential. Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain largely underexplored in multimodal RAGs compared to text-based RAGs Zhang et al. (2024g); Procko and Ochoa (2024). Retrieval biases such as position sensitivity Hu et al. (2024c), redundant retrieval Nan et al. (2024), and biases propagated from training data or retrieved content Zhai (2024), pose significant challenges that require further attention. Another promising direction is developing a unified embedding space for all modalities, enabling direct multimodal search without intermediary conversion models (e.g., ASRs). Despite some progress, mapping multimodal knowledge into a unified space remains an open challenge with significant potential.
Agent-Based and Self-Guided Systems
Recent trends indicate a shift towards agent-based multimodal RAGs that integrate retrieval, reasoning, and generation across diverse domains. Unlike static RAG systems, future multimodal RAGs should integrate interactive feedback and self-guided decision-making to iteratively refine outputs. Furthermore, existing feedback mechanisms often fail to accurately identify whether issues with responses arise during retrieval, generation, or other stages Dong et al. (2024b). The incorporation of reinforcement learning and end-to-end human-aligned feedback into multimodal RAGs remains largely unexplored but holds significant potential for enhancing these systems. These methods could enable multimodal RAGs to assess whether retrieval is necessary, evaluate the relevance of retrieved content, and dynamically determine the most suitable modalities for response generation. Achieving robust support for any-to-any modality is essential for adaptability in open-ended tasks Wu et al. (2024b). Future multimodal RAGs should incorporate data from diverse real-world sources, such as environmental sensors, alongside traditional modalities like text and images, to enhance situational awareness. This progression aligns with the trend toward embodied AI, where models integrate knowledge with physical interaction, enabling applications in robotics, navigation, and physics-informed reasoning. Bridging retrieval-based reasoning with real-world agency brings these systems closer to AGI.
Long-Context Processing, Scalability, and Personalization
High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing. Fixed extraction rates often fail to capture relevant video frames, necessitating dynamic adjustment based on content complexity and movement Kandhare and Gisselbrecht (2024). Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers highlight the need for more scalable architectures. Personalization mechanisms, such as adapting retrieval to user-specific contexts like medical history, remain underexplored. As personalization mechanisms evolve, ensuring privacy and mitigating the risks of sensitive data leakage in multimodal outputs remain critical challenges. Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples hinders robust evaluation.
7 Conclusion
In this study, we take a deep dive into the realm of multimodal Retrieval-Augmented Generation (RAG) literature. Specifically, we explore and categorize key advancements across different aspects of multimodal RAG systems, including retrieval, multimodal fusion, augmentation, generation, and training strategies. Additionally, we examine the tasks these systems address, their domain-specific applications, and the datasets, benchmarks, and evaluation methods. We also discuss open challenges and limitations in current approaches. We hope it could encourage future research in the domain, particularly in improving cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
8 Limitations
This study offers a comprehensive examination of multimodal RAG systems. Extended discussions, details of datasets and benchmarks, and additional relevant work are available in the Appendices. While we have made our maximum effort; however, some limits may persist. First, due to space constraints, our descriptions of individual methodologies are necessarily concise. Second, although we curate studies from major venues (e.g., ACL, EMNLP, NeurIPS, CVPR, ICLR, ICML, ACM Multimedia) and arXiv, our selection may inadvertently overlook emerging or domain-specific research, with a primary focus on recent advancements. Additionally, this work does not include a comparative performance evaluation of the various models, as task definitions, evaluation metrics, and implementation details vary significantly across studies, and executing these models requires substantial computational resources.
Furthermore, multimodal RAG is a rapidly evolving field with many open questions, such as optimizing fusion strategies for diverse modalities and addressing scalability challenges. As new paradigms emerge, our taxonomy and conclusions will inevitably evolve. To address these gaps, we plan to continuously monitor developments and update this survey and the corresponding GitHub repository to incorporate overlooked contributions and refine our perspectives.
Appendix A Taxonomy
In this section, we present a comprehensive taxonomy of multimodal retrieval-augmented generation (RAG) systems, along with a classification of their application domains. Figures 2 and 3 summarize these taxonomies, respectively.
Figure 2 provides an overview of recent advances in multimodal retrieval-augmented generation (RAG) systems. The taxonomy is organized into several key categories.
  • Retrieval strategies cover efficient search and similarity retrieval methods (including maximum inner product search (MIPS) variants and different multi-modal encoders) and modality-centric techniques that distinguish between text-, vision-, and video-centric as well as document retrieval models. Re-ranking strategies further refine these methods via optimized example selection, relevance scoring, and filtering.
  • Fusion mechanisms are implemented through score fusion and alignment, attention-based techniques, and unified frameworks that project multimodal information into common representations.
  • Augmentation techniques address context enrichment as well as adaptive and iterative retrieval.
  • Generation methods include in-context learning, reasoning, instruction tuning, and source attribution.
  • training strategies are characterized by approaches to alignment and robustness.
Detailed discussions of these categories are provided in the corresponding sections.
Figure 3
presents the taxonomy of application domains for multimodal RAG systems. The identified domains include healthcare and medicine, software engineering, fashion and e-commerce, entertainment and social computing, and emerging applications. This classification offers a concise overview of the diverse applications and serves as a framework for the more detailed analyses that follow.
Appendix B Dataset and Benchmark
Table 1 and Table 2 present a comprehensive overview of datasets and benchmarks commonly employed in multimodal RAG research. The table is organized into five columns:
  • Category: This column categorizes each dataset or benchmark based on its primary domain or modality. The datasets are grouped into eight categories: Image–Text General, Video–Text, Audio–Text, Medical, Fashion, 3D, Knowledge & QA, and Other. The benchmarks are grouped into three categories: Vision-Focused, Cross-Modal and Reasoning-Focused, and Text and Knowledge-Focused. This classification facilitates a clearer understanding of each dataset or benchmark’s role within a multimodal framework.
  • Name: The official name of the dataset or benchmarks is provided along with a citation for reference.
  • Statistics and Description: This column summarizes key details such as dataset size, the nature of the content (e.g., image–text pairs, video captions, QA pairs), and the specific tasks or applications for which the dataset or benchmarks are used. These descriptions are intended to convey the dataset’s scope and its relevance to various multimodal RAG tasks.
  • Modalities: The modalities covered by each dataset or benchmark are indicated (e.g., Image, Text, Video, Audio, or 3D). Notably, several datasets are unimodal; however, within multimodal RAG systems, these are combined with others to represent distinct aspects of a broader multimodal context.
  • Link: A hyperlink is provided to direct readers to the official repository or additional resources for the dataset or benchmark, thereby facilitating further exploration of its properties and applications.
Table 1: Overview of Popular Datasets in Multimodal RAG Research.
Category
Name Statistics and Description Modalities Link
Image-Text General LAION-400M Schuhmann et al. (2021) 200M image–text pairs; used for pre-training multimodal models. Image, Text LAION-400M
Conceptual-Captions (CC) Sharma et al. (2018) 15M image–caption pairs; multilingual English–German image descriptions. Image, Text Conceptual Captions
CIRR Liu et al. (2021) 36,554 triplets from 21,552 images; focuses on natural image relationships. Image, Text CIRR
MS-COCO Lin et al. (2014) 330K images with captions; used for caption–to–image and image–to–caption generation. Image, Text MS-COCO
Flickr30K Young et al. (2014) 31K images annotated with five English captions per image. Image, Text Flickr30K
Multi30K Elliott et al. (2016) 30k German captions from native speakers and human–translated captions. Image, Text Multi30K
NoCaps Agrawal et al. (2019) For zero–shot image captioning evaluation; 15K images. Image, Text NoCaps
Laion-5B Schuhmann et al. (2022) 5B image–text pairs used as external memory for retrieval. Image, Text LAION-5B
COCO-CN Author and Author (2018) 20,341 images for cross-lingual tagging and captioning with Chinese sentences. Image, Text COCO-CN
CIRCO Baldrati et al. (2023) 1,020 queries with an average of 4.53 ground truths per query; for composed image retrieval. Image, Text CIRCO
Video-Text BDD-X Xu et al. (2018) 77 hours of driving videos with expert textual explanations; for explainable driving behavior. Video, Text BDD-X
YouCook2 Zhou et al. (2018) 2,000 cooking videos with aligned descriptions; focused on video–text tasks. Video, Text YouCook2
ActivityNet Caba Heilbron et al. (2015) 20,000 videos with multiple captions; used for video understanding and captioning. Video, Text ActivityNet
SoccerNet Giancola et al. (2018) Videos and metadata for 550 soccer games; includes transcribed commentary and key event annotations. Video, Text SoccerNet
MSR-VTT Xu et al. (2016) 10,000 videos with 20 captions each; a large video description dataset. Video, Text MSR-VTT
MSVD Chen and Dolan (2011) 1,970 videos with approximately 40 captions per video. Video, Text MSVD
LSMDC Rohrbach et al. (2015) 118,081 video–text pairs from 202 movies; a movie description dataset. Video, Text LSMDC
DiDemo Anne Hendricks et al. (2017) 10,000 videos with four concatenated captions per video; with temporal localization of events. Video, Text DiDemo
Breakfast Kuehne et al. (2014) 1,712 videos of breakfast preparation; one of the largest fully annotated video datasets. Video, Text Breakfast
COIN Tang et al. (2019) 11,827 instructional YouTube videos across 180 tasks; for comprehensive instructional video analysis. Video, Text COIN
MSRVTT-QA Xu et al. (2017) Video question answering benchmark. Video, Text MSRVTT-QA
MSVD-QA Xu et al. (2017) 1,970 video clips with approximately 50.5K QA pairs; video QA dataset. Video, Text MSVD-QA
ActivityNet-QA Yu et al. (2019) 58,000 human–annotated QA pairs on 5,800 videos; benchmark for video QA models. Video, Text ActivityNet-QA
EpicKitchens-100 Dima (2020) 700 videos (100 hours of cooking activities) for online action prediction; egocentric vision dataset. Video, Text EPIC-KITCHENS-100
Ego4D Grauman et al. (2022) 4.3M video–text pairs for egocentric videos; massive–scale egocentric video dataset. Video, Text Ego4D
HowTo100M Miech et al. (2019) 136M video clips with captions from 1.2M YouTube videos; for learning text–video embeddings. Video, Text HowTo100M
CharadesEgo Sigurdsson et al. (2018) 68,536 activity instances from ego–exo videos; used for evaluation. Video, Text Charades-Ego
ActivityNet Captions Krishna et al. (2017) 20K videos with 3.7 temporally localized sentences per video; dense–captioning events in videos. Video, Text ActivityNet Captions
VATEX Wang et al. (2019) 34,991 videos, each with multiple captions; a multilingual video–and–language dataset. Video, Text VATEX
Charades Sigurdsson et al. (2016) 9,848 video clips with textual descriptions; a multimodal research dataset. Video, Text Charades
WebVid Bain et al. (2021) 10M video–text pairs (refined to WebVid-Refined-1M). Video, Text WebVid
Youku-mPLUG Xu et al. (2023) Chinese dataset with 10M video–text pairs (refined to Youku-Refined-1M). Video, Text Youku-mPLUG
Audio-Text LibriSpeech Panayotov et al. (2015) 1,000 hours of read English speech with corresponding text; ASR corpus based on audiobooks. Audio, Text LibriSpeech
AudioCap Kim et al. (2019) 46K audio clips paired with human-written text captions. Audio, Text AudioCaps
AudioSet Gemmeke et al. (2017) 2,084,320 human–labeled 10–second sound clips from YouTube; 632 audio event classes. Audio, Text AudioSet
Medical MIMIC-CXR Johnson et al. (2019) 125,417 training pairs of chest X–rays and reports. Image, Text MIMIC-CXR
CheXpert Irvin et al. (2019) 224,316 chest radiographs of 65,240 patients; focused on medical analysis. Image, Text CheXpert
MIMIC-III Johnson et al. (2016) Health-related data from over 40K patients (text data). Text MIMIC-III
IU-Xray Pavlopoulos et al. (2019) 7,470 pairs of chest X–rays and corresponding diagnostic reports. Image, Text IU X-ray
PubLayNet Zhong et al. (2019) 100,000 training samples and 2,160 test samples built from PubLayNet (tailored for the medical domain). Image, Text PubLayNet
Fashion Fashion-IQ Wu et al. (2019) 77,684 images across three categories; evaluated with Recall@10 and Recall@50. Image, Text Fashion IQ
FashionGen Hadi Kiapour et al. (2018) 260.5K image–text pairs of fashion images and item descriptions. Image, Text Fashion-Gen
VITON-HD Choi et al. (2021) 83K images for virtual try–on; high–resolution clothing items. Image, Text VITON-HD
Fashionpedia Author and Author (2023a) 48,000 fashion images annotated with segmentation masks and fine-grained attributes. Image, Text Fashionpedia
DeepFashion Liu et al. (2016) Approximately 800K diverse fashion images for pseudo triplet generation. Image, Text DeepFashion
3D ShapeNet Chang et al. (2015) 7,500 text–3D data pairs; repository for 3D CAD models. Text, 3D ShapeNet
Knowledge & QA VQA Antol et al. (2015) 400K QA pairs with images for visual question answering. Image, Text VQA
PAQ Lewis et al. (2021) 65M text–based QA pairs; a large–scale dataset. Text PAQ
ELI5 Fan et al. (2019) 270K complex and diverse questions augmented with web pages and images. Text ELI5
ViQuAE Biten et al. (2022) 11.8M passages from Wikipedia covering 2,397 unique entities; knowledge–intensive QA. Text ViQuAE
OK-VQA Marino et al. (2019) 14K questions requiring external knowledge for VQA. Image, Text OK-VQA
WebQA Li et al. (2022b) 46K queries that require reasoning across text and images. Text, Image WebQA
Infoseek Li et al. (2021) Fine-grained visual knowledge retrieval using a Wikipedia–based knowledge base ( 6M passages). Image, Text Infoseek
ClueWeb22 Callan et al. (2022) 10 billion web pages organized into three subsets; a large–scale web corpus. Text ClueWeb22
MOCHEG Yao et al. (2023) 15,601 claims annotated with truthfulness labels and accompanied by textual and image evidence. Text, Image MOCHEG
VQA v2 Goyal et al. (2017b) 1.1M questions (augmented with VG–QA questions) for fine-tuning VQA models. Image, Text VQA v2
ScienceQA Lu et al. (2022) High–school science curriculum dataset with 12K training samples and over 21K questions across various subjects. Text ScienceQA
A-OKVQA Schwenk et al. (2022) Benchmark for visual question answering using world knowledge; around 25K questions. Image, Text A-OKVQA
MMBench Liu et al. (2025) Bilingual benchmark for assessing the multimodal capabilities of VLMs; over 3,000 multiple–choice questions covering 20 dimensions. Text MMBench
XL-HeadTags Shohan et al. (2024) 415K news headline-article pairs consist of 20 languages across six diverse language families. Text XL-HeadTags
SEED-Bench Li et al. (2023a) 19K multiple–choice questions with accurate human annotations across 12 evaluation dimensions. Text SEED-Bench
Other ImageNet Deng et al. (2009) 14,197,122 images for perspective understanding; a hierarchical image database. Image ImageNet
Oxford Flowers102 Nilsback and Zisserman (2008) 102 flower categories with five examples per category; image classification dataset. Image Oxford Flowers102
Stanford Cars Krause et al. (2013) Images of different car models (five examples per model); for fine-grained categorization. Image Stanford Cars
GeoDE Author and Author (2023b) 61,940 images from 40 classes across 6 world regions; emphasizes geographic diversity in object recognition. Image GeoDE
Table 2: Overview of Popular Benchmarks in Multimodal RAG Research.
Category Name Statistics and Description Modalities Link
Vision-Focused MRAG-Bench Hu et al. (2024c)
Evaluates visual information retrieval and integration;
tests robustness to irrelevant visual information and visual reasoning.
Images MRAG-Bench
VQAv2 Goyal et al. (2017a)
Contains 265,016 images with 1,105,904 questions; assesses the ability to answer questions about images,
integrating image and question text, visual reasoning, and common-sense reasoning.
Images + Text VQAv2
OK-VQA Marino et al. (2019)
Comprises 14,031 questions; tests retrieval and integration of external information,
combining image, question, and external knowledge, focusing on reasoning with external knowledge.
Images + Text OK-VQA
Cross-Modal and Reasoning-Focused M2RAG Ma et al. (2024c)
Establishes a comprehensive benchmark for multimodal retrieval-augmented generation;
integrates fine-grained text-modal and multimodal metrics to evaluate foundation models on dynamic retrieval,
multi-hop reasoning, and multimodal integration.
Images + Text M2RAG
Dyn-VQA Li et al. (2024b)
Emphasizes dynamic retrieval with rapidly changing information;
requires integration of dynamic knowledge retrieval and multi-hop reasoning, testing robustness to changing information.
Images + Text Dyn-VQA
MMBench Liu et al. (2025)
Covers various tasks, including visual question answering, image captioning, and cross-modal retrieval;
focuses on integrating information across multiple modalities (vision, text, audio) for general multimodal understanding.
Images + Text + Audio MMBench
ScienceQA Saikh et al. (2022)
Contains 21,208 questions; tests scientific reasoning abilities using multimodal contexts (text, diagrams, images);
requires deep understanding and integration of information, scientific reasoning, and diagram understanding.
Images + Diagrams + Text ScienceQA
SK-VQA Su et al. (2024c)
Offers over 2 million question-answer pairs;
focuses on synthetic knowledge generation, multimodal reasoning, and external knowledge integration;
used for training context-augmented multimodal large language models.
Images + Text SK-VQA
Text and Knowledge-Focused TriviaQA Joshi et al. (2017)
Provides 650K question-answer pairs; a reading comprehension dataset where questions are based on evidence documents;
can be adapted to multimodal RAG by providing images or other visual context related to the trivia;
focuses on knowledge retrieval and factual accuracy.
Text TriviaQA
Natural Questions Kwiatkowski et al. (2019)
Contains 307,373 training examples; questions are naturally occurring Google search queries;
adapting this for multimodal RAG involves finding or creating relevant visual contexts for the queries;
emphasizes knowledge retrieval and handling noisy or ambiguous queries.
Text Natural Questions
SMMQG Wu et al. (2024a)
Includes 1,024 question-answer pairs; focuses on synthetic data generation,
multimodal retrieval-augmented generation, and control over question style and modality.
Images + Text SMMQG
Appendix C Metrics
Accuracy: Accuracy is typically defined as the ratio of correctly predicted instances to the total instances. In retrieval-based tasks, Top-K Accuracy is defined as:
[…]
FID (Fréchet inception distance): FID is a metric used to assess the quality of images created by a generative model. The formula for FID is:
FID $=\|\mu_{r}-\mu_{g}\|^{2}+tr(\Sigma_{r}+\Sigma_{g}-2\sqrt{\Sigma_{r}\Sigma_{g}})$ (6)
where $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of real images’ feature representations, respectively. $\mu_{g}$ and $\Sigma_{g}$ are the mean and covariance of generated images’ feature representations, respectively. To extract features, InceptionV3 is typically used.
ROUGE-N (N-gram Recall): The ROUGE metric is commonly used to evaluate text summarization and generation. ROUGE-N measures the overlap of N-grams between the generated and reference text. The formula for ROUGE-N is:
ROUGE-N $=\frac{\sum_{\text{gram}_{N}\in\text{Ref}}\text{Count}_{\text{match}}(\text{gram}_{N})}{\sum_{\text{gram}_{N}\in\text{Ref}}\text{Count}(\text{gram}_{N})}$ (7)
ROUGE-L measures the longest common subsequence (LCS) between generated and reference text. The formula for ROUGE-L is:
ROUGE-L $=\frac{LCS(X,Y)}{|Y|}$ (8)
BLEU: BLEU is another metric used for assessing text generation. The formula for calculating BLEU is:
$\text{BLEU}(p_{n},\text{BP})=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$ (9)
Here, $p_{n}$ represents the precision of n-grams, $w_{n}$ denotes the weight assigned to the n-gram precision, and the Brevity Penalty (BP) is defined as:
$\text{BP}=\begin{cases}1&\text{length}>\text{$rl$}\\ \exp\left(1-\frac{\text{$rl$}}{\text{$cl$}}\right)&\text{length}\leq\text{$rl$}\end{cases}$ (10)
Here, $rl$ represents the reference length and $cl$ represents the candidate length.
BERTScore: BERTScore is a metric for evaluating the quality of text generation, based on the similarity between the contextual embeddings of words in the candidate and reference texts. The formula for calculating BERTScore is:
$\text{BERTScore}(c,r)=\frac{1}{|c|}\sum_{i=1}^{|c|}\max_{j=1}^{|r|}\cos(\mathbf{e}_{i},\mathbf{e}_{j})$ (11)
$c$ is the candidate sentence, and $r$ is the reference sentence. $e_{i}$ and $e_{j}$ are the embeddings (e.g., from BERT) for words $c_{i}$ and $r_{j}$ in the candidate and reference sentences, respectively.
CLIPScore: CLIPScore is a metric that evaluates the similarity between the text and an image by using the CLIP model. The formula for calculating CLIPScore is:
$\text{CLIPScore}=\frac{\cos(\mathbf{t},\mathbf{i})}{\|\mathbf{t}\|\cdot\|\mathbf{i}\|}$ (12)
where t and i are text and image embedding respectively.
Mean Reciprocal Rank (MRR): MRR is a metric used to evaluate the performance of systems that return a list of results, such as search engines or recommendation systems. MRR measures the rank position of the first relevant result in the returned list. The formula for calculating MRR is:
$\text{MRR}=\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{\text{rank}_{q}}$ (13)
where $Q$ is the total number of queries. $rank_{q}$ is the rank of the first relevant result for query $q$.
Appendix D Loss Function
InfoNCE (Information Noise Contrastive Estimation): The InfoNCE loss is commonly used in self-supervised learning, especially in contrastive learning methods. The formula for InfoNCE loss is:
$\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(z_{i},z_{j})/\tau)}{\sum_{k=1}^{K}\exp(\text{sim}(z_{i},z_{k})/\tau)}$ (14)
where $z_{i}$ and $z_{j}$ are the embeddings of a positive pair and $\tau$ is a temperature parameter.
GAN (Generative Adversarial Network): The GAN loss consists of two parts: the discriminator loss and the generator loss. The discriminator loss formula is:
$\scriptstyle\mathcal{L}_{D}=-\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]-\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (15)
where $x$ is a real sample from the data distribution. $G(z)$ is the generated sample from the generator, where $z$ is a noise vector. $D(x)$ is the probability that $x$ is real.
The Generator loss formula is:
$\mathcal{L}_{G}=-\mathbb{E}_{z\sim p_{z}(z)}[\log D(G(z))]$ (16)
Triplet Loss: Triplet Loss is used in metric learning to ensure that similar data points are closer together while dissimilar ones are farther apart in the embedding space. The Triplet loss formula is:
$\scriptstyle\mathcal{L}=\sum_{i=1}^{N}\max(0,\|f(x_{a}^{i})-f(x_{p}^{i})\|^{2}-\|f(x_{a}^{i})-f(x_{n}^{i})\|^{2}+\alpha)$ (17)
where $x_{a}^{i}$ is the anchor sample. $x_{p}^{i}$ and $x_{n}^{i}$ are the positive and negative samples respectively. $f(x)$ is the neural network.

### v2

@@
 (2024), sophisticated reasoning Wei et al.
-(2024), In-context Learning Brown et al.
+(2024c), In-context Learning Brown et al.
 (2020), and multilingual machine translation Zhu et al.
@@
 Despite their remarkable achievements, LLMs face significant challenges, including hallucination, outdated internal knowledge, and a lack of verifiable reasoning Huang et al.
-(2024a); Xu et al.
+(2024); Xu et al.
 (2024b).
@@
 The retriever leverages embedding models Chen et al.
-(2024b); Rau et al.
+(2024a); Rau et al.
 (2024) to identify relevant passages from external knowledge bases and optionally applies re-ranking techniques to improve retrieval precision Dong et al.
@@
 These retrieved passages are then passed to the generator, incorporating this external context to produce informed responses.
-Recent advancements in RAG frameworks Asai et al.
-(2023); An et al.
-(2024); Lee et al.
-(2024); Liu et al.
-(2024c) have introduced iterative inference processes, enabling more accurate and reliable answers by refining both retrieval and generation stages.
+Recent advancements in RAG frameworks, such as planning-guided retrieval Lee et al.
+(2024), agentic RAG An et al.
+(2024), and feedback-driven iterative refinement Liu et al.
+(2024c); Asai et al.
+(2023), further enhance both retrieval and generation stages.
 However, traditional RAG architectures are primarily designed for textual information, limiting their ability to address multimodal challenges that require integrating diverse data formats.
@@
 Parallel to these developments, significant advances in multimodal learning have reshaped artificial intelligence by enabling systems to integrate and analyze heterogeneous data sources for a holistic representation of information.
-The introduction of CLIP (Contrastive Language-Image Pre-training) Radford et al.
+The introduction of CLIP (Contrastive Language-Image Pretraining) Radford et al.
 (2021) marked a pivotal moment in connecting visual and textual information through contrastive learning, inspiring numerous subsequent models and applications Alayrac et al.
@@
 (2024), demonstrating the value of multimodal approaches.
-By enabling systems to process and understand diverse data types such as text, images, audio, and video, multimodal learning has become essential to advancing artificial general intelligence (AGI) Song et al.
+By enabling systems to process and understand diverse data types such as text, images, audio, and video, multimodal learning plays a key role in advancing artificial general intelligence (AGI) Song et al.
 (2025).
-Multimodal RAG
-The extension of LLMs to multimodal LLMs (MLLMs) has further expanded their capabilities, allowing them to process, reason, and generate outputs across diverse modalities Liu et al.
-(2023a); Team et al.
-(2024); Li et al.
-(2023b).
-For instance, GPT-4 OpenAI et al.
-(2024) demonstrates human-level performance in various benchmarks by accepting inputs in both text and images, marking a significant milestone in multimodal perception and interaction.
-Building on this foundation, multimodal RAG systems extend traditional RAG frameworks by incorporating multimodal knowledge sources, such as images and audio, to provide enriched context for generation Hu et al.
-(2023); Chen et al.
-(2022a).
-This integration enhances the precision of generated outputs while leveraging multimodal cues to improve the reasoning capabilities of MLLMs.
-The general flow of the Multimodal RAG pipeline is illustrated in Figure 1.
-However, these multimodal systems also present unique challenges, including determining which modalities to retrieve, effectively fusing diverse data types, and addressing the complexities of cross-modal relevance Zhao et al.
-(2023).
-Task Formulation
-We present a mathematical formulation of the general task for multimodal RAG systems.
-These systems generate a multimodal response, denoted as $r$, in response to a query $q$, which is typically in textual format.
-Let $D=\{d_{1},d_{2},...,d_{n}\}$ be a corpus consisting of $n$ multimodal documents.
-Each document $d_{i}$ in $D$ is associated with a modality $M_{d_{i}}$ and is processed by a modality-specific encoder $Enc_{M_{d_{i}}}$:
-$z_{i}=Enc_{M_{d_{i}}}(d_{i})$ (1)
-The set of all encoded representations is denoted by:
-$Z=\{z_{1},z_{2},...,z_{n}\}$ (2)
-Modality-specific encoders map different modalities into a shared semantic space for cross-modal alignment.
-A retrieval model $R$ assesses the relevance of each encoded document representation $z$ with respect to the query $q$, represented as $R(q,z)$.
-To construct the retrieval-augmented multimodal context, the retrieval model selects the most relevant documents based on a modality-specific threshold:
-$X=\{d_{i}\mid s(e_{q},z_{i})\geq\tau_{M_{d_{i}}}\}$ (3)
-where $\tau_{M_{d_{i}}}$ is a relevancy threshold for the modality of $M_{d_{i}}$, $e_{q}$ is the encoded representation of $q$ in the shared semantic space, and $s$ is a scoring function that measures the relevance between the encoded query and document representations.
-The generative model $G$ produces the final multimodal response by conditioning on both the user query $q$ and the retrieved documents $X$ as context:
-$r=G(q,X)$ (4)
-Related Works
-As the field of multimodal RAGs is newly introduced and evolving rapidly, especially in recent years, there is a pressing need for a comprehensive survey that explores the current innovations and frontiers of these systems.
-While more than ten surveys have been published on RAG-related topics such as Agentic RAG Singh et al.
-(2025), none provide a detailed and comprehensive overview of advancements in multimodal RAGs.
-The only related survey to date Zhao et al.
-(2023) categorizes multimodal RAGs by grouping relevant papers based on their applications and modalities.
-However, our survey provides a more detailed and different innovation-driven perspective, offering a detailed taxonomy and exploring emerging trends and challenges in depth.
-Furthermore, significant advancements have been made in the field since its publication, and interest in this topic has grown substantially within the research community.
-In this survey, we review over 100 papers on multimodal RAGs published in recent years, primarily from the ACL Anthology and other repositories such as the ACM Digital Library.
-Contributions
-In this work, (i) we provide a comprehensive review of the multimodal RAG field, covering task formulation, datasets, benchmarks, tasks and domain-specific applications, evaluation, and key innovations across retrieval, fusion, augmentation, generation, training strategies, and loss functions.
-(ii) We introduce a precise structured taxonomy (Figure 2) that categorizes state-of-the-art models based on their primary contributions, highlighting methodological advancements and emerging frontiers.
-(iii) To support further research, we make resources, including datasets, benchmarks, and key innovations, publicly available.
-(iv) We identify current research trends and knowledge gaps, providing insights and recommendations to guide future advancements in this evolving field.
 Figure 1: Overview of the multimodal retrieval-augmented generation (RAG) pipeline, highlighting the advancements and techniques employed at each stage.
@@
 Noise management techniques are also applied to improve training stability and robustness.
+Multimodal RAG
+The extension of LLMs to multimodal LLMs (MLLMs) has further expanded their capabilities, allowing them to process, reason, and generate outputs across diverse modalities Liu et al.
+(2023a); Team et al.
+(2024); Li et al.
+(2023b).
+For example, GPT-4 OpenAI et al.
+(2024) demonstrates human-level performance in various benchmarks by processing both text and images, marking a significant milestone in multimodal perception and interaction.
+Building on this foundation, multimodal RAG systems extend traditional RAG frameworks by incorporating multimodal knowledge sources, such as images and audio, to provide enriched context for generation Hu et al.
+(2023); Chen et al.
+(2022a).
+This integration enhances the precision of generated outputs while leveraging multimodal cues to improve the reasoning capabilities of MLLMs.
+However, these multimodal systems also present unique challenges, including determining which modalities to retrieve, effectively fusing diverse data types, and addressing the complexities of cross-modal relevance Zhao et al.
+(2023a).
+Figure 1 illustrates the general pipeline of these systems.
+Task Formulation
+A mathematical formulation of the general task for multimodal RAG is presented in this section. These systems generate a multimodal response, denoted as $r$, in response to a multimodal query $q$.
+Let $D=\{d_{1},d_{2},...,d_{n}\}$ be a multimodal corpus. Each document $d_{i}\in D$ is associated with a modality $M_{d_{i}}$ and processed by a modality-specific encoder, yielding:
+$z_{i}=Enc_{M_{d_{i}}}(d_{i})$ (1)
+The set of all encoded representations is denoted by $Z=\{z_{1},z_{2},...,z_{n}\}$.
+Modality-specific encoders map different modalities into a shared semantic space for cross-modal alignment.
+A retrieval model $R$ assesses the relevance of each encoded document representation $z$ with respect to the query $q$, represented as $R(q,z)$.
+To construct the retrieval-augmented multimodal context, the retrieval model selects the most relevant documents based on a modality-specific threshold:
+$X=\{d_{i}\mid s(e_{q},z_{i})\geq\tau_{M_{d_{i}}}\}$ (2)
+where $\tau_{M_{d_{i}}}$ is a relevancy threshold for the modality of $M_{d_{i}}$, $e_{q}$ is the encoded representation of $q$ in the shared semantic space, and $s$ is a scoring function that measures the relevance between the encoded query and document representations.
+The generative model $G$ produces the final multimodal response, given the user query $q$ and the retrieved documents $X$ as context, denoted as $r=G(q,X)$.
+Related Works
+As the field of multimodal RAGs is newly introduced and evolving rapidly, especially in recent years, there is a pressing need for a comprehensive survey that explores the current innovations and frontiers of these systems.
+While more than ten surveys have been published on RAG-related topics such as Agentic RAG Singh et al.
+(2025), none provide a detailed and comprehensive overview of advancements in multimodal RAGs.
+The only related survey to date Zhao et al.
+(2023a) categorizes multimodal RAGs by grouping relevant papers based on their applications and modalities.
+However, our survey provides a more detailed and innovation-driven perspective, offering a detailed taxonomy and exploring emerging trends and challenges in depth.
+Furthermore, significant advancements have been made in the field since its publication, and interest in this topic has grown substantially within the research community.
+In this survey, we review over 100 papers on multimodal RAGs published in recent years, primarily from the ACL Anthology and other repositories such as the ACM Digital Library.
+Contributions
+In this work, (i) we provide a comprehensive review of multimodal RAG, covering task formulation, datasets, benchmarks, tasks and domain-specific applications, evaluation, and key innovations in retrieval, fusion, augmentation, generation, training strategies, and loss functions.
+(ii) We introduce a precise structured taxonomy (Figure 2) that categorizes state-of-the-art models based on their primary contributions, highlighting methodological advancements and emerging frontiers.
+(iii) To support further research, we make resources, including datasets, benchmarks, and key innovations, publicly available.
+(iv) We identify current research trends and knowledge gaps, providing insights and recommendations to guide future advancements in this evolving field.
 2 Datasets and Benchmarks
@@
 Advanced retrieval benchmarks such as RAG-Check Mortaheb et al.
-(2025a) evaluate retrieval relevance and system reliability, while specialized assessments like Counterfactual VQA Niu et al.
+(2025a) evaluate retrieval relevance and system reliability, while specialized assessments such as Counterfactual VQA Niu et al.
 (2021) test robustness against adversarial inputs.
@@
 (2024d) examine the cascading effects of errors on RAG systems.
-Additional details about datasets, benchmarks, and their categorization are presented in Table 1 and Table 2 in Appendix B.
+Additional details about datasets, benchmarks, and their categorization are presented in Table 1 and Table 2 in Appendix (§B).
+{forest}
+forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2em, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0.4em, inner ysep=0.6em, line width=0.8pt, text width=8.5em, where level=1 text width=5em, inner xsep=0.3em, inner ysep=0.5em, font=, , where level=2 text width=10em, , where level=3 text width=7em, , ver/.style= fill=white!50, rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=8em , leaf/.style= text opacity=1, inner sep=2pt, fill opacity=.5, fill=green!20, text=black, text width=44.5em font=, inner xsep=0.4em, inner ysep=0.6em, draw, , , [Multimodal RAG, ver [ Retrieval
+Strategies
+ (§4.1), fill=yellow!20 [ Efficient-Search and
+Similarity Retrieval
+ (§4.1.1), fill=red!15 [ Maximum Inner Product Search (MIPS), fill=blue!15 [ TPU-KNN Chern et al.
+(2022), ScaNN Guo et al.
+(2020), MAXSIM score Chan and Ng (2008), ADQ Li et al.
+(2024a), Zhang et al.
+(2023a), BanditMIPS Tiwari et al.
+(2024), MUST Wang et al.
+(2024c), FARGO Zhao et al.
+(2023b), MuRAG Chen et al.
+(2022a), RA-CM3 Yasunaga et al.
+(2023), Nguyen et al.
+(2024), Graph-based ANNs Zhang et al.
+(2024a), Zhai et al.
+(2023), Deeperimpact Basnet et al.
+(2024), RetrievalAttention Liu et al.
+(2024a), FactMM-RAG Sun et al.
+(2024b) , leaf, text width=53em ] ] [ Multimodal
+Encoders, fill=blue!15 [ CLIP Radford et al.
+(2021), BLIP Li et al.
+(2022a), MARVEL Zhou et al.
+(2024c), ALIGN Jia et al.
+(2021), FLAVA Singh et al.
+(2022), UniVL-DR Liu et al.
+(2023b), UniIR Wei et al.
+(2024a), GME Zhang et al.
+(2024i), VISTA Zhou et al.
+(2024b), ColPali Faysse et al.
+(2024), InternVideo Wang et al.
+(2022), Ovis Lu et al.
+(2024) Mi-RAG Adjali et al.
+(2024) , leaf, text width=53em ] ] ] [ Modality-Centric
+Retrieval (§4.1.2), fill=red!15 [ Text-Centric, fill=blue!15 [ Contriever Izacard et al.
+(2022), GTE Li et al.
+(2023c), Re-Imagen Chen et al.
+(2022b), BM25 Robertson and Zaragoza (2009), MiniLM Wang et al.
+(2020a), BGE-M3 Chen et al.
+(2024b), CapRet Shohan et al.
+(2024), OMG-QA Nan et al.
+(2024), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al.
+(2024b), RAFT Zhang et al.
+(2024h), CRAG Yan et al.
+(2024), M2-RAG Ma et al.
+(2024c) , leaf, text width=53em ] ] [ Vision-Centric, fill=blue!15 [ VQA4CIR Feng et al.
+(2023), Unifashion Zhao et al.
+(2024), Jang et al.
+Jang et al.
+(2024), Pic2word Saito et al.
+(2023), eClip Kumar and Marttinen (2024), RAMM Yuan et al.
+(2023), Joshi et al.
+Joshi et al.
+(2024), VISA Ma et al.
+(2024b), ImgRet Shohan et al.
+(2024), EchoSight Yan and Xie (2024), Xue et al.
+Xue et al.
+(2024) , leaf, text width=53em ] ] [ Video-Centric, fill=blue!15 [ iRAG Arefeen et al.
+(2024), VideoRAG Ren et al.
+(2025), VideoRAG Jeong et al.
+(2025), T-Mass Wang et al.
+(2024b), MV-Adapter Jin et al.
+(2024), OmAgent Zhang et al.
+(2024e), CM2 Kim et al.
+(2024), Video-RAG Luo et al.
+(2024b), CTCH Shen et al.
+(2024), RTime Du et al.
+(2024), VideoMAE Tong et al.
+(2022), DrVideo Ma et al.
+(2024d) , leaf, text width=53em ] ] [ Document Retrieval, fill=blue!15 [ ColPali Faysse et al.
+(2024), ColQwen2 Wang et al.
+(2024d), M3DocVQA Cho et al.
+(2024), ViTLP Mao et al.
+(2024), DocLLM Wang et al.
+(2024a), CREAM Zhang et al.
+(2024b), mPLUG-DocOwl 1.5 Hu et al.
+(2024a), mPLUG-DocOwl 2 Hu et al.
+(2024b), VisDom Suri et al.
+(2024), DSE Ma et al.
+(2024a) , leaf, text width=53em ] ] ] [ Re-ranking
+Strategies (§4.1.3), fill=red!15 [ Optimized Example Selection, fill=blue!15 [ MSIER Luo et al.
+(2024a), Hybrid RAG Su et al.
+(2024a), RULE Xia et al.
+(2024b), RAMM Yuan et al.
+(2023), M2RAAP Dong et al.
+(2024b) , leaf, text width=53em ] ] [ Relevance Score
+Evaluation, fill=blue!15 [ RAG-Check Mortaheb et al.
+(2025a, b), UniRaG Zhi Lim et al.
+(2024), MR2AG Zhang et al.
+(2024g), LDRE Yang et al.
+(2024), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al.
+(2024), OMG-QA Nan et al.
+(2024), EchoSight Yan and Xie (2024), EgoInstructor Xu et al.
+(2024a) , leaf, text width=53em ] ] [ Filtering
+Mechanisms, fill=blue!15 [ MAIN-RAG Chang et al.
+(2024), MM-Embed Lin et al.
+(2024a), GME Zhang et al.
+(2024i), MuRAR Zhu et al.
+(2025) RAFT Zhang et al.
+(2024h) , leaf, text width=53em ] ] ] ] [ Fusion
+Mechanisms
+ (§4.2), fill=yellow!20 [ Score Fusion
+and Alignment
+ (§4.2.1), fill=red!15 [ M3 Cai et al.
+(2025)  Zhi Lim et al.
+(2024),  Sharifymoghaddam et al.
+(2024), REVEAL Hu et al.
+(2023), RAG-Driver Yuan et al.
+(2024), C3Net Zhang et al.
+(2024c), LLM-RA Jian et al.
+(2024),  Riedler and Langer (2024), VISA Ma et al.
+(2024b), MA-LMM He et al.
+(2024),  Xue et al.
+(2024), RA-BLIP Ding et al.
+(2024b), Re-IMAGEN Chen et al.
+(2022b), MegaPairs Zhou et al.
+(2024a), Wiki-LLaVA Caffagni et al.
+(2024), VISRAG Yu et al.
+(2024) , leaf, text width=62.0em ] ] [ Attention-Based
+Mechanisms (§4.2.2), fill=red!15 [ RAMM Yuan et al.
+(2023), EMERGE Zhu et al.
+(2024b), MORE Cui et al.
+(2024), RAGTrans Cheng et al.
+(2024), AlzheimerRAG Lahiri and Hu (2024), MV-Adapter Jin et al.
+(2024),  Xu et al.
+(2024a),  Kim et al.
+(2024), M2-RAAP Dong et al.
+(2024b), Mu-RAG Chen et al.
+(2022a), Ou et al.
+Ou et al.
+(2025), CADMR Khalafaoui et al.
+(2024) , leaf, text width=62.0em ] ] [ Unified Frameworks
+and Projections (§4.2.3), fill=red!15 [ Hybrid-RAG Su et al.
+(2024a), Dense2Sparse Nguyen et al.
+(2024), IRAMIG Liu et al.
+(2024b), M3DocRAG Cho et al.
+(2024), DQU-CIR Wen et al.
+(2024), PDF-MVQA Ding et al.
+(2024c), SAM-RAG Zhai (2024), UFineBench Zuo et al.
+(2024),  Li et al.
+(2022a) , leaf, text width=62.0em ] ] ] [ Augmentation
+Techniques
+ (§4.3), fill=yellow!20 [ Context-Enrichment (§4.3.1), fill=red!15 [ EMERGE Zhu et al.
+(2024b), MiRAG Adjali et al.
+(2024), Wiki-LLaVA Caffagni et al.
+(2024), Video-RAG Luo et al.
+(2024b), Img2Loc Zhou et al.
+(2024e),  Xue et al.
+(2024) , leaf, text width=62.0em ] ] [ Adaptive and
+Iterative
+Retrieval (§4.3.2), fill=red!15 [ SKURG Yang et al.
+(2023), IRAMIG Liu et al.
+(2024b), OMG-QA Nan et al.
+(2024), SAM-RAG Zhai (2024), MMed-RAG Xia et al.
+(2024a), OmniSearch Li et al.
+(2024b), mR2AG Zhang et al.
+(2024f), RAGAR Khaliq et al.
+(2024) , leaf, text width=62.0em ] ] ] [ Generation
+Technique
+s (§4.4), fill=yellow!20 [ In-Context
+Learning (§4.4.1), fill=red!15 [ RMR Tan et al.
+(2024),  Sharifymoghaddam et al.
+(2024), RA-CM3 Yasunaga et al.
+(2023), RAG-Driver Yuan et al.
+(2024), MSIER Luo et al.
+(2024a), Raven Rao et al.
+(2024) , leaf, text width=62.0em ] ] [ Reasoning (§4.4.2), fill=red!15 [ RAGAR Khaliq et al.
+(2024), VisDoM Suri et al.
+(2024), SAM-RAG Zhai (2024), LDRE Yang et al.
+(2024) , leaf, text width=62.0em ] ] [ Instruction
+Tuning (§4.4.3), fill=red!15 [ RA-BLIP Ding et al.
+(2024b), RAGPT Lang et al.
+(2025), mR2AG Zhang et al.
+(2024f), RagVL Chen et al.
+(2024d), Jang et al.
+(2024), MMed-RAG Xia et al.
+(2024a), MegaPairs Zhou et al.
+(2024a), Surf Sun et al.
+(2024a), Rule Xia et al.
+(2024b) , leaf, text width=62.0em ] ] [ Source
+Attribution (§4.4.4), fill=red!15 [ MuRAR Zhu et al. (2025), VISA Ma et al. (2024b), OMG-QA Nan et al. (2024) , leaf, text width=62.0em ] ] ] [ Training
+Strategies
+ (§4.5), fill=yellow!20 [ Alignment (§4.5.1), fill=red!15 [ VISRAG Yu et al.
+(2024), MegaPairs Zhou et al.
+(2024a), SAM-RAG Zhai (2024), EchoSight Yan and Xie (2024), HACL Jiang et al.
+(2024),  Zhi Lim et al.
+(2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al.
+(2024) , leaf, text width=62.0em ] ] [ Robustness (§4.5.3), fill=red!15 [  Buettner and Kovashka (2024), MORE Cui et al.
+(2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al.
+(2024), RA-BLIP Ding et al.
+(2024b), RagVL Chen et al.
+(2024d), RA-CM3 Yasunaga et al.
+(2023) , leaf, text width=62.0em ] ] ] ] ]
+Figure 2: Taxonomy of recent advances in Multimodal RAG. Refer to Appendix (§A) for further details.
 3 Evaluation
-Assessing multimodal RAG models presents significant challenges due to their diverse modalities, intricate architectures, and the diverse evaluation dimensions involved.
-Given the broad spectrum of modalities and system components, evaluation requires a combination of metrics drawn from vision-language models (VLMs), generative AI, and retrieval systems.
-These metrics enable the assessment of various aspects of multimodal RAG, such as text and image generation or information retrieval, either independently or in combination.
-By analyzing prior works, we identified approximately 60 distinct evaluation metrics used in multimodal RAG studies.
-More details, including the formulas for the RAG evaluation metrics, can be found in Appendix C.
+Evaluating multimodal RAG models is complex due to their varied input types and complex structure.
+The evaluation combines metrics from VLMs, generative AI, and retrieval systems to assess capabilities like text/image generation and information retrieval.
+Our review found about 60 different metrics used in the field.
+More details, including the formulas for the RAG evaluation metrics, can be found in Appendix (§C).
 In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
 Retrieval Evaluation
-Evaluates how well a system retrieves relevant information from a database or knowledge source.
-Accuracy, recall, and precision are used to evaluate retrieval and downstream task performance, the F1 score also being used to assess both Recall and Precision simultaneously.
-A more commonly used metric than recall is Recall@K, which measures the proportion of relevant items appearing within the top K retrieved results.
-Another important metric Mean Reciprocal Rank (MRR), which is utilized by Adjali et al.
+Retrieval performance is measured through accuracy, recall, and precision metrics, with an F1 score combining recall and precision.
+Recall@K, which examines relevant items in top K results, is preferred over standard recall.
+Mean Reciprocal Rank (MRR) serves as another key metric for evaluation, which is utilized by Adjali et al.
 (2024); Nguyen et al.
@@
 Modality Evaluation
-These evaluations focus on criteria that rely on modalities.
-Text and image are the most commonly used modalities in multimodal RAG.
-This category of evaluations considers various aspects of these modalities, including the alignment between generated text and images, the fluency of the text, the quality of image captions, and other related factors.
-In text similarity and overlap metrics, Exact Match (EM), BLEU Papineni et al.
-(2002), ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005) are used.
+Modality-based evaluations primarily focus on text and image, assessing their alignment, text fluency, and image caption quality.
+For text evaluation, metrics include Exact Match (EM), BLEU Papineni et al.
+(2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005).
 MultiRAGen Shohan et al.
 (2024) uses Multilingual ROUGE for multilingual settings.
-For image captioning evaluation, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al.
-(2015) is a widely used metric for evaluating image captioning quality.
-It quantifies the alignment between a generated caption and a set of human-written reference captions by leveraging the term frequency-inverse document frequency (TF-IDF) and cosine similarity.
-Yasunaga et al.
+For image captioning, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al.
+(2015) measures caption quality using TF-IDF and cosine similarity Yasunaga et al.
 (2023); Zhao et al.
@@
 (2024); Zhang et al.
-(2024c) use CIDEr for evaluation.
-SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al.
-(2016) is a metric that emphasizes semantic aspects, in contrast to CIDEr.
+(2024c), while SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al.
+(2016) focuses on semantics.
 SPIDEr Liu et al.
-(2017), utilized in Zhang et al.
-(2024c), is a linear combination of both SPICE and CIDEr.
-In semantic alignment evaluations, BERTScore Zhang et al.
-(2020), employed in Sun et al.
+(2017), used in Zhang et al.
+(2024c), combines both metrics.
+For semantic alignment, BERTScore Zhang et al.
+(2020) compares BERT embeddings Sun et al.
 (2024b); Shohan et al.
-(2024), measures semantic similarity by comparing BERT-generated embeddings.
-Chen et al.
+(2024), and evaluates fluency Chen et al.
 (2022a); Zhi Lim et al.
 (2024); Ma et al.
-(2024c) evaluate Fluency by employing BERTScore.
+(2024c).
 CLIP Score Hessel et al.
@@
 (2024); Zhang et al.
-(2024c), measures image-text similarity using the CLIP model Radford et al.
+(2024c), measures image-text similarity using CLIP Radford et al.
 (2021).
-FID (Fréchet Inception Distance) Heusel et al.
-(2017) quantifies image quality by computing the distributional difference between the generated and original image features.
-Yasunaga et al.
+For image quality, FID (Fréchet Inception Distance) Heusel et al.
+(2017) compares feature distributions Yasunaga et al.
 (2023); Zhao et al.
@@
 (2024); Zhang et al.
-(2024c) use FID in their works.
-Zhao et al.
-(2024) use Kernel Inception Distance (KID) Bińkowski et al.
-(2018) for evaluation.
-KID is another metric used to evaluate the quality of generated images.
-It is similar to FID but offers an unbiased estimator.
-Zhi Lim et al.
-(2024) employs the Inception Score (IS) to assess the diversity and quality of generated images based on classification probabilities obtained from the Inception network.
-For the evaluation of audio synthesis, Zhang et al.
-(2024c) assesses overall sound quality (OVL) and relevance to the text (REL) with the assistance of human annotators.
-Additionally, this study employs the Fréchet Audio Distance (FAD) Kilgour et al.
-(2019) as an evaluation metric.
-FAD serves a similar purpose to the FID but is specifically designed for the audio domain.
-FLOPs, execution time, response time, and average retrieval time per query are utilized to assess efficiency and computational performance Nguyen et al.
+(2024c), while KID (Kernel Inception Distance) Bińkowski et al.
+(2018) provides an unbiased alternative.
+Inception Score (IS) evaluates image diversity and quality through classification probabilities Zhi Lim et al.
+(2024).
+For audio evaluation, Zhang et al.
+(2024c) uses human annotators to assess sound quality (OVL) and text relevance (REL), while also employing Fréchet Audio Distance (FAD) Kilgour et al.
+(2019), an audio-specific variant of FID.
+System efficiency is measured through FLOPs, execution time, response time, and retrieval time per query Nguyen et al.
 (2024); Strand et al.
 (2024); Dang (2024); Zhou (2024).
-Zhong et al.
-(2024) employs Spearman’s Rank Correlation (SRC), which calculates the Pearson correlation between the rank values of two variables.
-Several evaluation metrics are domain- and problem-specific.
-Geodesic distance measurement quantifies the model’s accuracy by calculating the geodesic distance between predicted and actual geographical coordinates Zhou et al.
-(2024e).
-Lahiri and Hu (2024) utilizes Clinical Relevance (CR) for evaluation.
+Domain-specific metrics include geodesic distance for geographical accuracy Zhou et al.
+(2024e), and Clinical Relevance for medical applications Lahiri and Hu (2024).
 4 Key Innovations and Methodologies
 4.1 Retrieval Strategy
-Efficient Search and Similarity Retrieval
-Modern multimodal RAG systems first encode diverse modalities (text, images, tables) into a unified embedding space to enable direct cross-modal retrieval.
-The foundational CLIP-Radford et al.
-(2021) or BLIP-based Li et al.
-(2022a) approaches leverage contrastive learning, which aligns text and images into a shared latent space via dual-tower encoders.
-Recent advancements extend this paradigm with novel multimodal retrieval architectures and training strategies Zhou et al.
+4.1.1 Efficient Search and Similarity Retrieval
+Modern multimodal RAG systems encode diverse input modalities into a unified embedding space to enable direct cross-modal retrieval.
+Recent advancements in CLIP-based Radford et al.
+(2021) and BLIP-inspired Li et al.
+(2022a) approaches have driven the evolution of contrastive learning strategies through novel multimodal retrieval architectures and training methodologies Zhou et al.
 (2024c); Wei et al.
-(2025); Zhang et al.
-(2024j).
-Multimodal RAG systems rely on efficient search strategies to retrieve relevant external knowledge from large-scale multimodal memory.
-Since dual-encoder models map multimodal inputs into a shared high-dimensional embedding space, maximum inner product search (MIPS) is commonly used for fast and direct similarity comparisons.
+(2024b); Zhang et al.
+(2024i).
+As these multi-encoder models project different modalities into a shared latent space, multimodal RAGs rely on efficient search strategies to retrieve relevant external knowledge.
+Maximum inner product search (MIPS) variants are widely used for fast and direct similarity comparisons Tiwari et al.
+(2024); Wang et al.
+(2024c); Zhao et al.
+(2023b).
 Systems such as MuRAG Chen et al.
@@
 Large-scale implementations leverage distributed MIPS techniques, such as TPU-KNN Chern et al.
-(2022), to enable high-speed retrieval across extensive datasets.
+(2022), for high-speed retrieval.
 Additionally, ScaNN (Scalable Nearest Neighbors) Guo et al.
@@
 (2024), and approximate KNN methods Caffagni et al.
-(2024) have been adopted for efficient image-text similarity computation.
+(2024) have been adopted for efficient similarity computation.
 Recent advancements in MIPS optimization focus on reducing retrieval latency and improving accuracy.
-These include adaptive quantization Li et al.
-(2024a); Huang et al.
-(2024b), hybrid sparse-dense representations Nguyen et al.
-(2024); Ma et al.
-(2024a); Zhang et al.
+These include adaptive quantization Zhang et al.
+(2023a); Li et al.
+(2024a), hybrid sparse-dense representations Nguyen et al.
+(2024); Zhang et al.
 (2024a), and learned index structures Zhai et al.
@@
 These techniques help optimize retrieval performance by balancing efficiency and precision in multimodal RAG systems.
-Modality-Based Retrieval
+4.1.2 Modality-Based Retrieval
 Modality-aware retrieval techniques optimize retrieval efficiency by leveraging the unique characteristics of each modality.
-Text-centric retrieval remains foundational in multimodal RAG systems.
-Traditional methods like BM25 Robertson and Zaragoza (2009), along with dense retrievers such as BGE-1.5 Xiao et al.
-(2023) and BGE-M3 Chen et al.
-(2024a) dominate text-based evidence retrieval Chen et al.
+Text-Centric Retrieval
+This remains foundational in multimodal RAG systems, with both traditional methods like BM25 Robertson and Zaragoza (2009), and dense retrievers such as MiniLM Wang et al.
+(2020a) and BGE-M3 Chen et al.
+(2024b) dominate text-based evidence retrieval Chen et al.
 (2022b); Suri et al.
@@
 (2024b) employ token-level interaction mechanisms that preserve nuanced textual details to improve precision for multimodal queries, while RAFT Zhang et al.
-(2024i) and CRAG Yan et al.
-(2024) enhance retrieval performance by ensuring precise citation of text spans, thereby optimizing domain-specific question answering within multimodal RAG pipelines.
-Vision-centric retrieval focuses on directly leveraging image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al.
+(2024h) and CRAG Yan et al.
+(2024) enhance retrieval performance by ensuring precise citation of text spans.
+Vision-Centric Retrieval
+This focuses on directly leveraging image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al.
 (2023).
@@
 (2024) retrieve visually similar content by using reference images as queries.
-Composed Image Retrieval (CMI) models Feng et al.
+In addition, composed Image Retrieval (CMI) models Feng et al.
 (2023); Zhao et al.
@@
 (2023) maps visual content to textual descriptions, enabling zero-shot image retrieval.
-Video-centric retrieval extends vision-based techniques by incorporating temporal dynamics and large video-language models (LVLMs), driven by novel frameworks like iRAG Arefeen et al.
+Video-Centric Retrieval
+These methods extend vision-based techniques by incorporating temporal dynamics and large video-language models (LVLMs), driven by novel frameworks like iRAG Arefeen et al.
 (2024), which introduces incremental retrieval for sequential video understanding, and MV-Adapter Jin et al.
@@
 Recent breakthroughs focus on long-context processing: Video-RAG Luo et al.
-(2024b) leverages visually aligned auxiliary texts (OCR/ASR) to enhance retrieval without proprietary models, while VideoRAG Ren et al.
-(2025) employs dual-channel architectures with graph-based knowledge grounding for extreme-length videos.
+(2024b) leverages visually aligned auxiliary texts (OCR/ASR) to enhance retrieval without proprietary models, T-Mass Wang et al.
+(2024b) models text as a stochastic embedding to enhance text-video retrieval, and VideoRAG Ren et al.
+(2025) -employs dual-channel architectures with graph-based knowledge grounding for extreme-length videos.
 For temporal reasoning, CTCH Shen et al.
-(2024) uses contrastive transformer hashing to model long-term dependencies, and RTime Du et al.
+(2024) uses contrastive transformer hashing to model long-term dependencies, while RTime Du et al.
 (2024) introduces reversed-video hard negatives to rigorously benchmark temporal causality.
 Meanwhile, for complex video understanding, OmAgent Zhang et al.
-(2024e) adopts a divide-and-conquer framework, while DRVideo Ma et al.
-(2024d) addresses long-video understanding by a document-based retrieval approach.
+(2024e) adopts a divide-and-conquer framework, and DRVideo Ma et al.
+(2024d) addresses long-video understanding with a document-based retrieval approach.
 Document Retrieval and Layout Understanding
 Recent research has advanced beyond traditional uni-modal retrieval methods toward models that directly process entire document pages.
-These approaches integrate textual, visual, and structural elements like tables, images, font styles, and page layouts to enhance retrieval performance in complex documents.
+These approaches integrate textual, visual, and structural elements like tables, font styles, and page layouts to enhance retrieval performance in complex documents.
 ColPali Faysse et al.
@@
 Models like ColQwen2 Wang et al.
-(2024b); Khattab and Zaharia (2020) and M3DocVQA Cho et al.
+(2024d); Khattab and Zaharia (2020) and M3DocVQA Cho et al.
 (2024) extend this paradigm with dynamic resolution handling and holistic multi-page reasoning.
@@
 (2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies.
-Together, these approaches highlight a shift toward holistic, layout-sensitive retrieval that leverages visual semantics and generative pre-training.
-Re-ranking and Selection Strategies
-Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also refining and prioritizing retrieved candidates.
+4.1.3 Re-ranking and Selection Strategies
+Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also prioritizing retrieved candidates.
 Re-ranking and selection strategies enhance retrieval quality by optimizing example selection, refining relevance scoring, and applying filtering mechanisms.
-Optimized example selection techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection strategies Luo et al.
-(2024a).
-Su et al.
-(2024a) refine multimodal inputs using probabilistic control keywords to improve credibility, while RULE Xia et al.
-(2024b) calibrates retrieved context selection through statistical methods like the Bonferroni correction to mitigate factuality risks.
-Furthermore, selective fusion mechanisms dynamically refine the retrieval outcomes Yuan et al.
-(2023), and key-frame extraction techniques ensure diversity in video-based retrieval through clustering-based frame selection Dong et al.
+Optimized Example Selection
+These techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection strategies Luo et al.
+(2024a); Yuan et al.
+(2023).
+For instance, Su et al.
+(2024a) enhance multimodal inputs using probabilistic control keywords to improve credibility, RULE Xia et al.
+(2024b) calibrates retrieved context selection via statistical methods like the Bonferroni correction to mitigate factuality risks, and clustering-based key-frame selection ensures diversity in video-based retrieval Dong et al.
 (2024b).
+Relevance Score Evaluation
 Several methods employ advanced scoring mechanisms to improve retrieval relevance Mortaheb et al.
-(2025b, a).
-Multimodal similarity measures, including negative log-likelihood (NLL), structural similarity index measure (SSIM) Wang et al.
-(2020), normalized cross-correlation (NCC), and BERTScore Zhang et al.
-(2020), aid in re-ranking retrieved documents based on textual and visual similarity.
+(2025b, a); Zhi Lim et al.
+(2024).
+Multimodal similarity measures, including structural similarity index measure (SSIM) Wang et al.
+(2020b), normalized cross-correlation (NCC), and BERTScore Zhang et al.
+(2020), aid in re-ranking retrieved documents.
 Cross-encoders trained on sequence classification tasks refine relevance scoring within retrieval pipelines Zhi Lim et al.
 (2024), while hierarchical post-processing techniques integrate retrieval, passage-level, and answer confidence scores for improved ranking Zhang et al.
-(2024h).
+(2024g); Yan and Xie (2024); Xu et al.
+(2024a).
 LDRE Yang et al.
@@
 (2024) and OMG-QA Nan et al.
-(2024) integrate traditional ranking functions like BM25 Robertson and Zaragoza (2009).
-Advanced re-ranking techniques also incorporate Q-Former-based image-to-text correspondence for fine-grained retrieval Yan and Xie (2024); Xu et al.
-(2024a).
-Advanced filtering methods are crucial for enhancing multi-modal RAG performance by ensuring that only high-quality, relevant data is processed.
-Hard negative mining techniques, as seen in GME Zhang et al.
-(2024j) and MM-Embed Lin et al.
-(2024a), use modality-aware sampling and synthesized negatives from imbalanced data to mitigate modality bias and refine contrastive learning.
-Similarly, consensus-based filtering approaches, exemplified by MuRAR Zhu et al.
-(2024d) and ColPali Faysse et al.
-(2024), employ source attribution and multi-vector mapping to filter out low-similarity candidates and reduce hallucinations in enterprise QA.
-Moreover, dynamic modality filtering methods—including RAFT Zhang et al.
-(2024i), Img2Loc Zhou et al.
+(2024) incorporate traditional ranking functions like BM25 Robertson and Zaragoza (2009).
+Filtering Mechasnism
+This ensures high-quality retrieval by eliminating irrelevant data.
+Hard negative mining, as used in GME Zhang et al.
+(2024i) and MM-Embed Lin et al.
+(2024a), mitigates modality bias through modality-aware sampling and synthesized negatives.
+Similarly, consensus-based filtering approaches, seen in MuRAR Zhu et al.
+(2025) and ColPali Faysse et al.
+(2024), employs source attribution and multi-vector mapping to filter out low-similarity candidates.
+Moreover, dynamic modality filtering methods, such as RAFT Zhang et al.
+(2024h), Img2Loc Zhou et al.
 (2024e), and MAIN-RAG Chang et al.
-(2024)—train retrievers to disregard confusing data, thereby enhancing the discriminative capacity and overall robustness of multi-modal retrieval systems.
+(2024), train retrievers to disregard confusing data, thereby enhancing the discriminative capacity and overall robustness of multi-modal retrieval systems.
 4.2 Fusion Mechanisms
-Score Fusion and Alignment
-Recent multimodal RAG models emphasize aligning embeddings or fusing scores across different modalities to enable adaptive retrieval.
+4.2.1 Score Fusion and Alignment
+Models in this category utilize distinct strategies to align multimodal representations.
 Zhi Lim et al.
-(2024) transform text, tables, and images into a unified textual representation using a cross-encoder trained for relevance scoring.
-To address multimodal LLM input constraints, Sharifymoghaddam et al.
-(2024) introduce interleaved image-text pairs, merging multiple few-shot images vertically for compatibility with models like LLaVA Liu et al.
-(2023a) while aligning modalities via CLIP score fusion Hessel et al.
+(2024) convert text, tables, and images into a single textual format using a cross-encoder trained for relevance scoring.
+Sharifymoghaddam et al.
+(2024) introduce interleaved image–text pairs that vertically merge multiple few-shot images (as in LLaVA Liu et al.
+(2023a)), while aligning modalities via CLIP score fusion Hessel et al.
 (2021) and BLIP feature fusion Li et al.
 (2022a).
-For fine-grained alignment, REVEAL Hu et al.
-(2023) injects retrieval scores into attention layers and minimizes L2-Norm distances between query and knowledge embeddings, and MA-LMM He et al.
-(2024) aligns video-text embeddings via a BLIP-inspired Query Transformer (Q-Former)Li et al.
+Riedler and Langer (2024), Wiki-LLaVA Caffagni et al.
+(2024), C3Net Zhang et al.
+(2024c), and MegaPairs Zhou et al.
+(2024a), embed images and queries into a shared CLIP space.
+VISA Ma et al.
+(2024b) employs the Document Screenshot Embedding (DSE) model to align textual queries with visual document representations by encoding both into a shared embedding space.
+REVEAL Hu et al.
+(2023) injects retrieval scores into attention layers to minimize L2-norm differences between query and knowledge embeddings, and MA-LMM He et al.
+(2024) aligns video-text embeddings via a BLIP-inspired Query Transformer Li et al.
 (2022a).
-In contrast, LLM-RA Jian et al.
-(2024) concatenates text and visual embeddings into joint queries to reduce retrieval noise.
-RA-BLIP Ding et al.
-(2024b) employs a 3-layer BERT-based adaptive fusion module to unify visual-textual semantics.
-CLIP-based alignment strategies are widely adopted; Riedler and Langer (2024) embed images and questions into a shared CLIP space, while VISA Ma et al.
-(2024b) uses the Document Screenshot Embedding (DSE) model as the retrieval component of the RAG system to align and fuse textual queries with visual document representations.
-DSE directly encodes both textual queries and document screenshots into a shared embedding space, allowing efficient retrieval based on cosine similarity.
+LLM-RA Jian et al.
+(2024) concatenates text and visual embeddings into joint queries to reduce retrieval noise, while RA-BLIP Ding et al.
+(2024b) employs a 3-layer BERT-based adaptive fusion module to unify visual–textual semantics.
 Xue et al.
-(2024) map object-predicate pairs into a shared semantic space using a prototype-based embedding network (PENET) Zheng et al.
-(2023), aligning visual features with textual prototypes.
+(2024) use a prototype-based embedding network Zheng et al.
+(2023) to map object-predicate pairs into a shared semantic space, aligning visual features with textual prototypes.
 Re-IMAGEN Chen et al.
-(2022b) refines text-to-image synthesis with interleaved classifier-free guidance during diffusion sampling, balancing creativity and entity fidelity.
-To improve multimodal alignment, MegaPairs Zhou et al.
-(2024a) integrates CLIP-based and MLLM-based retrieval through score fusion.
-Meanwhile, Wiki-LLaVA Caffagni et al.
-(2024) employs CLIP encoders for hierarchical document retrieval.
-VISRAG Yu et al.
-(2024) refines alignment via position-weighted mean pooling on VLM hidden states, prioritizing later tokens for relevance.
-RAG-Driver Yuan et al.
-(2024) aligns visual-language embeddings via visual instruction tuning and an MLP projector.
-Attention-Based Mechanisms
-Attention-based fusion dynamically weights cross-modal interactions for task-specific reasoning.
+(2022b) balances creativity and entity fidelity in text-to-image synthesis via interleaved classifier-free guidance during diffusion sampling.
+To improve multimodal alignment, VISRAG Yu et al.
+(2024) enhances alignment with position-weighted mean pooling on VLM hidden states, prioritizing later tokens for relevance, and RAG-Driver Yuan et al.
+(2024) aligns visual–language embeddings using visual instruction tuning and an MLP projector.
+4.2.2 Attention-Based Mechanisms
+Attention-based methods dynamically weight cross-modal interactions to support task-specific reasoning.
+EMERGE Zhu et al.
+(2024b), MORE Cui et al.
+(2024), and AlzheimerRAG Lahiri and Hu (2024) integrate heterogeneous data via cross-attention.
 RAMM Yuan et al.
-(2023) employs a dual-stream co-attention transformer, combining self-attention (intra-modal) and cross-attention (inter-modal) to fuse retrieved biomedical images/texts with input data.
-Similarly, EMERGE Zhu et al.
-(2024b) integrates Electronic Health Records (EHR) data with a cross-attention adaptive network, while MORE Cui et al.
-(2024) uses cross-attention to weigh multimodal results for commonsense reasoning.
-For user-centric tasks, RAGTrans Cheng et al.
-(2024) integrates user-aware attention to fuse social media features, and AlzheimerRAG Lahiri and Hu (2024) applies cross-modal attention to synthesize text, images, and tables.
-Video-text alignment is advanced by MV-Adapter Jin et al.
-(2024), which leverages a Cross Modality Tying (CMT) to align embeddings, and M2-RAAP Dong et al.
-(2024b), enhances video-text fusion using an Auxiliary-Caption-Guided (ACG) strategy.
-It employs Frame-Caption Re-weighting (FCR) to boost frames with high inter-modal consistency and Text-Caption Re-weighting (TCR) to adjust weights based on intra-modal similarity.
-A Mutual-guided (Mug) Alignment Head then filters misaligned features via a dot-product similarity and frame-to-token attention, producing refined frame-specific text representations.
-Generative refinement is demonstrated in Xu et al.
-(2024a), where gated cross-attention conditions text generation on visual features.
-Similarly, C3Net Zhang et al.
-(2024c) enhances alignment between queries and candidates within CLIP’s contrastive space.
-Mu-RAG Chen et al.
-(2022a) uses intermediate fusion with cross-attention for open QA, balancing modality-specific and joint representations.
+(2023) employs a dual-stream co-attention transformer, combining self-attention and cross-attention to fuse retrieved biomedical images/texts with input data.
+RAGTrans Cheng et al.
+(2024) applies user-aware attention to social media features.
+For video-text alignment, MV-Adapter Jin et al.
+(2024) leverages Cross Modality Tying to align embeddings, and M2-RAAP Dong et al.
+(2024b) enhances fusion through an auxiliary caption-guided strategy that re-weights frames and text captions based on intra-modal similarity.
+A mutual-guided alignment head then filters misaligned features using dot-product similarity and frame-to-token attention, generating refined frame-specific text representations.
+Xu et al.
+(2024a) condition text generation on visual features using gated cross-attention, and Mu-RAG Chen et al.
+(2022a) employs intermediate cross-attention for open-domain QA.
 Kim et al.
 (2024) leverage cross-modal memory retrieval with pre-trained CLIP ViT-L/14 to map video-text pairs into a shared space, enabling dense captioning through the attention-based fusion of retrieved memories.
-Unified Frameworks and Projections
-Architectures unifying multimodal inputs through frameworks or projections address scalability.
+4.2.3 Unified Frameworks and Projections
+Unified frameworks and projection methods consolidate multimodal inputs into coherent representations.
 Su et al.
 (2024a) employ hierarchical cross-chains and late fusion for healthcare data, while IRAMIG Liu et al.
-(2024b) iteratively consolidates multimodal results into knowledge representations.
-For document understanding, M3DocRAG Cho et al.
-(2024) flattens multi-page documents into a unified embedding tensor, and PDF-MVQA Ding et al.
-(2024c) combines Region-of-Interest (RoI)-based (VisualBERT) and patch-based (CLIP) VLPMs Long et al.
+(2024b) iteratively integrates multimodal results into unified knowledge representations.
+M3DocRAG Cho et al.
+(2024) flattens multi-page documents into a single embedding tensor, and PDF-MVQA Ding et al.
+(2024c) fuses Region-of-Interest (RoI)-based and patch-based (CLIP) vision-language models Long et al.
 (2022).
 DQU-CIR Wen et al.
-(2024) exemplifies raw-data unification by converting images into text captions for complex queries and overlaying text onto images for simple queries.
-The model then fuses embeddings using MLP-learned weights.
-SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input into unimodal text representation for subsequent processing.
+(2024) unifies raw data by converting images into text captions for complex queries and overlaying text onto images for simple ones, then fusing embeddings via MLP-learned weights.
+SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input into unimodal text for subsequent processing.
 UFineBench Zuo et al.
-(2024) uses a shared granularity decoder for ultra-fine text-person retrieval.
+(2024) utilizes a shared granularity decoder for ultra-fine text–person retrieval.
 Nguyen et al.
-(2024) introduce Dense2Sparse (D2S) projection, converting dense embeddings from models like BLIP/ALBEF Li et al.
-(2022a) into sparse lexical vectors via layer normalization and probabilistic expansion control, optimizing storage and interpretability.
+(2024) introduce Dense2Sparse projection, converting dense embeddings from models like BLIP/ALBEF Li et al.
+(2022a) into sparse lexical vectors using layer normalization and probabilistic expansion control to optimize storage and interpretability.
 4.3 Augmentation Techniques
 Basic RAG systems typically follow a single retrieval step and pass retrieved content directly to the generation phase without further refinement, which can lead to inefficiencies and suboptimal outputs.
-To address this, augmentation techniques have been introduced to enhance the utility of retrieved data before generation, ensuring that the model can better interpret, structure, and integrate multimodal information Gao et al.
+To address this, augmentation techniques have been introduced to refine retrieved data beforehand, improving multimodal interpretation, structuring, and integration Gao et al.
 (2023).
-Context Enrichment
-This focuses on enhancing the relevance of retrieved knowledge for the generation phase by refining external knowledge or retrieved data.
-General approaches such as EMERGE Zhu et al.
-(2024b) integrate entity relationships and semantic descriptions to provide richer context.
+4.3.1 Context Enrichment
+This focuses on enhancing the relevance of retrieved knowledge by refining or expanding retrieved data.
+General approaches incorporate additional contextual elements (e.g., text chunks, image tokens, structured data) to provide a richer grounding for generation Caffagni et al.
+(2024); Xue et al.
+(2024).
+EMERGE Zhu et al.
+(2024b) enriches context by integrating entity relationships and semantic descriptions.
 MiRAG Adjali et al.
-(2024) improves retrieval effectiveness by expanding initial queries through entity retrieval and query reformulation, enhancing subsequent stages for the visual question-answering task.
-Wiki-LLaVA Caffagni et al.
-(2024) incorporates text chunks and image tokens into multimodal inputs to enrich prompts for generation.
+(2024) expands initial queries through entity retrieval and reformulation, enhancing subsequent stages for the visual question-answering.
 Video-RAG Luo et al.
-(2024b) enhances long-video understanding through Query Decoupling, which reformulates user queries into structured retrieval requests to extract additional auxiliary multimodal context.
+(2024b) enhances long-video understanding through Query Decoupling, which reformulates user queries into structured retrieval requests to extract auxiliary multimodal context.
 Img2Loc Zhou et al.
 (2024e) enhances accuracy by including both the most similar and the most dissimilar points from the database in the prompt, allowing the model to rule out implausible locations for its predictions.
-Finally, Xue et al.
-(2024) introduce semantic-enhanced prompts, aligning structured image data with queries for contextual accuracy.
-Adaptive and Iterative Retrieval
+4.3.2 Adaptive and Iterative Retrieval
 For more complex queries, dynamic retrieval mechanisms have proven effective.
+Adaptive retrieval approaches optimize relevance by adjusting retrieval dynamically.
 SKURG Yang et al.
-(2023) employs adaptive retrieval steps to dynamically adjust the number of hops based on the query complexity.
+(2023) determines the number of retrieval hops based on query complexity.
+SAM-RAG Zhai (2024) and mR2AG Zhang et al.
+(2024f) dynamically assess the need for external knowledge and filter irrelevant content using MLLMs to retain only task-critical information.
+MMed-RAG Xia et al.
+(2024a) further improves retrieval precision by discarding low-relevance results, while OmniSearch Li et al.
+(2024b) introduces a self-adaptive retrieval agent that dynamically decomposes complex multimodal questions into sub-question chains and plans retrieval actions in real-time based on retrieved content.
+Iterative approaches refine results over multiple steps by incorporating feedback from prior iterations.
 IRAMIG Liu et al.
-(2024b) iteratively refines multimodal queries by incorporating feedback from previous iterations and context created by the fusion of retrieved items from text, images, and audio modalities.
+(2024b) improves multimodal retrieval by dynamically updating queries based on retrieved content.
 Similarly, OMG-QA Nan et al.
@@
 An evaluation module assesses retrieval effectiveness at each step, guiding the refinement of subsequent retrieval efforts through feedback.
-Taking a different approach, SAM-RAG Zhai (2024) further improves relevance by dynamically selecting the optimal number of retrieved contexts for each query.
-MMed-RAG Xia et al.
-(2024a) uses context filtering to discard low-relevance results.
-OmniSearch Li et al.
-(2024b) introduces a self-adaptive retrieval agent that dynamically decomposes complex multimodal questions into sub-question chains and plans retrieval actions in real-time based on retrieved content.
-mR2AG Zhang et al.
-(2024g) introduces reflection mechanisms to enhance retrieval effectiveness.
-Retrieval-Reflection dynamically determines when external knowledge is necessary, while Relevance-Reflection filters retrieved content, retaining only task-critical evidence.
-Similarly, RAGAR Khaliq et al.
-(2024) iteratively refines retrieved data by leveraging prior responses and multimodal analysis, ensuring contextual consistency.
+RAGAR Khaliq et al.
+(2024) further enhances contextual consistency by iteratively adjusting retrieval based on prior responses and multimodal analysis.
 4.4 Generation Techniques
 Recent advancements in multimodal RAG generation focus on robustness, cross-modal coherence, and task-specific adaptability. These innovations can be broadly categorized into these sections:
-In-Context Learning
+4.4.1 In-Context Learning
 In-context learning (ICL) with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining.
@@
 (2024) introduces Fusion-in-Context Learning, a novel approach that enriches ICL by incorporating a more diverse set of in-context examples, leading to better performance than standard ICL.
-Reasoning
+4.4.2 Reasoning
 Structured reasoning techniques, such as chain-of-thought (CoT), decompose complex reasoning into smaller sequential steps, enhancing coherence and robustness in multimodal RAG systems.
@@
 (2024) leverages LLMs for divergent compositional reasoning, generating refined captions by incorporating dense captions and modification text.
-Instruction Tuning
-Several works have fine-tuned generation components for specific applications.
+4.4.3 Instruction Tuning
+Several works have fine-tuned or instruct-tuned generation components for specific applications.
 RA-BLIP Ding et al.
-(2024b), uses Q-Former architecture from InstructBLIP Dai et al.
-(2023) to extract visual features based on question instructions.
-RAGPT Lang et al.
-(2025) uses a context-aware prompter, which captures contextual knowledge from relevant instances and generates dynamic prompts.
+(2024b) leverages the Q-Former architecture from InstructBLIP Dai et al.
+(2023) to extract visual features based on question instructions, while RAGPT Lang et al.
+(2025) employs a context-aware prompter to generate dynamic prompts from relevant instances.
 mR2AG Zhang et al.
-(2024g) uses instruction tuning with the mR2AG-IT dataset to train MLLMs to adaptively invoke retrieval, identify relevant evidence, and generate accurate answers for knowledge-based VQA tasks.
+(2024f) uses instruction tuning with the mR2AG-IT dataset to train MLLMs to adaptively invoke retrieval, identify relevant evidence, and generate accurate answers for knowledge-based VQA tasks.
 RagVL Chen et al.
@@
 Jang et al.
-(2024) propose a Visual Delta Generator to discriminate differences between image pairs and produce a textual response.
+(2024) focus on distinguishing image differences to generate descriptive textual responses.
 MMed-RAG Xia et al.
-(2024a) uses preference fine-tuning to help the model learn when to rely on retrieved information and when to trust its internal knowledge.
-Additionally, several methods have been proposed to improve generation quality by constructing datasets from the previous mistakes of LLMs and fine-tuning them accordingly.
-MegaPairs Zhou et al.
-(2024a) synthesizes large-scale multimodal instruction-tuning datasets for this purpose.
-Similarly, Surf Sun et al.
-(2024a) creates a dataset of positive and negative pairs derived from the prior errors of LVLMs and subsequently fine-tunes the LVLM using this dataset.
-Additionally, Rule Xia et al.
-(2024b) refines Med-LVLM by applying direct preference optimization on curated preference data, specifically focusing on instances where overreliance on retrieved contexts resulted in errors.
-Source Attribution and Evidence Transparency
+(2024a) applies preference fine-tuning to help models balance retrieved knowledge with internal reasoning.
+To improve generation quality, MegaPairs Zhou et al.
+(2024a) and Surf Sun et al.
+(2024a) construct multimodal instruction-tuning datasets from prior LLM errors, while Rule Xia et al.
+(2024b) refines Med-LVLM through direct preference optimization to mitigate overreliance on retrieved contexts.
+4.4.4 Source Attribution and Evidence Transparency
 Ensuring source attribution in multimodal RAG systems is a key focus of recent research.
 MuRAR Zhu et al.
-(2024d) integrates multimodal data, fetched by a source-based retriever, to refine LLM’s initial response, ensuring coherence and informativeness.
+(2025) integrates multimodal data, fetched by a source-based retriever, to refine LLM’s initial response, ensuring informativeness.
 VISA Ma et al.
-(2024b) uses large vision-language models (VLMs) to generate answers with visual source attribution by identifying and highlighting supporting evidence in retrieved document screenshots.
+(2024b) uses large vision-language models to generate answers with visual source attribution by identifying and highlighting supporting evidence in retrieved document screenshots.
 Similarly, OMG-QA Nan et al.
-(2024) enhances transparency by prompting the LLM to explicitly cite evidence in generated responses.
+(2024) ensures transparency by prompting the LLM to explicitly cite evidence in generated responses.
 4.5 Training Strategies
@@
 During fine-tuning, a cross-entropy objective trains the model for tasks like VQA and image captioning.
-Alignment
+Details of formulas for widely used RAG loss functions can be found in Appendix (§D).
+4.5.1 Alignment
 Contrastive learning improves representation quality by pulling positive pairs closer and pushing negative pairs apart in the embedding space.
@@
 It enforces sparsity through L1 regularization, optimizing retrieval precision by balancing dense and sparse representations.
-Details of formulas for widely used RAG loss functions can be found in Appendix D.
-Robustness
+4.5.2 Generation
+A key aspect of multimodal RAG is the generation ability.
+Autoregressive language models are typically trained using Cross-Entropy Loss Brown et al.
+(2020).
+For image generation, widely used approaches include Generative Adversarial Networks (GANs) Goodfellow et al.
+(2014) and Diffusion Models Ho et al.
+(2020).
+GANs employ various loss functions, such as Binary Cross-Entropy Loss, Minimax Loss, Wasserstein Loss (WGAN), and Hinge Loss.
+Diffusion Models utilize Mean Squared Error (MSE) Loss for noise prediction, a common approach in Denoising Diffusion Probabilistic Models (DDPMs) Ho et al.
+(2020).
+4.5.3 Robustness and Noise Management
 Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024).
@@
 (2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.
-{forest}
-forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2em, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0.4em, inner ysep=0.6em, line width=0.8pt, text width=8.5em, ver/.style= fill=gray!10, rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=8em , leaf/.style= text opacity=1, inner sep=2pt, fill opacity=.5, fill=hidden-blue!90, text=black, text width=44.5em font=, inner xsep=0.4em, inner ysep=0.6em, draw, , , [Multimodal RAG, ver [ Retrieval
-Strategies (§4.1) [ Efficient-Search and
-Similarity Retrieval
- (§4.1) [ Maximum Inner Product Search (MIPS) [ TPU-KNN Chern et al.
-(2022), ScaNN Guo et al.
-(2020), MAXSIM score Chan and Ng (2008), OmniSearch Li et al.
-(2024b), MuRAG Chen et al.
-(2022a), RetrievalAttention Liu et al.
-(2024a), RA-CM3 Yasunaga et al.
-(2023), RAVEN Rao et al.
-(2024) FactMM-RAG Sun et al.
-(2024b) , leaf, text width=44.5em ] ] [ Multimodal
-Encoders [ CLIP Radford et al.
-(2021), BLIP Li et al.
-(2022a), MARVEL Zhou et al.
-(2024c), ALIGN Jia et al.
-(2021), FLAVA Singh et al.
-(2022), UniVL-DR Liu et al.
-(2023b), UniIR Wei et al.
-(2025), GME Zhang et al.
-(2024j), VISTA Zhou et al.
-(2024b), ColPali Faysse et al.
-(2024), InternVideo Wang et al.
-(2022), Ovis Lu et al.
-(2024) Mi-RAG Adjali et al.
-(2024) , leaf, text width=44.5em ] ] ] [ Modality-Centric
-Retrieval (§4.1) [ Text-Centric [ Contriever Izacard et al.
-(2022), GTE Li et al.
-(2023c), Re-Imagen Chen et al.
-(2022b), BM25 Robertson and Zaragoza (2009), BGE-1.5 Xiao et al.
-(2023), BGE-M3 Chen et al.
-(2024a), CapRet Shohan et al.
-(2024), OMG-QA Nan et al.
-(2024), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al.
-(2024b), RAFT Zhang et al.
-(2024i), CRAG Yan et al.
-(2024), M2-RAG Ma et al.
-(2024c) , leaf, text width=44.5em ] ] [ Vision-Centric [ VQA4CIR Feng et al.
-(2023), Unifashion Zhao et al.
-(2024), Jang et al.
-Jang et al.
-(2024), Pic2word Saito et al.
-(2023), eClip Kumar and Marttinen (2024), RAMM Yuan et al.
-(2023), Joshi et al.
-Joshi et al.
-(2024), VISA Ma et al.
-(2024b), ImgRet Shohan et al.
-(2024), EchoSight Yan and Xie (2024), Xue et al.
-Xue et al.
-(2024) , leaf, text width=44.5em ] ] [ Video-Centric [ iRAG Arefeen et al.
-(2024), VideoRAG Ren et al.
-(2025), VideoRAG Jeong et al.
-(2025), MV-Adapter Jin et al.
-(2024), OmAgent Zhang et al.
-(2024f), CM2 Kim et al.
-(2024), Video-RAG Luo et al.
-(2024b), CTCH Shen et al.
-(2024), RTime Du et al.
-(2024), VideoMAE Tong et al.
-(2022), DrVideo Ma et al.
-(2024d) , leaf, text width=44.5em ] ] [ Document Retrieval [ ColPali Faysse et al.
-(2024), ColQwen2 Wang et al.
-(2024b), M3DocVQA Cho et al.
-(2024), ViTLP Mao et al.
-(2024), DocLLM Wang et al.
-(2024a), CREAM Zhang et al.
-(2024b), mPLUG-DocOwl 1.5 Hu et al.
-(2024a), mPLUG-DocOwl 2 Hu et al.
-(2024b), VisDom Suri et al.
-(2024), DSE Ma et al.
-(2024a) , leaf, text width=44.5em ] ] ] [ Re-ranking
-Strategies (§4.1) [ Optimized Example
-Selection [ MSIER Luo et al. (2024a), Hybrid RAG Su et al. (2024b), RULE Xia et al. (2024b), RAMM Yuan et al. (2023), M2RAAP Dong et al. (2024b) , leaf, text width=44.5em ] ] [ Relevance Score
-Evaluation [ RAG-Check Mortaheb et al.
-(2025a, b), UniRaG Zhi Lim et al.
-(2024), MR2AG Zhang et al.
-(2024h), LDRE Yang et al.
-(2024), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al.
-(2024), OMG-QA Nan et al.
-(2024), EchoSight Yan and Xie (2024), EgoInstructor Xu et al.
-(2024a) , leaf, text width=44.5em ] ] [ Filtering
-Mechanisms [ MAIN-RAG Chang et al.
-(2024), MM-Embed Lin et al.
-(2024a), GME Zhang et al.
-(2024j), Img2Loc Zhou (2024), MuRAR Zhu et al.
-(2024d) RAFT Zhang et al.
-(2024i) , leaf, text width=44.5em ] ] ] ] [ Fusion
-Mechanisms (§4.2) [ Score Fusion
-and Alignment
- (§4.2) [ M3 Cai et al.
-(2025)  Zhi Lim et al.
-(2024),  Sharifymoghaddam et al.
-(2024), REVEAL Hu et al.
-(2023), RAG-Driver Yuan et al.
-(2024), LLM-RA Jian et al.
-(2024),  Riedler and Langer (2024), VISA Ma et al.
-(2024b), MA-LMM He et al.
-(2024),  Xue et al.
-(2024), RA-BLIP Ding et al.
-(2024b), Re-IMAGEN Chen et al.
-(2022b), MegaPairs Zhou et al.
-(2024a), Wiki-LLaVA Caffagni et al.
-(2024), VISRAG Yu et al.
-(2024) , leaf, text width=55.0em ] ] [ Attention-Based
-Mechanisms (§4.2) [ RAMM Yuan et al.
-(2023), EMERGE Zhu et al.
-(2024b), MORE Cui et al.
-(2024), RAGTrans Cheng et al.
-(2024), AlzheimerRAG Lahiri and Hu (2024), MV-Adapter Jin et al.
-(2024), C3Net Zhang et al.
-(2024c),  Xu et al.
-(2024a),  Kim et al.
-(2024), M2-RAAP Dong et al.
-(2024b), Mu-RAG Chen et al.
-(2022a), Ou et al.
-Ou et al.
-(2025), CADMR Khalafaoui et al.
-(2024) , leaf, text width=55.0em ] ] [ Unified Frameworks
-and Projections
- (§4.2) [ Hybrid-RAG Su et al.
-(2024b), Dense2Sparse Nguyen et al.
-(2024), IRAMIG Liu et al.
-(2024b), M3DocRAG Cho et al.
-(2024), DQU-CIR Wen et al.
-(2024), PDF-MVQA Ding et al.
-(2024c), SAM-RAG Zhai (2024), UFineBench Zuo et al.
-(2024),  Li et al.
-(2022a), , leaf, text width=55.0em ] ] ] [ Augmentation
-Techniques (§4.3) [ Context-Enrichment (§4.3) [ EMERGE Zhu et al.
-(2024b), MiRAG Adjali et al.
-(2024), Wiki-LLaVA Caffagni et al.
-(2024), Video-RAG Luo et al.
-(2024b), Img2Loc Zhou et al.
-(2024e),  Xue et al.
-(2024) , leaf, text width=55.0em ] ] [ Adaptive and
-Iterative
-Retrieval (§4.3) [ SKURG Yang et al.
-(2023), IRAMIG Liu et al.
-(2024b), OMG-QA Nan et al.
-(2024), SAM-RAG Zhai (2024), MMed-RAG Xia et al.
-(2024a), OmniSearch Li et al.
-(2024b), mR2AG Zhang et al.
-(2024g), RAGAR Khaliq et al.
-(2024) , leaf, text width=55.0em ] ] ] [ Generation
-Techniques (§4.4) [ In-Context
-Learning (§4.4) [ RMR Tan et al.
-(2024),  Sharifymoghaddam et al.
-(2024), RA-CM3 Yasunaga et al.
-(2023), RAG-Driver Yuan et al.
-(2024), MSIER Luo et al.
-(2024a), Raven Rao et al.
-(2024) , leaf, text width=55.0em ] ] [ Reasoning (§4.4) [ RAGAR Khaliq et al.
-(2024), VisDoM Suri et al.
-(2024), SAM-RAG Zhai (2024), LDRE Yang et al.
-(2024) , leaf, text width=55.0em ] ] [ Instruction
-Tuning (§4.4) [ RA-BLIP Ding et al.
-(2024b), RAGPT Lang et al.
-(2025), mR2AG Zhang et al.
-(2024g), RagVL Chen et al.
-(2024d), Jang et al.
-(2024), MMed-RAG Xia et al.
-(2024a), MegaPairs Zhou et al.
-(2024a), Surf Sun et al.
-(2024a), Rule Xia et al.
-(2024b) , leaf, text width=55.0em ] ] [ Source
-Attribution (§4.4) [ MuRAR Zhu et al.
-(2024d), VISA Ma et al.
-(2024b), OMG-QA Nan et al.
-(2024) , leaf, text width=55.0em ] ] ] [ Training Strategies (§4.5) [ Alignment (§4.5) [ VISRAG Yu et al.
-(2024), MegaPairs Zhou et al.
-(2024a), SAM-RAG Zhai (2024), EchoSight Yan and Xie (2024), HACL Jiang et al.
-(2024),  Zhi Lim et al.
-(2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al.
-(2024) , leaf, text width=55.0em ] ] [ Robustness (§4.5) [  Buettner and Kovashka (2024), MORE Cui et al.
-(2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al.
-(2024), RA-BLIP Ding et al.
-(2024b), RagVL Chen et al.
-(2024d), RA-CM3 Yasunaga et al.
-(2023) , leaf, text width=55.0em ] ] ] ] ]
-Figure 2: Taxonomy of recent advances and enhancements in multimodal retrieval-augmented generation research.
 5 Tasks Addressed by Multimodal RAGs
@@
 Despite rapid advancements in multimodal RAG systems, fundamental challenges remain in achieving robust, efficient, and human-like reasoning across modalities.
-Generalization, Explainability, and Robustness
-Multimodal RAG systems often struggle with domain-specific training and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al.
+6.1 Generalization, Explainability, and Robustness
+Multimodal RAG systems often struggle with domain adaptation and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al.
 (2020).
-Explainability remains a key challenge, as these systems typically fail to attribute answers to precise sources.
+Explainability remains a major challenge, as these systems typically fail to attribute answers to precise sources.
 Current methods often cite entire documents or large visual regions as source attribution rather than identifying the exact part of an image, speech, or other modality that led to the answer Ma et al.
@@
 (2024d), enhancing the robustness of multimodal RAGs remains an open challenge and a promising research direction.
-Reasoning, Alignment, and Retrieval Enhancement
+6.2 Reasoning, Alignment, and Retrieval Enhancement
 Multimodal RAGs struggle with compositional reasoning, where information from different modalities must be logically integrated to generate coherent and contextually rich outputs.
-While cross-modal techniques such as Multimodal-CoT have been proposed Zhang et al.
-(2023), further innovations are needed to improve the coherence and contextual relevance of multimodal outputs.
+While cross-modal techniques such as Multimodal-CoT Zhang et al.
+(2023b) have been proposed, further innovations are needed to improve the coherence and contextual relevance of multimodal outputs.
 Enhancing modality alignment and retrieval strategies, particularly for entity-aware retrieval, is essential.
 Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain largely underexplored in multimodal RAGs compared to text-based RAGs Zhang et al.
-(2024g); Procko and Ochoa (2024).
+(2024f); Procko and Ochoa (2024).
 Retrieval biases such as position sensitivity Hu et al.
@@
 Despite some progress, mapping multimodal knowledge into a unified space remains an open challenge with significant potential.
-Agent-Based and Self-Guided Systems
+6.3 Agent-Based and Self-Guided Systems
 Recent trends indicate a shift towards agent-based multimodal RAGs that integrate retrieval, reasoning, and generation across diverse domains.
 Unlike static RAG systems, future multimodal RAGs should integrate interactive feedback and self-guided decision-making to iteratively refine outputs.
-Furthermore, existing feedback mechanisms often fail to accurately identify whether issues with responses arise during retrieval, generation, or other stages Dong et al.
+Existing feedback mechanisms often fail to accurately determine whether errors stem from retrieval, generation, or other stages Dong et al.
 (2024b).
@@
 Bridging retrieval-based reasoning with real-world agency brings these systems closer to AGI.
-Long-Context Processing, Scalability, and Personalization
+6.4 Long-Context Processing, Efficiency, Scalability, and Personalization
 High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing.
-Fixed extraction rates often fail to capture relevant video frames, necessitating dynamic adjustment based on content complexity and movement Kandhare and Gisselbrecht (2024).
-Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers highlight the need for more scalable architectures.
+Fixed extraction rates struggle to capture relevant frames, requiring adaptive selection based on content complexity and movement Kandhare and Gisselbrecht (2024).
+Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers emphasize the need for efficient, scalable architectures.
 Personalization mechanisms, such as adapting retrieval to user-specific contexts like medical history, remain underexplored.
 As personalization mechanisms evolve, ensuring privacy and mitigating the risks of sensitive data leakage in multimodal outputs remain critical challenges.
-Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples hinders robust evaluation.
+Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples limits robust evaluation.
 7 Conclusion
-In this study, we take a deep dive into the realm of multimodal Retrieval-Augmented Generation (RAG) literature.
+This study provides a comprehensive review of multimodal Retrieval-Augmented Generation (Multimodal RAG) literature.
 Specifically, we explore and categorize key advancements across different aspects of multimodal RAG systems, including retrieval, multimodal fusion, augmentation, generation, and training strategies.
 Additionally, we examine the tasks these systems address, their domain-specific applications, and the datasets, benchmarks, and evaluation methods.
-We also discuss open challenges and limitations in current approaches.
-We hope it could encourage future research in the domain, particularly in improving cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
+We also discuss open challenges and limitations in current approaches, along with promising future directions.
+We hope this work inspires future research, particularly in enhancing cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
 8 Limitations
@@
 As new paradigms emerge, our taxonomy and conclusions will inevitably evolve.
-To address these gaps, we plan to continuously monitor developments and update this survey and the corresponding GitHub repository to incorporate overlooked contributions and refine our perspectives.
+To address these gaps, we plan to continuously monitor developments and update this survey and the corresponding repository to incorporate overlooked contributions and refine our perspectives.
+9 Ethical Statement
+This survey provides a comprehensive review of research on multimodal RAG systems, offering insights that we believe will be valuable to researchers in this evolving field.
+All the studies, datasets, and benchmarks analyzed in this work are publicly available, with only a very small number of papers requiring institutional access.
+Additionally, this survey does not involve personal data or user interactions, and we adhere to ethical guidelines throughout.
+Since this work is purely a survey of existing literature and does not introduce new models, datasets, or experimental methodologies, it presents no potential risks.
+However, we acknowledge that multimodal RAG systems inherently raise ethical concerns, including bias, misinformation, privacy, and intellectual property issues.
+Bias can emerge from both retrieval and generation processes, potentially leading to skewed or unfair outputs.
+Additionally, these models may hallucinate or propagate misinformation, particularly when retrieval mechanisms fail or rely on unreliable sources.
+The handling of sensitive multimodal data also poses privacy risks, while content generation raises concerns about proper attribution and copyright compliance.
+Addressing these challenges requires careful dataset curation, bias mitigation strategies, and transparent evaluation of retrieval and generation mechanisms.
 Appendix A Taxonomy
-In this section, we present a comprehensive taxonomy of multimodal retrieval-augmented generation (RAG) systems, along with a classification of their application domains.
-Figures 2 and 3 summarize these taxonomies, respectively.
+In this section, we provide more details regarding the taxonomy of multimodal RAG systems, previously mentioned in Figure 2.
+Additionally, we present a classification of multimodal RAG application domains in Figure 3.
 Figure 2 provides an overview of recent advances in multimodal retrieval-augmented generation (RAG) systems. The taxonomy is organized into several key categories.
@@
 Detailed discussions of these categories are provided in the corresponding sections.
-Figure 3
-presents the taxonomy of application domains for multimodal RAG systems.
+Figure 3 presents the taxonomy of application domains for multimodal RAG systems.
 The identified domains include healthcare and medicine, software engineering, fashion and e-commerce, entertainment and social computing, and emerging applications.
@@
 The datasets are grouped into eight categories: Image–Text General, Video–Text, Audio–Text, Medical, Fashion, 3D, Knowledge & QA, and Other.
-The benchmarks are grouped into three categories: Vision-Focused, Cross-Modal and Reasoning-Focused, and Text and Knowledge-Focused.
+The benchmarks are grouped into two categories: Cross-Modal Understanding and Text-Focused.
 This classification facilitates a clearer understanding of each dataset or benchmark’s role within a multimodal framework.
@@
 Audio-Text LibriSpeech Panayotov et al. (2015) 1,000 hours of read English speech with corresponding text; ASR corpus based on audiobooks. Audio, Text LibriSpeech
+SpeechBrown Abootorabi and Asgari (2024) 55K paired speech-text samples; 15 categories covering diverse topics from religion to fiction. Audio, Text SpeechBrown
 AudioCap Kim et al. (2019) 46K audio clips paired with human-written text captions. Audio, Text AudioCaps
@@
 VQA v2 Goyal et al. (2017b) 1.1M questions (augmented with VG–QA questions) for fine-tuning VQA models. Image, Text VQA v2
-ScienceQA Lu et al. (2022) High–school science curriculum dataset with 12K training samples and over 21K questions across various subjects. Text ScienceQA
 A-OKVQA Schwenk et al. (2022) Benchmark for visual question answering using world knowledge; around 25K questions. Image, Text A-OKVQA
-MMBench Liu et al. (2025) Bilingual benchmark for assessing the multimodal capabilities of VLMs; over 3,000 multiple–choice questions covering 20 dimensions. Text MMBench
 XL-HeadTags Shohan et al. (2024) 415K news headline-article pairs consist of 20 languages across six diverse language families. Text XL-HeadTags
@@
 Table 2: Overview of Popular Benchmarks in Multimodal RAG Research.
-Category Name Statistics and Description Modalities Link
-Vision-Focused MRAG-Bench Hu et al. (2024c)
-Evaluates visual information retrieval and integration;
-tests robustness to irrelevant visual information and visual reasoning.
-Images MRAG-Bench
-VQAv2 Goyal et al. (2017a)
-Contains 265,016 images with 1,105,904 questions; assesses the ability to answer questions about images,
-integrating image and question text, visual reasoning, and common-sense reasoning.
-Images + Text VQAv2
-OK-VQA Marino et al. (2019)
-Comprises 14,031 questions; tests retrieval and integration of external information,
-combining image, question, and external knowledge, focusing on reasoning with external knowledge.
-Images + Text OK-VQA
-Cross-Modal and Reasoning-Focused M2RAG Ma et al. (2024c)
-Establishes a comprehensive benchmark for multimodal retrieval-augmented generation;
-integrates fine-grained text-modal and multimodal metrics to evaluate foundation models on dynamic retrieval,
-multi-hop reasoning, and multimodal integration.
-Images + Text M2RAG
-Dyn-VQA Li et al. (2024b)
-Emphasizes dynamic retrieval with rapidly changing information;
-requires integration of dynamic knowledge retrieval and multi-hop reasoning, testing robustness to changing information.
-Images + Text Dyn-VQA
-MMBench Liu et al. (2025)
-Covers various tasks, including visual question answering, image captioning, and cross-modal retrieval;
-focuses on integrating information across multiple modalities (vision, text, audio) for general multimodal understanding.
-Images + Text + Audio MMBench
-ScienceQA Saikh et al. (2022)
-Contains 21,208 questions; tests scientific reasoning abilities using multimodal contexts (text, diagrams, images);
-requires deep understanding and integration of information, scientific reasoning, and diagram understanding.
-Images + Diagrams + Text ScienceQA
-SK-VQA Su et al. (2024c)
-Offers over 2 million question-answer pairs;
-focuses on synthetic knowledge generation, multimodal reasoning, and external knowledge integration;
-used for training context-augmented multimodal large language models.
-Images + Text SK-VQA
-Text and Knowledge-Focused TriviaQA Joshi et al. (2017)
-Provides 650K question-answer pairs; a reading comprehension dataset where questions are based on evidence documents;
-can be adapted to multimodal RAG by providing images or other visual context related to the trivia;
-focuses on knowledge retrieval and factual accuracy.
-Text TriviaQA
-Natural Questions Kwiatkowski et al. (2019)
-Contains 307,373 training examples; questions are naturally occurring Google search queries;
-adapting this for multimodal RAG involves finding or creating relevant visual contexts for the queries;
-emphasizes knowledge retrieval and handling noisy or ambiguous queries.
-Text Natural Questions
-SMMQG Wu et al. (2024a)
-Includes 1,024 question-answer pairs; focuses on synthetic data generation,
-multimodal retrieval-augmented generation, and control over question style and modality.
-Images + Text SMMQG
+Category
+Name Statistics and Description Modalities Link
+Cross-Modal Understanding MRAG-Bench Hu et al. (2024c) Evaluates visual retrieval, integration, and robustness to irrelevant visual information. Images MRAG-Bench
+M2RAG Ma et al. (2024c) Benchmarks multimodal RAG; evaluates retrieval, multi-hop reasoning, and integration. Images + Text M2RAG
+Dyn-VQA Li et al. (2024b) Focuses on dynamic retrieval, multi-hop reasoning, and robustness to changing information. Images + Text Dyn-VQA
+MMBench Liu et al. (2025) Covers VQA, captioning, retrieval; evaluates cross-modal understanding across vision, text, and audio. Images + Text + Audio MMBench
+ScienceQA Saikh et al. (2022) Contains 21,208 questions; tests scientific reasoning with text, diagrams, and images. Images + Diagrams + Text ScienceQA
+SK-VQA Su et al. (2024b) Offers 2 million question-answer pairs; focuses on synthetic knowledge, multimodal reasoning, and external knowledge integration. Images + Text SK-VQA
+SMMQG Wu et al. (2024a) Includes 1,024 question-answer pairs; focuses on synthetic multimodal data and controlled question generation. Images + Text SMMQG
+Text-Focused TriviaQA Joshi et al. (2017) Provides 650K question-answer pairs; reading comprehension dataset, adaptable for multimodal RAG. Text TriviaQA
+Natural Questions Kwiatkowski et al. (2019) Contains 307,373 training examples; real-world search queries, adaptable with visual contexts. Text Natural Questions
 Appendix C Metrics
@@
 The Generator loss formula is:
-$\mathcal{L}_{G}=-\mathbb{E}_{z\sim p_{z}(z)}[\log D(G(z))]$ (16)
+$\mathcal{L}_{G}=\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (14)
 Triplet Loss: Triplet Loss is used in metric learning to ensure that similar data points are closer together while dissimilar ones are farther apart in the embedding space.

### v3

@@
-Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation
+A Comprehensive Survey on Multimodal Retrieval-Augmented Generation
 Abstract
-Large Language Models (LLMs) struggle with hallucinations and outdated knowledge due to their reliance on static training data.
-Retrieval-Augmented Generation (RAG) mitigates these issues by integrating external dynamic information enhancing factual and updated grounding.
-Recent advances in multimodal learning have led to the development of Multimodal RAG, incorporating multiple modalities such as text, images, audio, and video to enhance the generated outputs.
-However, cross-modal alignment and reasoning introduce unique challenges to Multimodal RAG, distinguishing it from traditional unimodal RAG.
-This survey offers a structured and comprehensive analysis of Multimodal RAG systems, covering datasets, metrics, benchmarks, evaluation, methodologies, and innovations in retrieval, fusion, augmentation, and generation.
-We precisely review training strategies, robustness enhancements, and loss functions, while also exploring the diverse Multimodal RAG scenarios.
-Furthermore, we discuss open challenges and future research directions to support advancements in this evolving field.
+Large Language Models (LLMs) suffer from hallucinations and outdated knowledge due to their reliance on static training data.
+Retrieval-Augmented Generation (RAG) mitigates these issues by integrating external dynamic information for improved factual grounding.
+With advances in multimodal learning, Multimodal RAG extends this approach by incorporating multiple modalities such as text, images, audio, and video to enhance the generated outputs.
+However, cross-modal alignment and reasoning introduce unique challenges beyond those in unimodal RAG.
+This survey offers a structured and comprehensive analysis of Multimodal RAG systems, covering datasets, benchmarks, metrics, evaluation, methodologies, and innovations in retrieval, fusion, augmentation, and generation.
+We review training strategies, robustness enhancements, loss functions, and agent-based approaches, while also exploring the diverse Multimodal RAG scenarios.
+In addition, we outline open challenges and future directions to guide research in this evolving field.
 This survey lays the foundation for developing more capable and reliable AI systems that effectively leverage multimodal dynamic external knowledge bases.
-Resources are available at https://github.com/llm-lab-org/Multimodal-RAG-Survey.
-Ask in Any Modality:
+All resources are publicly available  [^1] https://github.com/llm-lab-org/Multimodal-RAG-Survey.
+spacing=nonfrench
+Ask in Any Modality
 A Comprehensive Survey on Multimodal Retrieval-Augmented Generation
@@
 1 Introduction & Background
-In recent years, significant breakthroughs have been achieved in language models, driven primarily by the advent of transformers Vaswani et al.
-(2017), enhanced computational capabilities, and the availability of large-scale training data Naveed et al.
-(2024).
+Recent advancements in transformer architectures Vaswani et al.
+(2017), coupled with increased computational resources and the availability of large-scale training datasets Naveed et al.
+(2024), have significantly accelerated progress in the development of language models.
 The emergence of foundational Large Language Models (LLMs) Ouyang et al.
@@
 (2025); Anil et al.
-(2023) has revolutionized natural language processing (NLP), demonstrating unprecedented capabilities in a wide range of tasks including instruction following Qin et al.
-(2024), sophisticated reasoning Wei et al.
-(2024c), In-context Learning Brown et al.
-(2020), and multilingual machine translation Zhu et al.
+(2023), has revolutionized natural language processing (NLP), excelling in tasks such as instruction following Qin et al.
+(2024), reasoning Wei et al.
+(2024b), in-context learning Brown et al.
+(2020), and multilingual translation Zhu et al.
 (2024a).
-These advancements have elevated the performance of various NLP tasks, opening new avenues for research and application.
-Despite their remarkable achievements, LLMs face significant challenges, including hallucination, outdated internal knowledge, and a lack of verifiable reasoning Huang et al.
+Despite these achievements, LLMs face challenges such as hallucinations, outdated knowledge, and a lack of verifiable reasoning Huang et al.
 (2024); Xu et al.
 (2024b).
-Their reliance on parametric memory restricts their ability to access up-to-date knowledge, making them less effective for knowledge-intensive tasks compared to task-specific architectures.
-Moreover, providing provenance for their decisions and updating their world knowledge remain critical open problems Lewis et al.
-(2020).
+Their reliance on parametric memory limits access to up-to-date information, reducing their effectiveness in knowledge-intensive tasks.
 Retrieval-Augmented Generation (RAG)
 RAG Lewis et al.
-(2020) has emerged as a promising solution to these limitations by enabling LLMs to retrieve and incorporate external knowledge, improving factual accuracy and reducing hallucinations Shuster et al.
+(2020) addresses these limitations by enabling LLMs to retrieve and incorporate external knowledge, improving factual accuracy and reducing hallucinations Shuster et al.
 (2021); Ding et al.
 (2024a).
-By dynamically accessing vast external knowledge repositories, RAG systems enhance knowledge-intensive tasks while ensuring responses remain grounded in verifiable sources Gao et al.
+By dynamically accessing external knowledge sources, RAG enhances knowledge-intensive tasks while grounding responses in verifiable sources Gao et al.
 (2023).
-In practice, RAG systems operate through a retriever-generator pipeline.
-The retriever leverages embedding models Chen et al.
+In practice, RAG systems follow a retriever-generator pipeline: the retriever uses embedding models Chen et al.
 (2024a); Rau et al.
-(2024) to identify relevant passages from external knowledge bases and optionally applies re-ranking techniques to improve retrieval precision Dong et al.
+(2024) to identify relevant passages from external knowledge bases and may apply re-ranking techniques to improve precision Dong et al.
 (2024a).
-These retrieved passages are then passed to the generator, incorporating this external context to produce informed responses.
+The retrieved passages are then provided to the generator, which leverages this contextual information to produce more informed and coherent responses.
 Recent advancements in RAG frameworks, such as planning-guided retrieval Lee et al.
@@
 (2024c); Asai et al.
-(2023), further enhance both retrieval and generation stages.
-However, traditional RAG architectures are primarily designed for textual information, limiting their ability to address multimodal challenges that require integrating diverse data formats.
+(2023), have further improved both the retrieval and generation components of these systems.
+Figure 1: Overview of the multimodal RAG pipeline, illustrating key techniques and recent advancements.
 Multimodal Learning
-Parallel to these developments, significant advances in multimodal learning have reshaped artificial intelligence by enabling systems to integrate and analyze heterogeneous data sources for a holistic representation of information.
-The introduction of CLIP (Contrastive Language-Image Pretraining) Radford et al.
-(2021) marked a pivotal moment in connecting visual and textual information through contrastive learning, inspiring numerous subsequent models and applications Alayrac et al.
+In parallel with advances in language modeling, multimodal learning has emerged as a transformative area in artificial intelligence, enabling systems to integrate and reason over heterogeneous data sources for more comprehensive representations.
+A pivotal breakthrough was the introduction of CLIP model Radford et al.
+(2021), which aligned visual and textual modalities through contrastive learning and inspired a wave of subsequent models Alayrac et al.
 (2024); Wang et al.
-(2023); Pramanick et al.
+(2023a); Pramanick et al.
 (2023).
-These breakthroughs have driven progress in various domains, including sentiment analysis Das and Singh (2023) and cutting-edge biomedical research Hemker et al.
-(2024), demonstrating the value of multimodal approaches.
-By enabling systems to process and understand diverse data types such as text, images, audio, and video, multimodal learning plays a key role in advancing artificial general intelligence (AGI) Song et al.
+These developments have catalyzed progress across diverse domains, including sentiment analysis Das and Singh (2023) and biomedical applications Hemker et al.
+(2024), highlighting the effectiveness of multimodal approaches.
+By facilitating the joint processing of text, images, audio, and video, multimodal learning is increasingly recognized as a critical enabler of progress toward artificial general intelligence (AGI) Song et al.
 (2025).
-Figure 1: Overview of the multimodal retrieval-augmented generation (RAG) pipeline, highlighting the advancements and techniques employed at each stage.
-The flow begins with query preprocessing, where user queries are refined and then encoded into a shared embedding space alongside a multimodal database.
-Retrieval strategies, such as modality-centric retrieval, similarity search, and re-ranking, enhance document selection, while fusion mechanisms align and integrate data from multiple modalities using score fusion or attention-based methods.
-Augmentation techniques such as iterative retrieval with feedback mechanisms, further refine the retrieved documents for the multimodal LLM.
-The generation stage incorporates innovations like Chain-of-Thought reasoning and source attribution for better outputs, with loss functions combining alignment loss and generation loss to optimize both retrieval and generation components.
-Noise management techniques are also applied to improve training stability and robustness.
 Multimodal RAG
-The extension of LLMs to multimodal LLMs (MLLMs) has further expanded their capabilities, allowing them to process, reason, and generate outputs across diverse modalities Liu et al.
+The extension of large language models to multimodal LLMs (MLLMs) has significantly broadened their capabilities, enabling reasoning and generation across multiple data modalities Liu et al.
 (2023a); Team et al.
@@
 (2023b).
-For example, GPT-4 OpenAI et al.
-(2024) demonstrates human-level performance in various benchmarks by processing both text and images, marking a significant milestone in multimodal perception and interaction.
-Building on this foundation, multimodal RAG systems extend traditional RAG frameworks by incorporating multimodal knowledge sources, such as images and audio, to provide enriched context for generation Hu et al.
+Notably, GPT-4 OpenAI et al.
+(2024) demonstrates human-level performance by jointly processing text and images, marking a milestone in multimodal understanding.
+Building on this progress, multimodal RAG incorporates diverse sources, such as images, audio, and structured data, to enrich contextual grounding and enhance generation quality Hu et al.
 (2023); Chen et al.
 (2022a).
-This integration enhances the precision of generated outputs while leveraging multimodal cues to improve the reasoning capabilities of MLLMs.
-However, these multimodal systems also present unique challenges, including determining which modalities to retrieve, effectively fusing diverse data types, and addressing the complexities of cross-modal relevance Zhao et al.
+This approach improves reasoning by leveraging cross-modal cues, but also introduces challenges, including modality selection, effective fusion, and managing cross-modal relevance Zhao et al.
 (2023a).
 Figure 1 illustrates the general pipeline of these systems.
-Task Formulation
-A mathematical formulation of the general task for multimodal RAG is presented in this section. These systems generate a multimodal response, denoted as $r$, in response to a multimodal query $q$.
-Let $D=\{d_{1},d_{2},...,d_{n}\}$ be a multimodal corpus. Each document $d_{i}\in D$ is associated with a modality $M_{d_{i}}$ and processed by a modality-specific encoder, yielding:
-$z_{i}=Enc_{M_{d_{i}}}(d_{i})$ (1)
-The set of all encoded representations is denoted by $Z=\{z_{1},z_{2},...,z_{n}\}$.
-Modality-specific encoders map different modalities into a shared semantic space for cross-modal alignment.
-A retrieval model $R$ assesses the relevance of each encoded document representation $z$ with respect to the query $q$, represented as $R(q,z)$.
-To construct the retrieval-augmented multimodal context, the retrieval model selects the most relevant documents based on a modality-specific threshold:
-$X=\{d_{i}\mid s(e_{q},z_{i})\geq\tau_{M_{d_{i}}}\}$ (2)
-where $\tau_{M_{d_{i}}}$ is a relevancy threshold for the modality of $M_{d_{i}}$, $e_{q}$ is the encoded representation of $q$ in the shared semantic space, and $s$ is a scoring function that measures the relevance between the encoded query and document representations.
-The generative model $G$ produces the final multimodal response, given the user query $q$ and the retrieved documents $X$ as context, denoted as $r=G(q,X)$.
+Multimodal RAG Formulation
+We present a mathematical formulation of multimodal RAG.
+These systems aim to generate a multimodal response $r$ given a multimodal query $q$.
+Let $D=\{d_{1},d_{2},\ldots,d_{n}\}$ denote a multimodal corpus.
+For clarity, we assume each document $d_{i}\in D$ is associated with a single modality $M_{d_{i}}$.
+In practice, however, documents often span multiple modalities—for example, a scientific article containing both text and images.
+Such cases are typically addressed by either decomposing the document into modality-specific sub-documents or employing universal encoders capable of jointly processing multiple modalities.
+Each document $d_{i}$ is encoded using its corresponding modality-specific encoder, yielding $z_{i}=\text{Enc}_{M_{d_{i}}}(d_{i})$.
+The collection of all encoded representations is denoted as $Z=\{z_{1},z_{2},\ldots,z_{n}\}$.
+These modality-specific encoders project diverse input modalities into a shared semantic space, enabling cross-modal alignment.
+A retrieval model $R$ computes a relevance score $s(e_{q},z_{i})$ between the encoded query representation $e_{q}$ (obtained by encoding $q$ using the appropriate encoders) and each document representation $z_{i}$.
+The retrieval-augmented multimodal context $X$ is constructed by selecting documents whose relevance scores exceed a modality-specific threshold:
+$X=\{d_{i}\in D\mid s(e_{q},z_{i})\geq\tau_{M_{d_{i}}}\},\vspace{-0.1cm}$
+where $\tau_{M_{d_{i}}}$ is the relevance threshold for the modality $M_{d_{i}}$, and $s$ is the scoring function that measures semantic relevance.
+Finally, the generative model $G$ produces the response conditioned on the original query $q$ and the retrieved context $X$, formally defined as $r=G(q,X)$.
 Related Works
-As the field of multimodal RAGs is newly introduced and evolving rapidly, especially in recent years, there is a pressing need for a comprehensive survey that explores the current innovations and frontiers of these systems.
-While more than ten surveys have been published on RAG-related topics such as Agentic RAG Singh et al.
-(2025), none provide a detailed and comprehensive overview of advancements in multimodal RAGs.
-The only related survey to date Zhao et al.
-(2023a) categorizes multimodal RAGs by grouping relevant papers based on their applications and modalities.
-However, our survey provides a more detailed and innovation-driven perspective, offering a detailed taxonomy and exploring emerging trends and challenges in depth.
-Furthermore, significant advancements have been made in the field since its publication, and interest in this topic has grown substantially within the research community.
-In this survey, we review over 100 papers on multimodal RAGs published in recent years, primarily from the ACL Anthology and other repositories such as the ACM Digital Library.
+Multimodal RAG is a rapidly emerging field, yet a comprehensive survey dedicated to its recent advancements remains lacking.
+While over ten surveys discuss RAG-related topics such as Agentic RAG Singh et al.
+(2025), none specifically focus on the multimodal setting.
+To our knowledge, the only relevant work Zhao et al.
+(2023a) categorizes multimodal RAGs by application and modality.
+In contrast, our survey adopts a more innovation-driven perspective, offering a detailed taxonomy and addressing recent trends and open challenges.
+We review over 100 recent papers, primarily from the ACL Anthology, reflecting the growing interest and progress in this domain.
 Contributions
-In this work, (i) we provide a comprehensive review of multimodal RAG, covering task formulation, datasets, benchmarks, tasks and domain-specific applications, evaluation, and key innovations in retrieval, fusion, augmentation, generation, training strategies, and loss functions.
-(ii) We introduce a precise structured taxonomy (Figure 2) that categorizes state-of-the-art models based on their primary contributions, highlighting methodological advancements and emerging frontiers.
-(iii) To support further research, we make resources, including datasets, benchmarks, and key innovations, publicly available.
-(iv) We identify current research trends and knowledge gaps, providing insights and recommendations to guide future advancements in this evolving field.
-2 Datasets and Benchmarks
-Multimodal RAG research employs diverse datasets and benchmarks to evaluate retrieval, integration, and generation across heterogeneous sources.
-Image–text tasks, including captioning and retrieval, commonly use MS-COCO Lin et al.
-(2014), Flickr30K Young et al.
-(2014), and LAION-400M Schuhmann et al.
-(2021), while visual question answering with external knowledge is supported by OK-VQA Marino et al.
-(2019) and WebQA Chang et al.
-(2022).
-For complex multimodal reasoning, MultimodalQA Talmor et al.
-(2021) integrates text, images, and tables, whereas video-text tasks leverage ActivityNet Caba Heilbron et al.
-(2015) and YouCook2 Zhou et al.
-(2018).
-In the medical domain, MIMIC-CXR Johnson et al.
-(2019) and CheXpert Irvin et al.
-(2019) facilitate tasks such as medical report generation.
-It is noteworthy that a number of these datasets are unimodal (e.g., solely text-based or image-based).
-Unimodal datasets are frequently employed to represent a specific modality and are subsequently integrated with complementary datasets from other modalities.
-This modular approach allows each dataset to contribute its domain-specific strengths, thereby enhancing the overall performance of the multimodal retrieval and generation processes.
-Benchmarks assess multimodal RAG systems on visual reasoning, external knowledge integration, and dynamic retrieval.
-The $M^{2}RAG$ Ma et al.
-(2024c) benchmark provides a unified evaluation framework that combines fine-grained text-modal and multimodal metrics to jointly assess both the quality of generated language and the effective integration of visual elements Vision-focused evaluations, including MRAG-Bench Hu et al.
-(2024c), VQAv2 Goyal et al.
-(2017a) and VisDoMBench Suri et al.
-(2024), test models on complex visual tasks.
-Dyn-VQA Li et al.
-(2024b), MMBench Liu et al.
-(2025), and ScienceQA Saikh et al.
-(2022) evaluate dynamic retrieval and multi-hop reasoning across textual, visual, and diagrammatic inputs.
-Knowledge-intensive benchmarks, such as TriviaQA Joshi et al.
-(2017) and Natural Questions Kwiatkowski et al.
-(2019), together with document-oriented evaluations such as OmniDocBench Ouyang et al.
-(2024), measure integration of unstructured and structured data.
-Advanced retrieval benchmarks such as RAG-Check Mortaheb et al.
-(2025a) evaluate retrieval relevance and system reliability, while specialized assessments such as Counterfactual VQA Niu et al.
-(2021) test robustness against adversarial inputs.
-Additionally, OCR impact studies such as OHRBench Zhang et al.
-(2024d) examine the cascading effects of errors on RAG systems.
-Additional details about datasets, benchmarks, and their categorization are presented in Table 1 and Table 2 in Appendix (§B).
+In this work, (i) we present a comprehensive review of multimodal RAG, covering task formulation, datasets, benchmarks, applications, and key innovations across retrieval, fusion, augmentation, generation, training strategies, loss functions, and agent frameworks.
+(ii) We propose a structured taxonomy (Figure 2) that categorizes state-of-the-art models by their core contributions, highlighting methodological advances and emerging trends.
+(iii) We provide open-access resources, including datasets, benchmarks, and implementation details, to facilitate future research.
+(iv) Finally, we identify research gaps and offer insights to guide future directions in this rapidly evolving field.
+2 Datasets, Evaluation, and Applications
+We review diverse datasets and benchmarks supporting tasks such as multimodal summarization, visual QA, video understanding, and more.
+For full details, see Appendix (§B) and Tables 1 and 2.
+Multimodal RAG has been applied across various domains, including healthcare, software engineering, fashion, entertainment, and emerging fields.
+An overview of tasks and applications are detailed in Appendix (§E) and  Figure 3.
+Evaluating these systems requires multiple metrics, covering retrieval performance, generation quality, and modality alignment.
+The complete evaluation methods, metrics, and their definitions and formulations are in Appendix (§C).
 {forest}
@@
 Similarity Retrieval
- (§4.1.1), fill=red!15 [ Maximum Inner Product Search (MIPS), fill=blue!15 [ TPU-KNN Chern et al.
+ (§3.1), fill=red!15 [ Maximum Inner Product Search (MIPS), fill=blue!15 [ TPU-KNN Chern et al.
 (2022), ScaNN Guo et al.
@@
 (2021), BLIP Li et al.
-(2022a), MARVEL Zhou et al.
+(2022), MARVEL Zhou et al.
 (2024c), ALIGN Jia et al.
@@
 (2022), Ovis Lu et al.
+(2024), LLaVE Lan et al.
 (2025) Mi-RAG Adjali et al.
 (2024) , leaf, text width=53em ] ] ] [ Modality-Centric
-Retrieval (§4.1.2), fill=red!15 [ Text-Centric, fill=blue!15 [ Contriever Izacard et al.
+Retrieval (§3.1), fill=red!15 [ Text-Centric, fill=blue!15 [ Contriever Izacard et al.
 (2022), GTE Li et al.
@@
 (2024), OMG-QA Nan et al.
-(2024), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al.
+(2024b), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al.
 (2024b), RAFT Zhang et al.
@@
 (2024), M2-RAG Ma et al.
-(2024c) , leaf, text width=53em ] ] [ Vision-Centric, fill=blue!15 [ VQA4CIR Feng et al.
+(2024d) , leaf, text width=53em ] ] [ Vision-Centric, fill=blue!15 [ VQA4CIR Feng et al.
 (2023), Unifashion Zhao et al.
@@
 Xue et al.
-(2024) , leaf, text width=53em ] ] [ Video-Centric, fill=blue!15 [ iRAG Arefeen et al.
+(2024b) , leaf, text width=53em ] ] [ Video-Centric, fill=blue!15 [ iRAG Arefeen et al.
 (2024), VideoRAG Ren et al.
@@
 (2022), DrVideo Ma et al.
-(2024d) , leaf, text width=53em ] ] [ Document Retrieval, fill=blue!15 [ ColPali Faysse et al.
+(2024e) , leaf, text width=53em ] ] [ Audio-Centric, fill=blue!15 [ CA-CLAP Xue et al.
+(2024a), Recap Ghosh et al.
+(2024), SpeechRAG Min et al.
+(2025), WavRAG Chen et al.
+(2025b), SEAL Sun et al.
+(2025), Audiobox TTA-RAG Yang et al.
+(2024a), DRCap Li et al.
+(2025c), P2PCAP Changin et al.
+(2024), LA-RAG Li et al.
+(2024b), Xiao et al.
+(2025) , leaf, text width=53em ] ] [ Document
+Retrieval, fill=blue!15 [ ColPali Faysse et al.
 (2024), ColQwen2 Wang et al.
@@
 (2024), DSE Ma et al.
+(2024a) , SV-RAG Chen et al.
 (2025a) , leaf, text width=53em ] ] ] [ Re-ranking
-Strategies (§4.1.3), fill=red!15 [ Optimized Example Selection, fill=blue!15 [ MSIER Luo et al.
+Strategies (§3.1), fill=red!15 [ Optimized
+Example
+Selection, fill=blue!15 [ MSIER Luo et al.
 (2024a), Hybrid RAG Su et al.
@@
 (2024g), LDRE Yang et al.
-(2024), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al.
+(2024b), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al.
 (2024), OMG-QA Nan et al.
-(2024), EchoSight Yan and Xie (2024), EgoInstructor Xu et al.
-(2024a) , leaf, text width=53em ] ] [ Filtering
+(2024b), EchoSight Yan and Xie (2024), EgoInstructor Xu et al.
+(2024a), VR-RAG Khan et al.
+(2025) , leaf, text width=53em ] ] [ Filtering
 Mechanisms, fill=blue!15 [ MAIN-RAG Chang et al.
@@
  (§3.2), fill=yellow!20 [ Score Fusion
-and Alignment
- (§4.2.1), fill=red!15 [ M3 Cai et al.
+and Alignment (§3.2), fill=red!15 [ M3 Cai et al.
 (2025)  Zhi Lim et al.
@@
 (2024),  Xue et al.
-(2024), RA-BLIP Ding et al.
+(2024b), RA-BLIP Ding et al.
 (2024b), Re-IMAGEN Chen et al.
@@
 (2024) , leaf, text width=62.0em ] ] [ Attention-Based
-Mechanisms (§4.2.2), fill=red!15 [ RAMM Yuan et al.
+Mechanisms (§3.2), fill=red!15 [ RAMM Yuan et al.
 (2023), EMERGE Zhu et al.
@@
 (2024) , leaf, text width=62.0em ] ] [ Unified Frameworks
-and Projections (§4.2.3), fill=red!15 [ Hybrid-RAG Su et al.
+and Projections (§3.2), fill=red!15 [ Hybrid-RAG Su et al.
 (2024a), Dense2Sparse Nguyen et al.
@@
 (2024), PDF-MVQA Ding et al.
-(2024c), SAM-RAG Zhai (2024), UFineBench Zuo et al.
+(2024d), SAM-RAG Zhai (2024), UFineBench Zuo et al.
 (2024),  Li et al.
-(2022a) , leaf, text width=62.0em ] ] ] [ Augmentation
+(2022) , leaf, text width=62.0em ] ] ] [ Augmentation
 Techniques
- (§4.3), fill=yellow!20 [ Context-Enrichment (§4.3.1), fill=red!15 [ EMERGE Zhu et al.
+ (§3.3), fill=yellow!20 [ Context-Enrichment (§3.3), fill=red!15 [ EMERGE Zhu et al.
 (2024b), MiRAG Adjali et al.
@@
 (2024e),  Xue et al.
-(2024) , leaf, text width=62.0em ] ] [ Adaptive and
-Iterative
-Retrieval (§4.3.2), fill=red!15 [ SKURG Yang et al.
+(2024b) , leaf, text width=62.0em ] ] [ Adaptive and Iterative
+Retrieval (§3.3), fill=red!15 [ SKURG Yang et al.
 (2023), IRAMIG Liu et al.
 (2024b), OMG-QA Nan et al.
-(2024), SAM-RAG Zhai (2024), MMed-RAG Xia et al.
+(2024b), SAM-RAG Zhai (2024), MMed-RAG Xia et al.
 (2024a), OmniSearch Li et al.
-(2024b), mR2AG Zhang et al.
+(2024d), mR2AG Zhang et al.
 (2024f), RAGAR Khaliq et al.
+(2024), UniversalRAG Yeo et al.
+(2025), OMGM Yang et al.
 (2025) , leaf, text width=62.0em ] ] ] [ Generation
-Technique
-s (§4.4), fill=yellow!20 [ In-Context
-Learning (§4.4.1), fill=red!15 [ RMR Tan et al.
+Techniques
+ (§3.4), fill=yellow!20 [ In-Context Learning  (§3.4), fill=red!15 [ RMR Tan et al.
 (2024),  Sharifymoghaddam et al.
@@
 (2024a), Raven Rao et al.
-(2024) , leaf, text width=62.0em ] ] [ Reasoning (§4.4.2), fill=red!15 [ RAGAR Khaliq et al.
+(2024) , leaf, text width=62.0em ] ] [ Reasoning (§3.4), fill=red!15 [ RAGAR Khaliq et al.
 (2024), VisDoM Suri et al.
 (2024), SAM-RAG Zhai (2024), LDRE Yang et al.
-(2024) , leaf, text width=62.0em ] ] [ Instruction
-Tuning (§4.4.3), fill=red!15 [ RA-BLIP Ding et al.
+(2024b) , leaf, text width=62.0em ] ] [ Instruction Tuning  (§3.4), fill=red!15 [ RA-BLIP Ding et al.
 (2024b), RAGPT Lang et al.
@@
 (2024f), RagVL Chen et al.
-(2024d), Jang et al.
+(2024e), Jang et al.
 (2024), MMed-RAG Xia et al.
@@
 (2024a), Rule Xia et al.
-(2024b) , leaf, text width=62.0em ] ] [ Source
-Attribution (§4.4.4), fill=red!15 [ MuRAR Zhu et al. (2025), VISA Ma et al. (2024b), OMG-QA Nan et al. (2024) , leaf, text width=62.0em ] ] ] [ Training
+(2024b) , leaf, text width=62.0em ] ] [ Source Attribution  (§3.4), fill=red!15 [ MuRAR Zhu et al.
+(2025), VISA Ma et al.
+(2024b), OMG-QA Nan et al.
+(2024b) , leaf, text width=62.0em ] ] [ Agentic Generation and Interaction  (§3.4), fill=red!15 [ AppAgent v2  Li et al.
+(2024c), USER-LLM R1 Rahimi et al.
+(2025), MMAD Jiang et al.
+(2025),  Yi et al.
+(2025), CollEX Schneider et al.
+(2025), HM-RAG Liu et al.
+(2025a), CogPlanner Yu et al.
+(2025) , leaf, text width=62.0em ] ] ] [ Training
 Strategies
- (§4.5), fill=yellow!20 [ Alignment (§4.5.1), fill=red!15 [ VISRAG Yu et al.
+ (§3.5), fill=yellow!20 [ Alignment (§3.5), fill=red!15 [ VISRAG Yu et al.
 (2024), MegaPairs Zhou et al.
@@
 (2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al.
-(2024) , leaf, text width=62.0em ] ] [ Robustness (§4.5.3), fill=red!15 [  Buettner and Kovashka (2024), MORE Cui et al.
+(2024) , leaf, text width=62.0em ] ] [ Robustness (§D.1), fill=red!15 [  Buettner and Kovashka (2024), MORE Cui et al.
 (2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al.
@@
 (2024b), RagVL Chen et al.
-(2024d), RA-CM3 Yasunaga et al.
+(2024e), RA-CM3 Yasunaga et al.
 (2023) , leaf, text width=62.0em ] ] ] ] ]
 Figure 2: Taxonomy of recent advances in Multimodal RAG. Refer to Appendix (§A) for further details.
-3 Evaluation
-Evaluating multimodal RAG models is complex due to their varied input types and complex structure.
-The evaluation combines metrics from VLMs, generative AI, and retrieval systems to assess capabilities like text/image generation and information retrieval.
-Our review found about 60 different metrics used in the field.
-More details, including the formulas for the RAG evaluation metrics, can be found in Appendix (§C).
-In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
-Retrieval Evaluation
-Retrieval performance is measured through accuracy, recall, and precision metrics, with an F1 score combining recall and precision.
-Recall@K, which examines relevant items in top K results, is preferred over standard recall.
-Mean Reciprocal Rank (MRR) serves as another key metric for evaluation, which is utilized by Adjali et al.
-(2024); Nguyen et al.
-(2024).
-Modality Evaluation
-Modality-based evaluations primarily focus on text and image, assessing their alignment, text fluency, and image caption quality.
-For text evaluation, metrics include Exact Match (EM), BLEU Papineni et al.
-(2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005).
-MultiRAGen Shohan et al.
-(2024) uses Multilingual ROUGE for multilingual settings.
-For image captioning, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al.
-(2015) measures caption quality using TF-IDF and cosine similarity Yasunaga et al.
-(2023); Zhao et al.
-(2024); Luo et al.
-(2024a); Yuan et al.
-(2024); Sharifymoghaddam et al.
-(2024); Hu et al.
-(2023); Rao et al.
-(2024); Xu et al.
-(2024a); Kim et al.
-(2024); Zhang et al.
-(2024c), while SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al.
-(2016) focuses on semantics.
-SPIDEr Liu et al.
-(2017), used in Zhang et al.
-(2024c), combines both metrics.
-For semantic alignment, BERTScore Zhang et al.
-(2020) compares BERT embeddings Sun et al.
-(2024b); Shohan et al.
-(2024), and evaluates fluency Chen et al.
-(2022a); Zhi Lim et al.
-(2024); Ma et al.
-(2024c).
-CLIP Score Hessel et al.
-(2021), used in Sharifymoghaddam et al.
-(2024); Zhang et al.
-(2024c), measures image-text similarity using CLIP Radford et al.
-(2021).
-For image quality, FID (Fréchet Inception Distance) Heusel et al.
-(2017) compares feature distributions Yasunaga et al.
-(2023); Zhao et al.
-(2024); Sharifymoghaddam et al.
-(2024); Zhang et al.
-(2024c), while KID (Kernel Inception Distance) Bińkowski et al.
-(2018) provides an unbiased alternative.
-Inception Score (IS) evaluates image diversity and quality through classification probabilities Zhi Lim et al.
-(2024).
-For audio evaluation, Zhang et al.
-(2024c) uses human annotators to assess sound quality (OVL) and text relevance (REL), while also employing Fréchet Audio Distance (FAD) Kilgour et al.
-(2019), an audio-specific variant of FID.
-System efficiency is measured through FLOPs, execution time, response time, and retrieval time per query Nguyen et al.
-(2024); Strand et al.
-(2024); Dang (2024); Zhou (2024).
-Domain-specific metrics include geodesic distance for geographical accuracy Zhou et al.
-(2024e), and Clinical Relevance for medical applications Lahiri and Hu (2024).
 3 Key Innovations and Methodologies
 3.1 Retrieval Strategy
-4.1.1 Efficient Search and Similarity Retrieval
+Efficient Search and Similarity Retrieval
 Modern multimodal RAG systems encode diverse input modalities into a unified embedding space to enable direct cross-modal retrieval.
-Recent advancements in CLIP-based Radford et al.
-(2021) and BLIP-inspired Li et al.
-(2022a) approaches have driven the evolution of contrastive learning strategies through novel multimodal retrieval architectures and training methodologies Zhou et al.
-(2024c); Wei et al.
-(2024b); Zhang et al.
-(2024i).
-As these multi-encoder models project different modalities into a shared latent space, multimodal RAGs rely on efficient search strategies to retrieve relevant external knowledge.
-Maximum inner product search (MIPS) variants are widely used for fast and direct similarity comparisons Tiwari et al.
+Early CLIP-based Radford et al.
+(2021) methods often struggled to balance retrieval precision and computational cost.
+BLIP-inspired Li et al.
+(2022) approaches addressed some of these trade-offs by integrating cross-modal attention during training, yielding richer alignments between visual and textual features.
+To reconcile high accuracy with efficiency, contrastive retrieval frameworks such as MARVEL Zhou et al.
+(2024c) and Uni-IR Wei et al.
+(2024a) improved inter-modal discrimination through hard-negative mining and balanced sampling strategies Zhang et al.
+(2024i); Lan et al.
+(2025).
+Despite these representational gains, direct search over millions of embeddings demands fast similarity computation.
+Maximum inner product search (MIPS) variants offer sublinear lookup by approximating top-$k$ inner products Tiwari et al.
 (2024); Wang et al.
@@
 (2023b).
+However, coarse quantization can degrade recall.
+To mitigate this, adaptive quantization methods Zhang et al.
+(2023a); Li et al.
+(2024a) dynamically allocate bits where the embedding distribution is dense, resulting in recall improvements over static schemes.
+Hybrid sparse–dense retrieval Nguyen et al.
+(2024); Zhang et al.
+(2024a) further complements dense embeddings with sparse lexical signals.
 Systems such as MuRAG Chen et al.
 (2022a) and RA-CM3 Yasunaga et al.
-(2023) employ approximate MIPS to efficiently retrieve the top-k candidates by maximizing the inner product between the query vector and a large collection of image–text embeddings.
+(2023) employ approximate MIPS for efficient top-k candidate retrieval from large collections of image–text embeddings.
 Large-scale implementations leverage distributed MIPS techniques, such as TPU-KNN Chern et al.
 (2022), for high-speed retrieval.
-Additionally, ScaNN (Scalable Nearest Neighbors) Guo et al.
+Other efficient similarity computation methods include ScaNN (Scalable Nearest Neighbors) Guo et al.
 (2020), MAXSIM score Chan and Ng (2008); Cho et al.
 (2024), and approximate KNN methods Caffagni et al.
-(2024) have been adopted for efficient similarity computation.
-Recent advancements in MIPS optimization focus on reducing retrieval latency and improving accuracy.
-These include adaptive quantization Zhang et al.
-(2023a); Li et al.
-(2024a), hybrid sparse-dense representations Nguyen et al.
-(2024); Zhang et al.
-(2024a), and learned index structures Zhai et al.
+(2024).
+Emerging approaches explore learned index structures Zhai et al.
 (2023); Basnet et al.
-(2024).
-These techniques help optimize retrieval performance by balancing efficiency and precision in multimodal RAG systems.
-4.1.2 Modality-Based Retrieval
-Modality-aware retrieval techniques optimize retrieval efficiency by leveraging the unique characteristics of each modality.
-Text-Centric Retrieval
-This remains foundational in multimodal RAG systems, with both traditional methods like BM25 Robertson and Zaragoza (2009), and dense retrievers such as MiniLM Wang et al.
+(2024), which embed the search tree itself in neural parameters, aiming to adapt retrieval paths to the data distribution and reduce both latency and storage overhead.
+Modality-Based Retrieval
+Modality-aware retrieval techniques optimize efficiency by leveraging the unique characteristics of each modality.
+(i) Text-centric retrieval remains foundational in multimodal RAG systems, with both traditional methods like BM25 Robertson and Zaragoza (2009) and dense retrievers such as MiniLM Wang et al.
 (2020a) and BGE-M3 Chen et al.
-(2024b) dominate text-based evidence retrieval Chen et al.
+(2024b) dominating text-based evidence retrieval Chen et al.
 (2022b); Suri et al.
 (2024); Nan et al.
-(2024).
+(2024b).
 Novel approaches also address the need for fine-grained semantic matching and domain specificity: For instance, ColBERT Khattab and Zaharia (2020) and PreFLMR Lin et al.
@@
 (2024h) and CRAG Yan et al.
-(2024) enhance retrieval performance by ensuring precise citation of text spans.
-Vision-Centric Retrieval
-This focuses on directly leveraging image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al.
+(2024) enhance retrieval by ensuring accurate citation of text spans.
+(ii) Vision-centric retrieval leverages image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al.
 (2023).
@@
 (2024) retrieve visually similar content by using reference images as queries.
-In addition, composed Image Retrieval (CMI) models Feng et al.
+In addition, composed image retrieval methods Feng et al.
 (2023); Zhao et al.
 (2024); Jang et al.
-(2024) enhance retrieval by integrating multiple image features into a unified query representation.
-Similarly, Pic2word Saito et al.
-(2023) maps visual content to textual descriptions, enabling zero-shot image retrieval.
-Video-Centric Retrieval
-These methods extend vision-based techniques by incorporating temporal dynamics and large video-language models (LVLMs), driven by novel frameworks like iRAG Arefeen et al.
-(2024), which introduces incremental retrieval for sequential video understanding, and MV-Adapter Jin et al.
-(2024), optimizing multimodal transfer learning for video-text retrieval.
-Recent breakthroughs focus on long-context processing: Video-RAG Luo et al.
-(2024b) leverages visually aligned auxiliary texts (OCR/ASR) to enhance retrieval without proprietary models, T-Mass Wang et al.
-(2024b) models text as a stochastic embedding to enhance text-video retrieval, and VideoRAG Ren et al.
-(2025) -employs dual-channel architectures with graph-based knowledge grounding for extreme-length videos.
-For temporal reasoning, CTCH Shen et al.
-(2024) uses contrastive transformer hashing to model long-term dependencies, while RTime Du et al.
-(2024) introduces reversed-video hard negatives to rigorously benchmark temporal causality.
-Meanwhile, for complex video understanding, OmAgent Zhang et al.
-(2024e) adopts a divide-and-conquer framework, and DRVideo Ma et al.
-(2024d) addresses long-video understanding with a document-based retrieval approach.
+(2024); Saito et al.
+(2023) integrate multiple image features into unified query representations, enabling zero-shot image retrieval.
+(iii) Video-centric retrieval extends vision-based techniques by incorporating temporal dynamics and large video-language models.
+For instance, iRAG Arefeen et al.
+(2024) enables incremental retrieval for sequential video understanding, addressing the need for temporal coherence, while T-Mass Wang et al.
+(2024b) uses stochastic text embeddings to improve robustness in text-video alignment.
+Tackling long-context processing, Video-RAG Luo et al.
+(2024b) avoids reliance on proprietary models by using auxiliary OCR/ASR texts, whereas VideoRAG Ren et al.
+(2025) employs dual-channel architectures and graph-based knowledge grounding for extreme-length videos.
+To capture temporal reasoning, CTCH Shen et al.
+(2024) applies contrastive transformer hashing for long-term dependencies, which RTime Du et al.
+(2024) further refines by introducing reversed-video hard negatives for more robust causality benchmarking.
+Finally, OmAgent Zhang et al.
+(2024e) addresses the challenge of complex video understanding with a divide-and-conquer framework, while DRVideo Ma et al.
+(2024e) takes a complementary document-centric approach to enhance narrative preservation.
+(iv) Audio-centric retrieval aims to bypass traditional ASR pipelines while improving contextual alignment and real-time processing Xue et al.
+(2024a); Ghosh et al.
+(2024); Min et al.
+(2025).
+Pioneering frameworks like WavRAG Chen et al.
+(2025b) and SEAL Sun et al.
+(2025) introduce unified embedding architectures, directly mapping raw audio into a shared latent space to enable retrieval from hybrid knowledge bases.
+Audiobox TTA-RAG Yang et al.
+(2024a) conditions text-to-audio synthesis on retrieved acoustic samples, thereby enhancing zero-shot performance by enriching prompts with unlabeled audio context.
+For audio captioning, DRCap Li et al.
+(2025c) bridges the audio-text latent space of CLAP Wu et al.
+(2023) via text-only training for domain-adaptable descriptions without paired data.
+In parallel, P2PCAP Changin et al.
+(2024) improves retrieval precision by regenerating captions as dynamic queries.
+Further innovations address error correction and efficiency.
+LA-RAG Li et al.
+(2024b) utilizes fine-grained speech-to-speech retrieval and forced alignment to enhance ASR accuracy through LLM in-context learning.
+Meanwhile, hybrid systems, such as Xiao et al.
+(2025), integrate LLMs to correct errors in noisy environments using retrieved text/audio context.
 Document Retrieval and Layout Understanding
-Recent research has advanced beyond traditional uni-modal retrieval methods toward models that directly process entire document pages.
-These approaches integrate textual, visual, and structural elements like tables, font styles, and page layouts to enhance retrieval performance in complex documents.
+Recent research has moved beyond traditional unimodal retrieval, developing models that process entire documents by integrating textual, visual, and layout information.
 ColPali Faysse et al.
@@
 Models like ColQwen2 Wang et al.
-(2024d); Khattab and Zaharia (2020) and M3DocVQA Cho et al.
+(2024d); Faysse et al.
+(2024) and M3DocVQA Cho et al.
 (2024) extend this paradigm with dynamic resolution handling and holistic multi-page reasoning.
@@
 (2024a) pre-train generative models to align spatial layouts with text, while CREAM Zhang et al.
-(2024b) employs coarse-to-fine retrieval with multimodal efficient tuning to balance accuracy and computation costs.
+(2024b) employs coarse-to-fine retrieval with multimodal efficient tuning to balance accuracy and computational costs.
 Finally, mPLUG-DocOwl 1.5 Hu et al.
 (2024a) and 2 Hu et al.
-(2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies.
-4.1.3 Re-ranking and Selection Strategies
+(2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies, while SV-RAG Chen et al.
+(2025a) leverages MLLMs’ intrinsic retrieval capabilities via dual LoRA adapters: one for evidence page retrieval and the other for question answering.
+Re-ranking and Selection Strategies
 Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also prioritizing retrieved candidates.
-Re-ranking and selection strategies enhance retrieval quality by optimizing example selection, refining relevance scoring, and applying filtering mechanisms.
-Optimized Example Selection
-These techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection strategies Luo et al.
+Re-ranking and selection strategies improve retrieval quality through optimized example selection, refined relevance scoring, and filtering mechanisms.
+(i) Optimized example selection techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection approaches Luo et al.
 (2024a); Yuan et al.
 (2023).
-For instance, Su et al.
-(2024a) enhance multimodal inputs using probabilistic control keywords to improve credibility, RULE Xia et al.
-(2024b) calibrates retrieved context selection via statistical methods like the Bonferroni correction to mitigate factuality risks, and clustering-based key-frame selection ensures diversity in video-based retrieval Dong et al.
+Supervised methods like Su et al.
+(2024a) enhance multimodal inputs using probabilistic control keywords, whereas RULE Xia et al.
+(2024b) calibrates retrieved context via statistical methods like the Bonferroni correction Haynes (2013) to mitigate factuality risks.
+Clustering-based key-frame selection ensures diversity in video-based retrieval Dong et al.
 (2024b).
-Relevance Score Evaluation
-Several methods employ advanced scoring mechanisms to improve retrieval relevance Mortaheb et al.
+Advanced (ii) scoring mechanisms are employed by several methods to improve retrieval relevance Mortaheb et al.
 (2025b, a); Zhi Lim et al.
@@
 (2020b), normalized cross-correlation (NCC), and BERTScore Zhang et al.
-(2020), aid in re-ranking retrieved documents.
-Cross-encoders trained on sequence classification tasks refine relevance scoring within retrieval pipelines Zhi Lim et al.
-(2024), while hierarchical post-processing techniques integrate retrieval, passage-level, and answer confidence scores for improved ranking Zhang et al.
+(2020), aid in re-ranking documents.
+Some frameworks combine similarity scores derived from various modalities for more robust re-ranking.
+For example, VR-RAG Khan et al.
+(2025) proposes a visual re-ranking framework that combines cross-modal text-image similarity with intra-modal visual similarity using DINOv2 Oquab et al.
+(2023), demonstrating significant improvements in open-vocabulary recognition tasks.
+Hierarchical post-processing integrates passage-level and answer confidence scores for improved ranking Zhang et al.
 (2024g); Yan and Xie (2024); Xu et al.
@@
 LDRE Yang et al.
-(2024) employs semantic ensemble methods to adaptively weigh multiple caption features, whereas RAGTrans Cheng et al.
+(2024b) employs semantic ensemble methods to adaptively weigh multiple caption features, while RAGTrans Cheng et al.
 (2024) and OMG-QA Nan et al.
-(2024) incorporate traditional ranking functions like BM25 Robertson and Zaragoza (2009).
-Filtering Mechasnism
-This ensures high-quality retrieval by eliminating irrelevant data.
+(2024b) incorporate traditional ranking functions like BM25 Robertson and Zaragoza (2009).
+(iii) Filtering methods ensure high-quality retrieval by eliminating irrelevant data.
 Hard negative mining, as used in GME Zhang et al.
@@
 (2024a), mitigates modality bias through modality-aware sampling and synthesized negatives.
-Similarly, consensus-based filtering approaches, seen in MuRAR Zhu et al.
+Similarly, consensus-based filtering, seen in MuRAR Zhu et al.
 (2025) and ColPali Faysse et al.
 (2024), employs source attribution and multi-vector mapping to filter out low-similarity candidates.
-Moreover, dynamic modality filtering methods, such as RAFT Zhang et al.
-(2024h), Img2Loc Zhou et al.
-(2024e), and MAIN-RAG Chang et al.
-(2024), train retrievers to disregard confusing data, thereby enhancing the discriminative capacity and overall robustness of multi-modal retrieval systems.
+Dynamic modality filtering methods, such as RAFT Zhang et al.
+(2024h) and MAIN-RAG Chang et al.
+(2024), train retrievers to disregard confusing data, improving multimodal retrieval robustness.
 3.2 Fusion Mechanisms
-4.2.1 Score Fusion and Alignment
+Score Fusion and Alignment
 Models in this category utilize distinct strategies to align multimodal representations.
@@
 (2021) and BLIP feature fusion Li et al.
-(2022a).
-Riedler and Langer (2024), Wiki-LLaVA Caffagni et al.
+(2022).
+Wiki-LLaVA Caffagni et al.
 (2024), C3Net Zhang et al.
-(2024c), and MegaPairs Zhou et al.
-(2024a), embed images and queries into a shared CLIP space.
+(2024c), Riedler and Langer (2024), and MegaPairs Zhou et al.
+(2024a) embed images and queries into a shared CLIP space.
+In particular, MegaPairs Zhou et al.
+(2024a) scales this approach by integrating both CLIP-based and MLLM-based retrieval, fusing their scores to leverage complementary strengths, but at the cost of increased inference complexity.
 VISA Ma et al.
@@
 (2024) aligns video-text embeddings via a BLIP-inspired Query Transformer Li et al.
-(2022a).
+(2022).
 LLM-RA Jian et al.
@@
 Xue et al.
-(2024) use a prototype-based embedding network Zheng et al.
+(2024b) use a prototype-based embedding network Zheng et al.
 (2023) to map object-predicate pairs into a shared semantic space, aligning visual features with textual prototypes.
@@
 To improve multimodal alignment, VISRAG Yu et al.
-(2024) enhances alignment with position-weighted mean pooling on VLM hidden states, prioritizing later tokens for relevance, and RAG-Driver Yuan et al.
-(2024) aligns visual–language embeddings using visual instruction tuning and an MLP projector.
-4.2.2 Attention-Based Mechanisms
-Attention-based methods dynamically weight cross-modal interactions to support task-specific reasoning.
-EMERGE Zhu et al.
+(2024) applies position-weighted mean pooling over VLM hidden states, giving higher relevance to later tokens.
+In contrast, RAG-Driver Yuan et al.
+(2024) aligns visual and language embeddings through visual instruction tuning and an MLP projector.
+Attention-Based Mechanisms
+Attention-based methods dynamically modulate cross-modal interactions to enable fine-tuned reasoning across tasks, balancing specificity and interpretability.
+Cross-attention is frequently used to integrate heterogeneous modalities, as in EMERGE Zhu et al.
 (2024b), MORE Cui et al.
-(2024), and AlzheimerRAG Lahiri and Hu (2024) integrate heterogeneous data via cross-attention.
+(2024), and AlzheimerRAG Lahiri and Hu (2024), though often requiring task-specific attention heads.
 RAMM Yuan et al.
@@
 (2024) applies user-aware attention to social media features.
-For video-text alignment, MV-Adapter Jin et al.
-(2024) leverages Cross Modality Tying to align embeddings, and M2-RAAP Dong et al.
-(2024b) enhances fusion through an auxiliary caption-guided strategy that re-weights frames and text captions based on intra-modal similarity.
-A mutual-guided alignment head then filters misaligned features using dot-product similarity and frame-to-token attention, generating refined frame-specific text representations.
+MV-Adapter Jin et al.
+(2024) introduces Cross-Modality Tying to align video-text embeddings by sharing latent factors, improving robustness but reducing granularity of modality-specific features.
+M2-RAAP Dong et al.
+(2024b) enhances fusion through an auxiliary caption-guided strategy that re-weights frames and text captions based on intra-modal similarity, then uses a mutual-guided alignment head to filter misaligned features via dot-product similarity and frame-to-token attention; however, this method is computationally intensive.
 Xu et al.
-(2024a) condition text generation on visual features using gated cross-attention, and Mu-RAG Chen et al.
+(2024a) condition text generation on visual features using gated cross-attention, optimizing controllability but requiring aligned supervision, and Mu-RAG Chen et al.
 (2022a) employs intermediate cross-attention for open-domain QA.
@@
 (2024) leverage cross-modal memory retrieval with pre-trained CLIP ViT-L/14 to map video-text pairs into a shared space, enabling dense captioning through the attention-based fusion of retrieved memories.
-4.2.3 Unified Frameworks and Projections
+Unified Frameworks and Projections
 Unified frameworks and projection methods consolidate multimodal inputs into coherent representations.
@@
 (2024a) employ hierarchical cross-chains and late fusion for healthcare data, while IRAMIG Liu et al.
-(2024b) iteratively integrates multimodal results into unified knowledge representations.
+(2024b) iteratively integrates multimodal results into unified knowledge representations, enhancing consistency but requiring multiple reasoning passes.
 M3DocRAG Cho et al.
 (2024) flattens multi-page documents into a single embedding tensor, and PDF-MVQA Ding et al.
-(2024c) fuses Region-of-Interest (RoI)-based and patch-based (CLIP) vision-language models Long et al.
-(2022).
+(2024d) proposes a joint-grained retriever that fuses coarse-grained semantic entity representations with their fine-grained token-level textual content, creating a richer, unified representation.
 DQU-CIR Wen et al.
 (2024) unifies raw data by converting images into text captions for complex queries and overlaying text onto images for simple ones, then fusing embeddings via MLP-learned weights.
-SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input into unimodal text for subsequent processing.
+SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input to unimodal text for subsequent processing.
 UFineBench Zuo et al.
-(2024) utilizes a shared granularity decoder for ultra-fine text–person retrieval.
+(2024) uses a shared granularity decoder for ultra-fine text–person retrieval.
 Nguyen et al.
 (2024) introduce Dense2Sparse projection, converting dense embeddings from models like BLIP/ALBEF Li et al.
-(2022a) into sparse lexical vectors using layer normalization and probabilistic expansion control to optimize storage and interpretability.
+(2022) into sparse lexical vectors using layer normalization and probabilistic expansion control to optimize storage and interpretability.
 3.3 Augmentation Techniques
-Basic RAG systems typically follow a single retrieval step and pass retrieved content directly to the generation phase without further refinement, which can lead to inefficiencies and suboptimal outputs.
-To address this, augmentation techniques have been introduced to refine retrieved data beforehand, improving multimodal interpretation, structuring, and integration Gao et al.
+Basic RAG systems typically retrieve content in a single step, directly passing it to generation, often leading to inefficiencies and suboptimal outputs.
+Augmentation techniques refine retrieved data beforehand, improving multimodal interpretation, structuring, and integration Gao et al.
 (2023).
-4.3.1 Context Enrichment
+Context Enrichment
 This focuses on enhancing the relevance of retrieved knowledge by refining or expanding retrieved data.
@@
 (2024); Xue et al.
-(2024).
+(2024b).
 EMERGE Zhu et al.
@@
 Img2Loc Zhou et al.
-(2024e) enhances accuracy by including both the most similar and the most dissimilar points from the database in the prompt, allowing the model to rule out implausible locations for its predictions.
-4.3.2 Adaptive and Iterative Retrieval
+(2024e) boosts accuracy by including both similar and dissimilar points in prompts, helping rule out implausible locations.
+Adaptive and Iterative Retrieval
 For more complex queries, dynamic retrieval mechanisms have proven effective.
 Adaptive retrieval approaches optimize relevance by adjusting retrieval dynamically.
+For instance, UniversalRAG Yeo et al.
+(2025) introduces a framework that adapts retrieval by dynamically routing queries to the most suitable corpus based on both the required modality and granularity (e.g., paragraph vs.
+document, clip vs.
+full video), thereby addressing the specific knowledge type and scope demanded by the query.
 SKURG Yang et al.
@@
 (2024a) further improves retrieval precision by discarding low-relevance results, while OmniSearch Li et al.
-(2024b) introduces a self-adaptive retrieval agent that dynamically decomposes complex multimodal questions into sub-question chains and plans retrieval actions in real-time based on retrieved content.
+(2024d) decomposes multimodal queries into structured sub-questions, planning retrieval actions in real time.
 Iterative approaches refine results over multiple steps by incorporating feedback from prior iterations.
+For example, OMGM Yang et al.
+(2025) orchestrates a multi-step, coarse-to-fine retrieval process for knowledge-based visual question answering, starting with a broad entity search and progressively refining the selection through multimodal reranking and fine-grained textual filtering to pinpoint the most relevant knowledge, achieving superior retrieval performance in comparison to prior methods.
 IRAMIG Liu et al.
 (2024b) improves multimodal retrieval by dynamically updating queries based on retrieved content.
-Similarly, OMG-QA Nan et al.
-(2024) introduces a multi-round retrieval strategy, where each retrieval step incorporates episodic memory to refine subsequent queries.
-An evaluation module assesses retrieval effectiveness at each step, guiding the refinement of subsequent retrieval efforts through feedback.
+OMG-QA Nan et al.
+(2024b) integrates episodic memory to refine retrieval across multiple rounds, ensuring continuity in reasoning.
 RAGAR Khaliq et al.
@@
 3.4 Generation Techniques
-Recent advancements in multimodal RAG generation focus on robustness, cross-modal coherence, and task-specific adaptability. These innovations can be broadly categorized into these sections:
-4.4.1 In-Context Learning
-In-context learning (ICL) with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining.
+In-Context Learning (ICL)
+ICL with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining.
 Models such as RMR Tan et al.
@@
 RAG-Driver Yuan et al.
-(2024) refines ICL by retrieving relevant driving experiences from a memory database, ensuring scenario-specific contextual alignment.
+(2024) refines ICL by retrieving relevant driving experiences from a memory database.
 MSIER Luo et al.
-(2024a) further enhances example selection through a Multimodal Supervised In-Context Examples Retrieval framework, leveraging a foundation MLLM scorer to evaluate both textual and visual relevance.
-Meanwhile, Raven Rao et al.
-(2024) introduces Fusion-in-Context Learning, a novel approach that enriches ICL by incorporating a more diverse set of in-context examples, leading to better performance than standard ICL.
-4.4.2 Reasoning
-Structured reasoning techniques, such as chain-of-thought (CoT), decompose complex reasoning into smaller sequential steps, enhancing coherence and robustness in multimodal RAG systems.
+(2024a) improves example selection with a multimodal supervised in-context examples retrieval framework, using an MLLM scorer to assess textual and visual relevance.
+Raven Rao et al.
+(2024) introduces Fusion-in-Context Learning, integrating diverse in-context examples for superior performance over standard ICL.
+Reasoning
+Reasoning methods, like chain of thought (CoT), decompose complex reasoning into sequential steps, improving coherence and robustness in multimodal RAG.
 RAGAR Khaliq et al.
-(2024) introduces Chain of RAG and Tree of RAG to iteratively refine fact-checking queries and explore branching reasoning paths for more robust evidence generation.
-VisDoM Suri et al.
-(2024) integrates CoT with evidence curation to ensure logical and contextual accuracy, while SAM-RAG Zhai (2024) employs reasoning chains alongside multi-stage answer verification to enhance the relevance, utility, and support of generated responses.
-Meanwhile, LDRE Yang et al.
-(2024) leverages LLMs for divergent compositional reasoning, generating refined captions by incorporating dense captions and modification text.
-4.4.3 Instruction Tuning
+(2024) refines fact-checking queries and explores branching reasoning paths by introducing Chain of RAG and Tree of RAG, while VisDoM Suri et al.
+(2024) and SAM-RAG Zhai (2024) integrate CoT with evidence curation and multi-stage verification to enhance accuracy and support.
+Notably, VisDoM performs well in scenarios where key information is distributed across modalities.
+LDRE Yang et al.
+(2024b) applies LLMs for divergent compositional reasoning by refining captions using dense descriptions and textual modifications, achieving superior zero-shot composed image retrieval results.
+Instruction Tuning
 Several works have fine-tuned or instruct-tuned generation components for specific applications.
@@
 MR2AG Zhang et al.
-(2024f) uses instruction tuning with the mR2AG-IT dataset to train MLLMs to adaptively invoke retrieval, identify relevant evidence, and generate accurate answers for knowledge-based VQA tasks.
-RagVL Chen et al.
-(2024d) employs instruction tuning to enhance the ranking capability of MLLMs, serving them as a re-ranker for filtering the top-k retrieved images.
+(2024f) and RagVL Chen et al.
+(2024e) train MLLMs to invoke retrieval adaptively, identify relevant evidence, and enhance ranking capabilities for improved response accuracy.
 Jang et al.
@@
 (2024a) construct multimodal instruction-tuning datasets from prior LLM errors, while Rule Xia et al.
-(2024b) refines Med-LVLM through direct preference optimization to mitigate overreliance on retrieved contexts.
-4.4.4 Source Attribution and Evidence Transparency
-Ensuring source attribution in multimodal RAG systems is a key focus of recent research.
+(2024b) refines a medical large vision language model through direct preference optimization to mitigate overreliance on retrieved contexts.
+Source Attribution and Evidence Transparency
+Ensuring source attribution in multimodal RAG systems is a significant research focus.
+OMG-QA Nan et al.
+(2024b) prompts LLMs for explicit evidence citation in generated responses.
 MuRAR Zhu et al.
-(2025) integrates multimodal data, fetched by a source-based retriever, to refine LLM’s initial response, ensuring informativeness.
-VISA Ma et al.
-(2024b) uses large vision-language models to generate answers with visual source attribution by identifying and highlighting supporting evidence in retrieved document screenshots.
-Similarly, OMG-QA Nan et al.
-(2024) ensures transparency by prompting the LLM to explicitly cite evidence in generated responses.
+(2025) refines an LLM’s initial response by integrating multimodal information from a source-based retriever to improve informativeness.
+However, its recall is constrained, as the retriever may miss evidence spanning different sections or web documents.
+Similarly, VISA Ma et al.
+(2024b) employs vision-language models to generate answers with visual source attribution by highlighting evidence in retrieved screenshots.
+Nevertheless, its attribution accuracy degrades when evidence spans multiple sections or requires cross-modal integration.
+Agentic Generation and Interaction
+Agent-driven multimodal RAG uses versatile autonomous/semi-autonomous systems across diverse interaction paradigms and specialized domains, often generating complex outputs.
+For user interaction, AppAgent v2 Li et al.
+(2024c) enables mobile GUI navigation while USER-LLM R1 Rahimi et al.
+(2025) creates personalized conversational agents via dynamic profiling, particularly for elderly users.
+In specialized applications, MMAD Jiang et al.
+(2025) addresses industrial anomaly detection with training-free enhancement strategies, Yi et al.
+(2025) improve clinical report generation while reducing hallucination, and CollEX Schneider et al.
+(2025) facilitates scientific collection exploration for researchers and learners.
+For complex reasoning, HM-RAG Liu et al.
+(2025a) coordinates hierarchical multi-agent collaboration across multimodal data streams, while CogPlanner Yu et al.
+(2025) introduces a cognitively inspired planning framework that iteratively refines queries and selects retrieval strategies adaptively.
 3.5 Training Strategies
-Training multimodal RAG models involves a multi-stage process to effectively handle cross-modal interactions Chen et al.
+Training multimodal RAG models follows a multi-stage process to effectively capture cross-modal interactions Chen et al.
 (2022a).
-Pretraining establishes the foundation using large paired datasets to learn cross-modal relationships while fine-tuning adapts models to specific tasks by leveraging cross-modal attention Ye et al.
+Pretraining on large paired datasets establishes cross-modal relationships, while fine-tuning adapts models to task-specific objectives by aligning outputs with task requirements Ye et al.
 (2019).
-For instance, REVEAL Hu et al.
+For example, REVEAL Hu et al.
 (2023) integrates multiple training objectives.
-Its pretraining phase optimizes Prefix Language Modeling Loss ($L_{\text{PrefixLM}}$), where the model predicts text continuations from a prefix and an image.
-Supporting losses include Contrastive Loss ($L_{\text{contra}}$) for aligning queries with pseudo-ground-truth knowledge, Disentangled Regularization Loss ($L_{\text{decor}}$) to improve embedding expressiveness, and Alignment Regularization Loss ($L_{\text{align}}$) to align query and knowledge embeddings.
-During fine-tuning, a cross-entropy objective trains the model for tasks like VQA and image captioning.
-Details of formulas for widely used RAG loss functions can be found in Appendix (§D).
-4.5.1 Alignment
+Its pretraining phase optimizes Prefix Language Modeling Loss ($L_{\text{PrefixLM}}$), where text is predicted from a given prefix and an associated image.
+Supporting losses include Contrastive Loss ($L_{\text{contra}}$) which aligns queries with pseudo-ground-truth knowledge, Disentangled Regularization Loss ($L_{\text{decor}}$) to enhance embedding expressiveness, and Alignment Regularization Loss ($L_{\text{align}}$) to refine query-knowledge alignment.
+Fine-tuning employs a cross-entropy objective for downstream tasks like visual question answering or image captioning.
+Details on robustness advancements and loss formulations are in Appendix (§D).
+Alignment
 Contrastive learning improves representation quality by pulling positive pairs closer and pushing negative pairs apart in the embedding space.
-A common objective is the InfoNCE loss van den Oord et al.
-(2019), which maximizes the mutual information between positive pairs while minimizing similarity to negatives.
-Several multimodal RAG models, such as VISRAG Yu et al.
+The InfoNCE loss van den Oord et al.
+(2019) is widely employed in multimodal RAG models, including VISRAG Yu et al.
 (2024), MegaPairs Zhou et al.
-(2024a) and SAM-RAG Zhai (2024) utilize InfoNCE loss to improve retrieval-augmented generation.
-Furthermore, EchoSight Yan and Xie (2024) enhances retrieval accuracy by selecting visually similar yet contextually distinct negatives, while HACL Jiang et al.
-(2024) improves generation by introducing hallucinative captions as distractors.
+(2024a), and SAM-RAG Zhai (2024), to improve retrieval-augmented generation.
+Several models introduce refinements to contrastive training.
+EchoSight Yan and Xie (2024) enhances retrieval accuracy by selecting visually similar yet semantically distinct negatives, while HACL Jiang et al.
+(2024) mitigates hallucinations by incorporating adversarial captions as distractors.
 Similarly, UniRaG Zhi Lim et al.
-(2024) strengthens retrieval by incorporating hard negative documents, helping the model distinguish relevant contexts from noise.
-The eCLIP loss Kumar and Marttinen (2024) extends contrastive training by integrating expert-annotated data and an auxiliary Mean Squared Error (MSE) loss to refine embedding quality.
-Mixup strategies further augment training by generating synthetic positive pairs, improving generalization in contrastive learning Kumar and Marttinen (2024).
+(2024) improves retrieval robustness by leveraging hard negative documents to help the model discriminate between relevant and irrelevant contexts.
+The eCLIP loss Kumar and Marttinen (2024) extends contrastive learning by integrating expert-annotated data and an auxiliary MSE loss to refine embedding quality.
+Mixup strategies further improve generalization by generating synthetic positive pairs Kumar and Marttinen (2024).
 Dense2Sparse Nguyen et al.
-(2024) incorporates two unidirectional losses: an image-to-caption loss $\ell(I\to C)$ and a caption-to-image loss $\ell(C\to I)$.
-It enforces sparsity through L1 regularization, optimizing retrieval precision by balancing dense and sparse representations.
-4.5.2 Generation
-A key aspect of multimodal RAG is the generation ability.
-Autoregressive language models are typically trained using Cross-Entropy Loss Brown et al.
-(2020).
-For image generation, widely used approaches include Generative Adversarial Networks (GANs) Goodfellow et al.
-(2014) and Diffusion Models Ho et al.
-(2020).
-GANs employ various loss functions, such as Binary Cross-Entropy Loss, Minimax Loss, Wasserstein Loss (WGAN), and Hinge Loss.
-Diffusion Models utilize Mean Squared Error (MSE) Loss for noise prediction, a common approach in Denoising Diffusion Probabilistic Models (DDPMs) Ho et al.
-(2020).
-4.5.3 Robustness and Noise Management
-Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024).
-Managing noisy retrieval inputs is critical for maintaining model performance.
-MORE Cui et al.
-(2024) injects irrelevant results during training to enhance focus on relevant inputs.
-AlzheimerRAG Lahiri and Hu (2024) uses progressive knowledge distillation to reduce noise while maintaining multimodal alignment.
-RAGTrans Cheng et al.
-(2024) leverages hypergraph-based knowledge aggregation to refine multimodal representations, ensuring more effective propagation of relevant information.
-RA-BLIP Ding et al.
-(2024b) introduces the Adaptive Selection Knowledge Generation (ASKG) strategy, which leverages the implicit capabilities of LLMs to filter relevant knowledge for generation through a denoising-enhanced loss term, eliminating the need for fine-tuning.
-This approach achieves strong performance compared to baselines while significantly reducing computational overhead by minimizing trainable parameters.
-RagVL Chen et al.
-(2024d) improves robustness through noise-injected training by adding hard negative samples at the data level and applying Gaussian noise with loss reweighting at the token level, enhancing the model’s resilience to multimodal noise.
-Finally, RA-CM3 Yasunaga et al.
-(2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.
-5 Tasks Addressed by Multimodal RAGs
-Multimodal RAG systems extend RAG beyond unimodal settings to tasks requiring cross-modal integration, enhancing performance across modalities such as text, images, and audio.
-In content generation, these models enhance image captioning Zhi Lim et al.
-(2024); Hu et al.
-(2023); Rao et al.
-(2024) and text-to-image synthesis Yasunaga et al.
-(2023); Chen et al.
-(2022b) by retrieving relevant contextual information.
-They also improve coherence in visual storytelling and ensure factual alignment in multimodal summarization Tonmoy et al.
-(2024).
-In knowledge-intensive applications, multimodal RAG supports open-domain and knowledge-seeking question answering Chen et al.
-(2024d); Ding et al.
-(2024b); Yuan et al.
-(2023), video-based QA Luo et al.
-(2024b), and fact verification Khaliq et al.
-(2024), grounding responses in retrieved knowledge and thereby mitigating hallucinations.
-Cross-modal retrieval further advances zero-shot image–text retrieval Yang et al.
-(2024); Dong et al.
-(2024b).
-Additionally, the recent incorporation of chain-of-thought reasoning Zhai (2024); Khaliq et al.
-(2024) has enhanced its ability to support complex problem solving and logical inference.
-Finally, integrating multimodal RAG into interactive agents and AI assistants such as Gemini Team et al.
-(2024) enables natural language-driven visual search, document understanding, and multimodal reasoning.
-The taxonomy of application domains can be seen in Figure 3.
-The following sections explore domain-specific adaptations of these techniques in greater depth.
-Healthcare and Medicine
-Multimodal RAG enhances clinical decision-making through integrated analysis of medical imaging, electronic health records, and biomedical literature.
-Systems like MMED-RAG Xia et al.
-(2024a) address diagnostic uncertainty in medical visual question answering by aligning radiology images with contextual patient data.
-RULE Xia et al.
-(2024b) mitigates hallucinations in automated report generation through dynamic retrieval of clinically similar cases.
-AsthmaBot Bahaj and Ghogho (2024) introduces a multimodal RAG-based approach for supporting asthma patients across multiple languages, enabling structured, language-specific semantic searches.
-Predictive frameworks such as Realm Zhu et al.
-(2024c) demonstrate robust risk assessment by fusing heterogeneous patient data streams, while Hybrid RAG Su et al.
-(2024a) advances privacy-preserving architectures for federated clinical data integration.
-FactMM-RAG Sun et al.
-(2024b) automates radiology report drafting by retrieving biomarker correlations from medical ontologies, exemplifying the paradigm’s capacity to operationalize expert knowledge at scale.
-Software Engineering
-Code generation systems leverage multimodal RAG to synthesize context-aware solutions from technical documentation and version histories.
-DocPrompting Zhou et al.
-(2023) improves semantic coherence in code completion by retrieving API specifications and debugging patterns.
-Commit message generation models like RACE Shi et al.
-(2022) contextualize code diffs against historical repository activity, while CEDAR Nashid et al.
-(2023) optimizes few-shot learning through retrieval-based prompt engineering.
-REDCODER Parvez et al.
-(2021) enhances code summarization via semantic search across open-source repositories, preserving syntactic conventions across programming paradigms.
-Fashion and E-Commerce
-Cross-modal alignment drives advancements in product discovery and design automation.
-UniFashion Zhao et al.
-(2024) enables style-aware retrieval by jointly embedding garment images and textual descriptors, while Dang (2024) reduces search friction through multimodal query expansion.
-LLM4DESIGN Chen et al.
-(2024c) demonstrates architectural design automation by retrieving compliance constraints and environmental impact assessments, underscoring RAG’s adaptability to creative domains.
-Entertainment and Social Computing
-Multimedia analytics benefit from RAG’s capacity to correlate heterogeneous signals.
-SoccerRAG Strand et al.
-(2024) derives tactical insights by linking match footage with player statistics.
-MMRA Zhong et al.
-(2024) predicts content virality through joint modeling of visual aesthetics and linguistic engagement patterns.
-Emerging Applications
-Autonomous systems adopt multimodal RAG for explainable decision-making, as seen in RAG-Driver’s Yuan et al.
-(2024) real-time retrieval of traffic scenarios during navigation.
-ENWAR Nazar et al.
-(2024) enhances wireless network resilience through multi-sensor fusion, while Riedler and Langer (2024) streamline equipment maintenance by retrieving schematics during fault diagnosis.
-Geospatial systems such as Img2Loc Zhou et al.
-(2024e) advance image geolocalization through cross-modal landmark correlation.
+(2024) employs image-to-caption $\ell(I\to C)$ and caption-to-image $\ell(C\to I)$ losses, while enforcing sparsity through $\ell 1$ regularization, optimizing retrieval precision by balancing dense and sparse representations.
 4 Open Problems and Future Directions
-Despite rapid advancements in multimodal RAG systems, fundamental challenges remain in achieving robust, efficient, and human-like reasoning across modalities.
-6.1 Generalization, Explainability, and Robustness
+Additional challenges and future directions about long-context processing, scalability, efficiency, and personalization are discussed in Appendix (§F).
+Generalization, Explainability, and Robustness
 Multimodal RAG systems often struggle with domain adaptation and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al.
 (2020).
-Explainability remains a major challenge, as these systems typically fail to attribute answers to precise sources.
-Current methods often cite entire documents or large visual regions as source attribution rather than identifying the exact part of an image, speech, or other modality that led to the answer Ma et al.
+Explainability remains a major challenge, as these systems often attribute responses to broad sources, citing entire documents or large visual regions instead of pinpointing exact contributing elements across modalities Ma et al.
 (2024b); Hu et al.
 (2023).
-In addition, the interplay between modalities affects the quality of outcomes produced by these models; for example, answers derived solely from text sources may differ in quality compared to those requiring a combination of text and image inputs Baltrusaitis et al.
+Moreover, the interplay between modalities affects the outcome quality; for example, answers derived solely from text sources may differ in quality compared to those requiring a combination of text and image inputs Baltrusaitis et al.
 (2019).
-They are also vulnerable to adversarial perturbations, such as misleading images influencing textual outputs, and their performance can degrade when relying on low-quality or outdated sources Chen et al.
+They are also vulnerable to adversarial perturbations, such as misleading images influencing textual outputs, and their performance degrades when relying on low-quality or outdated sources Chen et al.
 (2022b).
+MM-PoisonRAG Ha et al.
+(2025) and Poisoned-MRAG Liu et al.
+(2025b) demonstrate that even a few adversarial knowledge injections can hijack cross-modal retrieval and derail generation, underscoring the imperative for robust defense mechanisms against knowledge poisoning in multimodal RAG systems.
 While the trustworthiness of unimodal RAGs has been studied Zhou et al.
-(2024d), enhancing the robustness of multimodal RAGs remains an open challenge and a promising research direction.
-6.2 Reasoning, Alignment, and Retrieval Enhancement
-Multimodal RAGs struggle with compositional reasoning, where information from different modalities must be logically integrated to generate coherent and contextually rich outputs.
-While cross-modal techniques such as Multimodal-CoT Zhang et al.
-(2023b) have been proposed, further innovations are needed to improve the coherence and contextual relevance of multimodal outputs.
-Enhancing modality alignment and retrieval strategies, particularly for entity-aware retrieval, is essential.
-Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain largely underexplored in multimodal RAGs compared to text-based RAGs Zhang et al.
+(2024d), ensuring robustness in multimodal RAGs remains an open challenge and a crucial research direction.
+Reasoning, Alignment, and Retrieval Enhancement
+Multimodal RAGs struggle with compositional reasoning, requiring logical integration of information across modalities for coherent, context-rich outputs.
+While cross-modal techniques like Multimodal-CoT Zhang et al.
+(2023b) have emerged, further advancements are needed to enhance coherence and contextual relevance.
+Improving modality alignment and entity-aware retrieval is crucial.
+Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain underexplored in multimodal RAGs compared to text-based RAGs Zhang et al.
 (2024f); Procko and Ochoa (2024).
 Retrieval biases such as position sensitivity Hu et al.
-(2024c), redundant retrieval Nan et al.
-(2024), and biases propagated from training data or retrieved content Zhai (2024), pose significant challenges that require further attention.
-Another promising direction is developing a unified embedding space for all modalities, enabling direct multimodal search without intermediary conversion models (e.g., ASRs).
-Despite some progress, mapping multimodal knowledge into a unified space remains an open challenge with significant potential.
-6.3 Agent-Based and Self-Guided Systems
+(2024c), redundancy Nan et al.
+(2024b), and biases from training data or retrieved content Zhai (2024), pose significant challenges.
+A promising direction is a unified embedding space for all modalities, enabling direct multimodal search without intermediary models (e.g., ASRs).
+Despite progress, mapping multimodal knowledge into a unified space remains an open challenge with substantial potential.
+Agent-Based and Self-Guided Systems
 Recent trends indicate a shift towards agent-based multimodal RAGs that integrate retrieval, reasoning, and generation across diverse domains.
-Unlike static RAG systems, future multimodal RAGs should integrate interactive feedback and self-guided decision-making to iteratively refine outputs.
-Existing feedback mechanisms often fail to accurately determine whether errors stem from retrieval, generation, or other stages Dong et al.
+Unlike static RAGs, future systems should incorporate interactive feedback and self-guided decision-making to iteratively refine outputs.
+Existing feedback mechanisms often fail to determine whether errors stem from retrieval, generation, or other stages Dong et al.
 (2024b).
-The incorporation of reinforcement learning and end-to-end human-aligned feedback into multimodal RAGs remains largely unexplored but holds significant potential for enhancing these systems.
-These methods could enable multimodal RAGs to assess whether retrieval is necessary, evaluate the relevance of retrieved content, and dynamically determine the most suitable modalities for response generation.
-Achieving robust support for any-to-any modality is essential for adaptability in open-ended tasks Wu et al.
+The incorporation of reinforcement learning and end-to-end human-aligned feedback remains largely overlooked but holds significant potential for assessing whether retrieval is necessary, evaluating the relevance of retrieved content, and dynamically determining the most suitable modalities for response generation.
+Robust support for any-to-any modality is crucial for open-ended tasks Wu et al.
 (2024b).
-Future multimodal RAGs should incorporate data from diverse real-world sources, such as environmental sensors, alongside traditional modalities like text and images, to enhance situational awareness.
+Future multimodal RAGs should incorporate data from diverse real-world sources, such as environmental sensors, alongside traditional modalities to enhance situational awareness.
 This progression aligns with the trend toward embodied AI, where models integrate knowledge with physical interaction, enabling applications in robotics, navigation, and physics-informed reasoning.
 Bridging retrieval-based reasoning with real-world agency brings these systems closer to AGI.
-6.4 Long-Context Processing, Efficiency, Scalability, and Personalization
-High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing.
-Fixed extraction rates struggle to capture relevant frames, requiring adaptive selection based on content complexity and movement Kandhare and Gisselbrecht (2024).
-Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers emphasize the need for efficient, scalable architectures.
-Personalization mechanisms, such as adapting retrieval to user-specific contexts like medical history, remain underexplored.
-As personalization mechanisms evolve, ensuring privacy and mitigating the risks of sensitive data leakage in multimodal outputs remain critical challenges.
-Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples limits robust evaluation.
 5 Conclusion
-This study provides a comprehensive review of multimodal Retrieval-Augmented Generation (Multimodal RAG) literature.
-Specifically, we explore and categorize key advancements across different aspects of multimodal RAG systems, including retrieval, multimodal fusion, augmentation, generation, and training strategies.
-Additionally, we examine the tasks these systems address, their domain-specific applications, and the datasets, benchmarks, and evaluation methods.
-We also discuss open challenges and limitations in current approaches, along with promising future directions.
+This study provides a comprehensive review of multimodal RAG, categorizing key advancements in retrieval, multimodal fusion, augmentation, generation, training strategies, and agents.
+We also examine task-specific applications, datasets, benchmarks, and evaluation methods while highlighting open challenges and promising future directions.
 We hope this work inspires future research, particularly in enhancing cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
@@
 Additionally, we present a classification of multimodal RAG application domains in Figure 3.
-Figure 2 provides an overview of recent advances in multimodal retrieval-augmented generation (RAG) systems. The taxonomy is organized into several key categories.
-  • Retrieval strategies cover efficient search and similarity retrieval methods (including maximum inner product search (MIPS) variants and different multi-modal encoders) and modality-centric techniques that distinguish between text-, vision-, and video-centric as well as document retrieval models.
+Figure 2 provides an overview of recent advances in multimodal RAG systems. The taxonomy is organized into several key categories.
+  • Retrieval strategies cover efficient search and similarity retrieval methods (including maximum inner product search (MIPS) variants and different multimodal encoders) and modality-centric techniques that distinguish between text-, vision-, audio-, and video-centric as well as document retrieval models.
 Re-ranking strategies further refine these methods via optimized example selection, relevance scoring, and filtering.
-  • Fusion mechanisms are implemented through score fusion and alignment, attention-based techniques, and unified frameworks that project multimodal information into common representations.
+  • Fusion mechanisms cover score fusion and alignment techniques, including CLIP score fusion and prototype-based embeddings that unify multimodal representations, attention-based methods such as cross-attention and co-attention transformers that dynamically weight cross-modal interactions, and unified frameworks and projections like hierarchical fusion and dense-to-sparse projections that consolidate multimodal inputs.
   • Augmentation techniques address context enrichment as well as adaptive and iterative retrieval.
-  • Generation methods include in-context learning, reasoning, instruction tuning, and source attribution.
+  • Generation methods include in-context learning, reasoning, instruction tuning, source attribution, and agentic frameworks.
   • training strategies are characterized by approaches to alignment and robustness.
@@
 Appendix B Dataset and Benchmark
+Multimodal RAG research employs diverse datasets and benchmarks to evaluate retrieval, integration, and generation across heterogeneous sources.
+Image–text tasks, including captioning and retrieval, commonly use MS-COCO Lin et al.
+(2014), Flickr30K Young et al.
+(2014), and LAION-400M Schuhmann et al.
+(2021), while visual question answering (QA) with external knowledge is supported by OK-VQA Marino et al.
+(2019) and WebQA Chang et al.
+(2022).
+For complex multimodal reasoning, MultimodalQA Talmor et al.
+(2021) integrates text, images, and tables, whereas video-text tasks leverage ActivityNet Caba Heilbron et al.
+(2015) and YouCook2 Zhou et al.
+(2018).
+In the medical domain, MIMIC-CXR Johnson et al.
+(2019) and CheXpert Irvin et al.
+(2019) facilitate tasks such as medical report generation.
+It should be noted that a number of these datasets are unimodal (e.g., solely text-based or image-based).
+Unimodal datasets are frequently employed to represent a specific modality and are subsequently integrated with complementary datasets from other modalities.
+This modular approach allows each dataset to contribute its domain-specific strengths, thereby enhancing the overall performance of the multimodal retrieval and generation processes.
+Benchmarks assess multimodal RAG systems on visual reasoning, external knowledge integration, and dynamic retrieval.
+The $M^{2}RAG$ Ma et al.
+(2024d) benchmark provides a unified evaluation framework that combines fine-grained text-modal and multimodal metrics to jointly assess both the quality of generated language and the effective integration of visual elements.
+In addition, Liu et al.
+(2025d) introduce another specialized benchmark for multimodal RAG that evaluates performance across image captioning, multi-modal question answering, fact verification, and image reranking in an open-domain retrieval setting.
+Vision-focused evaluations, including MRAG-Bench Hu et al.
+(2024c), VQAv2 Goyal et al.
+(2017a) and VisDoMBench Suri et al.
+(2024), test models on complex visual tasks.
+Dyn-VQA Li et al.
+(2024d), MMBench Liu et al.
+(2025c), and ScienceQA Lu et al.
+(2022) evaluate dynamic retrieval and multi-hop reasoning across textual, visual, and diagrammatic inputs.
+Knowledge-intensive benchmarks, such as TriviaQA Joshi et al.
+(2017) and Natural Questions Kwiatkowski et al.
+(2019), together with document-oriented evaluations such as OmniDocBench Ouyang et al.
+(2024), measure integration of unstructured and structured data.
+Advanced retrieval benchmarks such as RAG-Check Mortaheb et al.
+(2025a) evaluate retrieval relevance and system reliability, while specialized assessments like Counterfactual VQA Niu et al.
+(2021) test robustness against adversarial inputs.
+Additionally, OCR impact studies such as OHRBench Zhang et al.
+(2024d) examine the cascading effects of errors on RAG systems.
+The choice of dataset significantly influences the evaluation focus, ranging from foundational pre-training on large-scale image-text corpora like LAION-5B Schuhmann et al.
+(2022) (5.85 billion pairs) or MINT-1T Awadalla et al.
+(2024) (3.4 billion images with 1 trillion text tokens), to more specialized tasks such as video understanding with HowTo100M Miech et al.
+(2019) (136 million video clips) or medical report generation using MIMIC-CXR Johnson et al.
+(2019) (125,417 image-report pairs).
+Datasets are often tailored for specific downstream tasks.
+For visual question answering, VQA Antol et al.
+(2015) and A-OKVQA Schwenk et al.
+(2022) specifically require external knowledge, making them suitable for evaluating RAG systems’ ability to retrieve and reason over such knowledge.
+For document understanding, datasets such as DocVQA Mathew et al.
+(2021) and M3DocVQA Cho et al.
+(2024) are essential.
+As discussed in the benchmarks overview above, unified evaluation frameworks like $M^{2}RAG$ Ma et al.
+(2024d) provide a comprehensive assessment across multiple tasks, including image captioning, visual question answering, and fact verification.
+Evaluating complex reasoning capabilities in multimodal RAG systems has become increasingly important.
+Datasets such as MultimodalQA Talmor et al.
+(2021), WebQA Chang et al.
+(2022), and ScienceQA Lu et al.
+(2022) are specifically designed to benchmark multi-hop reasoning abilities crucial for advanced RAG systems, with Dyn-VQA Li et al.
+(2024d) additionally focusing on robustness to changing information.
+Comparative Analysis of Datasets
+Understanding the strategic trade-offs in dataset design is crucial for multimodal RAG development, as different dataset characteristics serve distinct purposes across the model development pipeline.
+(i) Scale and Diversity vs.
+Curation: Large-scale datasets such as LAION-5B Schuhmann et al.
+(2022) and Conceptual Captions Sharma et al.
+(2018) provide substantial scale essential for pre-training, enabling models to learn generalizable representations across diverse domains.
+However, their reliance on web-crawled data introduces inherent noise that can compromise training quality.
+Conversely, smaller, meticulously curated datasets like Flickr30K Young et al.
+(2014) (31,000 images with human annotations) and domain-specific collections such as Fashionpedia Jia et al.
+(2020) (48,000 images with segmentation masks) prioritize annotation quality over scale, making them essential for fine-tuning models and assessing specialized performance.
+(ii) Modality Focus and Combination: While many systems aggregate unimodal datasets to construct multimodal contexts, datasets explicitly designed for multimodal tasks demonstrate superior alignment between modalities.
+Foundational datasets like MS-COCO Lin et al.
+(2014) and VQA Antol et al.
+(2015) establish benchmarks for image-text understanding, while specialized collections such as AudioSet Gemmeke et al.
+(2017) (2 million audio clips) and AudioCaps Kim et al.
+(2019) (46,000 audio clips with captions) address audio-language integration.
+Emerging modalities like 3D (e.g., ShapeNet Chang et al.
+(2015)) remain underrepresented, yet are essential for expanding RAG applications into spatial reasoning domains.
 Table 1 and  Table 2 present a comprehensive overview of datasets and benchmarks commonly employed in multimodal RAG research. The table is organized into five columns:
@@
   • Link: A hyperlink is provided to direct readers to the official repository or additional resources for the dataset or benchmark, thereby facilitating further exploration of its properties and applications.
+Limitations of Existing Datasets and Benchmarks
+While the datasets and benchmarks discussed above have significantly advanced multimodal RAG research, several limitations persist that offer important avenues for future work:
+(i) Bias and Fairness: Large datasets, especially those scraped from the web, can inherit societal biases related to gender, race, or culture.
+This can lead to skewed model behavior and unfair outcomes.
+Efforts to create more balanced datasets are crucial, but comprehensive bias auditing across modalities remains a challenge.
+(ii) Annotation Quality and Noise: The trade-off between dataset scale and annotation quality remains a persistent challenge.
+While large datasets facilitate broad learning, their often noisy or weakly supervised labels (e.g., alt-text for images) can hinder precise model training.
+As demonstrated by OHRBench Zhang et al.
+(2024d), OCR errors exemplify how noise in one modality can cascade and affect overall RAG system performance.
+(iii) Coverage and Generalization Gaps: Many datasets are domain-specific, which can limit the generalization of models to out-of-domain scenarios.
+There is a need for more datasets covering a wider array of real-world contexts and less common modalities.
+(iv) Real-World Complexity and Long-Context Understanding: Current datasets inadequately capture real-world multimodal information complexity.
+Challenges include efficient sampling of relevant video frames, handling multi-page documents with numerous images, and processing dynamic information environments; benchmarks like Dyn-VQA Li et al.
+(2024d) are, however, beginning to address this latter challenge.
+(v) Lack of Adversarial and Robustness Testing: While benchmarks like Counterfactual VQA Niu et al.
+(2021) specifically test robustness against certain perturbations, there is a general scarcity of datasets containing multimodal adversarial examples or structured negative instances.
+Such datasets are vital for developing more robust and reliable RAG systems that can handle out-of-distribution inputs or misleading information.
+(vi) Retrieval-Generation Integration: Many benchmarks evaluate retrieval and generation components separately rather than assessing their synergistic interplay.
+More holistic evaluation frameworks are needed that jointly measure retrieval accuracy, relevance of retrieved multimodal context, and final output quality, as aimed by benchmarks like MRAG-Bench Hu et al.
+(2024c) for visual integration and RAG-Check Mortaheb et al.
+(2025a) for retrieval relevance.
+(vii) Limited Support for "Any-to-Any" Modalities: While current research primarily focuses on text, image, video, and audio, future RAG systems are envisioned to support any-to-any modality interactions.
+Existing datasets offer limited support for such comprehensive multimodality.
 Table 1: Overview of Popular Datasets in Multimodal RAG Research.
@@
 Name Statistics and Description Modalities Link
-Image-Text General LAION-400M Schuhmann et al. (2021) 200M image–text pairs; used for pre-training multimodal models. Image, Text LAION-400M
-Conceptual-Captions (CC) Sharma et al. (2018) 15M image–caption pairs; multilingual English–German image descriptions. Image, Text Conceptual Captions
+Image-Text General LAION-400M Schuhmann et al. (2021) 400M image–text pairs; used for pre-training multimodal models. Image, Text LAION-400M
+Conceptual-Captions (CC) Sharma et al. (2018) More than 3M image–caption pairs; multilingual English–German image descriptions. Image, Text Conceptual Captions
 CIRR Liu et al. (2021) 36,554 triplets from 21,552 images; focuses on natural image relationships. Image, Text CIRR
@@
 NoCaps Agrawal et al. (2019) For zero–shot image captioning evaluation; 15K images. Image, Text NoCaps
-Laion-5B Schuhmann et al. (2022) 5B image–text pairs used as external memory for retrieval. Image, Text LAION-5B
-COCO-CN Author and Author (2018) 20,341 images for cross-lingual tagging and captioning with Chinese sentences. Image, Text COCO-CN
+Laion-5B Schuhmann et al. (2022) 5.85B image–text pairs used as external memory for retrieval. Image, Text LAION-5B
+COCO-CN Li et al. (2019) 20,341 images for cross-lingual tagging and captioning with Chinese sentences. Image, Text COCO-CN
 CIRCO Baldrati et al. (2023) 1,020 queries with an average of 4.53 ground truths per query; for composed image retrieval. Image, Text CIRCO
-Video-Text BDD-X Xu et al. (2018) 77 hours of driving videos with expert textual explanations; for explainable driving behavior. Video, Text BDD-X
+MINT-1T Awadalla et al. (2024) 1T text tokens and 3.4B images; 10x larger than existing open-source datasets. Image, Text MINT-1T
+ShareGPT4V Chen et al. (2024c) 1.2M images with GPT-4-generated captions, including spatial and factual details. Image, Text ShareGPT4V
+OmniCorpus Li et al. (2025b) 8.6B images and 1.7T tokens across 2.2B web documents; interleaved text-image layout. Image, Text OmniCorpus
+Video-Text BDD-X Kim et al. (2018) 77 hours of driving videos with expert textual explanations; for explainable driving behavior. Video, Text BDD-X
 YouCook2 Zhou et al. (2018) 2,000 cooking videos with aligned descriptions; focused on video–text tasks. Video, Text YouCook2
@@
 SoccerNet Giancola et al. (2018) Videos and metadata for 550 soccer games; includes transcribed commentary and key event annotations. Video, Text SoccerNet
-MSR-VTT Xu et al. (2016) 10,000 videos with 20 captions each; a large video description dataset. Video, Text MSR-VTT
 MSVD Chen and Dolan (2011) 1,970 videos with approximately 40 captions per video. Video, Text MSVD
@@
 DiDemo Anne Hendricks et al. (2017) 10,000 videos with four concatenated captions per video; with temporal localization of events. Video, Text DiDemo
-Breakfast Kuehne et al. (2014) 1,712 videos of breakfast preparation; one of the largest fully annotated video datasets. Video, Text Breakfast
 COIN Tang et al. (2019) 11,827 instructional YouTube videos across 180 tasks; for comprehensive instructional video analysis. Video, Text COIN
 MSRVTT-QA Xu et al. (2017) Video question answering benchmark. Video, Text MSRVTT-QA
-MSVD-QA Xu et al. (2017) 1,970 video clips with approximately 50.5K QA pairs; video QA dataset. Video, Text MSVD-QA
 ActivityNet-QA Yu et al. (2019) 58,000 human–annotated QA pairs on 5,800 videos; benchmark for video QA models. Video, Text ActivityNet-QA
-EpicKitchens-100 Dima (2020) 700 videos (100 hours of cooking activities) for online action prediction; egocentric vision dataset. Video, Text EPIC-KITCHENS-100
+EpicKitchens-100 Damen et al. (2022) 700 videos (100 hours of cooking activities) for online action prediction; egocentric vision dataset. Video, Text EPIC-KITCHENS-100
 Ego4D Grauman et al. (2022) 4.3M video–text pairs for egocentric videos; massive–scale egocentric video dataset. Video, Text Ego4D
@@
 VATEX Wang et al. (2019) 34,991 videos, each with multiple captions; a multilingual video–and–language dataset. Video, Text VATEX
-Charades Sigurdsson et al. (2016) 9,848 video clips with textual descriptions; a multimodal research dataset. Video, Text Charades
 WebVid Bain et al. (2021) 10M video–text pairs (refined to WebVid-Refined-1M). Video, Text WebVid
+InternVid Wang et al. (2023b) 7M YouTube videos (760K hours), 234M clips, 4.1B words; used for video-text pretraining and representation learning. Video, Text InternVid
+OpenVid-1M Nan et al. (2024a) 1 million video-text pairs for multimodal learning. Video, Text OpenVid-1M
 Youku-mPLUG Xu et al. (2023) Chinese dataset with 10M video–text pairs (refined to Youku-Refined-1M). Video, Text Youku-mPLUG
@@
 SpeechBrown Abootorabi and Asgari (2024) 55K paired speech-text samples; 15 categories covering diverse topics from religion to fiction. Audio, Text SpeechBrown
-AudioCap Kim et al. (2019) 46K audio clips paired with human-written text captions. Audio, Text AudioCaps
+AudioCaps Kim et al. (2019) 46K audio clips paired with human-written text captions. Audio, Text AudioCaps
+MusicCaps Agostinelli et al. (2023) It is composed of 5.5k music-text pairs, with rich text descriptions provided by human experts. Audio, Text MusicCaps
+Clotho Drossos et al. (2020) Audio captioning dataset with diverse soundscapes. Audio, Text Clotho
+WavCaps Mei et al. (2024) Large-scale weakly-labeled audio-text dataset, comprising approximately 400k audio clips with paired captions. Audio, Text WavCaps
+Spoken SQuAD Li et al. (2018) Audio version of the SQuAD dataset for spoken question answering, focusing on the listening comprehension task. Audio, Text Spoken SQuAD
 AudioSet Gemmeke et al. (2017) 2,084,320 human–labeled 10–second sound clips from YouTube; 632 audio event classes. Audio, Text AudioSet
@@
 PubLayNet Zhong et al. (2019) 100,000 training samples and 2,160 test samples built from PubLayNet (tailored for the medical domain). Image, Text PubLayNet
+Quilt-1M Ikezogwo et al. (2023) 438K medical images with 768K text pairs; includes microscopic images and UMLS entities. Image, Text Quilt-1M
 Fashion Fashion-IQ Wu et al. (2021) 77,684 images across three categories; evaluated with Recall@10 and Recall@50. Image, Text Fashion IQ
-FashionGen Hadi Kiapour et al. (2018) 260.5K image–text pairs of fashion images and item descriptions. Image, Text Fashion-Gen
+FashionGen Rostamzadeh et al. (2018) 260.5K image–text pairs of fashion images and item descriptions. Image, Text Fashion-Gen
 VITON-HD Choi et al. (2021) 83K images for virtual try–on; high–resolution clothing items. Image, Text VITON-HD
-Fashionpedia Author and Author (2023a) 48,000 fashion images annotated with segmentation masks and fine-grained attributes. Image, Text Fashionpedia
+Fashionpedia Jia et al. (2020) 48,000 fashion images annotated with segmentation masks and fine-grained attributes. Image, Text Fashionpedia
 DeepFashion Liu et al. (2016) Approximately 800K diverse fashion images for pseudo triplet generation. Image, Text DeepFashion
-3D ShapeNet Chang et al. (2015) 7,500 text–3D data pairs; repository for 3D CAD models. Text, 3D ShapeNet
+3D ShapeNet Chang et al. (2015) Covering 55 common object categories with  51,300 unique 3D models. Text, 3D ShapeNet
 Knowledge & QA VQA Antol et al. (2015) 400K QA pairs with images for visual question answering. Image, Text VQA
@@
 ELI5 Fan et al. (2019) 270K complex and diverse questions augmented with web pages and images. Text ELI5
-ViQuAE Biten et al. (2022) 11.8M passages from Wikipedia covering 2,397 unique entities; knowledge–intensive QA. Text ViQuAE
+MultimodalQA Talmor et al. (2021) 29,918 questions requiring multi-modal multi-hop reasoning over text, tables, and images. Image, Text, Table MultimodalQA
+ViQuAE Lerner et al. (2022) 11.8M passages from Wikipedia covering 2,397 unique entities; knowledge–intensive QA. Text ViQuAE
 OK-VQA Marino et al. (2019) 14K questions requiring external knowledge for VQA. Image, Text OK-VQA
-WebQA Li et al. (2022b) 46K queries that require reasoning across text and images. Text, Image WebQA
-Infoseek Li et al. (2021) Fine-grained visual knowledge retrieval using a Wikipedia–based knowledge base ( 6M passages). Image, Text Infoseek
-ClueWeb22 Callan et al. (2022) 10 billion web pages organized into three subsets; a large–scale web corpus. Text ClueWeb22
+WebQA Chang et al. (2022) 46K queries that require reasoning across text and images. Text, Image WebQA
+Infoseek Chen et al. (2023) Fine-grained visual knowledge retrieval using a Wikipedia–based knowledge base ( 6M passages). Image, Text Infoseek
+ClueWeb22 Overwijk et al. (2022) 10 billion web pages organized into three subsets; a large–scale web corpus. Text ClueWeb22
 MOCHEG Yao et al. (2023) 15,601 claims annotated with truthfulness labels and accompanied by textual and image evidence. Text, Image MOCHEG
@@
 XL-HeadTags Shohan et al. (2024) 415K news headline-article pairs consist of 20 languages across six diverse language families. Text XL-HeadTags
+DocVQA Mathew et al. (2021) 12,767 diverse document images with 50K QA pairs, categorized by reasoning type to evaluate DocVQA methods. Image, Text DocVQA
+ChartQA Masry et al. (2022) 9.6K human-written QA pairs + 23.1K generated from chart summaries. Image, Text ChartQA
+DVQA Kafle et al. (2018) 3.5M QA pairs on 300K diagrams, evaluating structure, data retrieval, and reasoning. Image, Text DVQA
+RETVQA Penamakuri et al. (2023) 416,000 QA samples where retrieval from a large image set is needed to answer questions; emphasizes RAG pipeline. Image, Text RETVQA
 SEED-Bench Li et al. (2023a) 19K multiple–choice questions with accurate human annotations across 12 evaluation dimensions. Text SEED-Bench
-Other ImageNet Deng et al. (2009) 14,197,122 images for perspective understanding; a hierarchical image database. Image ImageNet
-Oxford Flowers102 Nilsback and Zisserman (2008) 102 flower categories with five examples per category; image classification dataset. Image Oxford Flowers102
-Stanford Cars Krause et al. (2013) Images of different car models (five examples per model); for fine-grained categorization. Image Stanford Cars
-GeoDE Author and Author (2023b) 61,940 images from 40 classes across 6 world regions; emphasizes geographic diversity in object recognition. Image GeoDE
+M3DocVQA Cho et al. (2024) 2,441 multi-hop questions across 3,368 PDF documents; evaluates open-domain DocVQA. Image, Text M3DocVQA
+MMLongBench-Doc Ma et al. (2024c) 135 lengthy PDFs with 1,091 questions; focuses on multi-hop reasoning in single documents. Image, Text MMLongBench-Doc
+Other GeoDE Ramaswamy et al. (2023) 61,940 images from 40 classes across 6 world regions; emphasizes geographic diversity in object recognition. Image GeoDE
+RU-AI Huang et al. (2025) 1.47M samples of real vs AI-generated content for fake detection robustness. Image, Text, Audio RU-AI
+MIMIC-IT Li et al. (2025a) 2.8M multimodal instruction-response pairs for model alignment. Image, Video, Text MIMIC-IT
+MMVQA Ding et al. (2024c) 262K question-answer pairs across 3,146 multipage research PDFs for robust multimodal information retrieval. Image, Text MMVQA
 Table 2: Overview of Popular Benchmarks in Multimodal RAG Research.
@@
 Cross-Modal Understanding MRAG-Bench Hu et al. (2024c) Evaluates visual retrieval, integration, and robustness to irrelevant visual information. Images MRAG-Bench
-M2RAG Ma et al. (2024c) Benchmarks multimodal RAG; evaluates retrieval, multi-hop reasoning, and integration. Images + Text M2RAG
-Dyn-VQA Li et al. (2024b) Focuses on dynamic retrieval, multi-hop reasoning, and robustness to changing information. Images + Text Dyn-VQA
-MMBench Liu et al. (2025) Covers VQA, captioning, retrieval; evaluates cross-modal understanding across vision, text, and audio. Images + Text + Audio MMBench
-ScienceQA Saikh et al. (2022) Contains 21,208 questions; tests scientific reasoning with text, diagrams, and images. Images + Diagrams + Text ScienceQA
+$M^{2}RAG$ Ma et al. (2024d) Benchmarks multimodal RAG; evaluates retrieval, multi-hop reasoning, and integration. Images + Text $M^{2}RAG$
+Dyn-VQA Li et al. (2024d) Focuses on dynamic retrieval, multi-hop reasoning, and robustness to changing information. Images + Text Dyn-VQA
+MMBench Liu et al. (2025c) Covers VQA, captioning, retrieval; evaluates cross-modal understanding across vision, text, and audio. Images + Text + Audio MMBench
+ScienceQA Lu et al. (2022) Contains 21,208 questions; tests scientific reasoning with text, diagrams, and images. Images + Diagrams + Text ScienceQA
 SK-VQA Su et al. (2024b) Offers 2 million question-answer pairs; focuses on synthetic knowledge, multimodal reasoning, and external knowledge integration. Images + Text SK-VQA
@@
 Natural Questions Kwiatkowski et al. (2019) Contains 307,373 training examples; real-world search queries, adaptable with visual contexts. Text Natural Questions
-Appendix C Metrics
-Accuracy: Accuracy is typically defined as the ratio of correctly predicted instances to the total instances. In retrieval-based tasks, Top-K Accuracy is defined as:
+{forest}
+forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2em, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0.4em, inner ysep=0.6em, line width=0.8pt, text width=8.5em, ver/.style= rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=17em, font=, minimum height=3em, minimum width=5em, , leaf/.style= fill=green!20, text= black, text width=44.5em, font=, inner xsep=0.4em, inner ysep=0.6em, draw, , intermediate/.style= font=, minimum height=2.5em, minimum width=18em, text width=18em, draw, , [ Multimodal RAG Application Domains, ver , fill=white!20 [ Healthcare and Medicine (§E) , fill=yellow!20, intermediate [ MMED-RAG Xia et al.
+(2024a), RULE Xia et al.
+(2024b), AsthmaBot Bahaj and Ghogho (2024), Realm Zhu et al.
+(2024c),  Su et al.
+(2024a), FactMM-RAG Sun et al.
+(2024b), RA-RRG Choi et al.
+(2025) , leaf ] ] [ Software Engineering (§E) , fill=yellow!20, intermediate [ DocPrompting Zhou et al.
+(2023), RACE Shi et al.
+(2022), CEDAR Nashid et al.
+(2023), REDCODER Parvez et al.
+(2021) , leaf ] ] [ Fashion and E-Commerce (§E) , fill=yellow!20, intermediate [ Unifashion Zhao et al.
+(2024),  Dang (2024), Fashion-RAG Sanguigni et al.
+(2025), LLM4DESIGN Chen et al.
+(2024d) , leaf ] ] [ Entertainment and Social Computing (§E) , fill=yellow!20, intermediate [ SoccerRAG Strand et al.
+(2024), MMRA Zhong et al.
+(2024) , leaf ] ] [ Emerging Applications (§E) , fill=yellow!20, intermediate [ RAG-Driver Yuan et al.
+(2024), ENWAR Nazar et al.
+(2024),  Riedler and Langer (2024), Img2Loc Zhou et al.
+(2024e) , leaf ] ] ]
+Figure 3: Taxonomy of application domains for Multimodal Retrieval-Augmented Generation systems.
+Appendix C Evaluation and Metrics
+Evaluating multimodal RAG models is complex due to their varied input types and complex structure.
+The evaluation combines metrics from VLMs, generative AI, and retrieval systems to assess capabilities like text/image generation and information retrieval.
+Our review found about 60 different metrics used in the field.
+In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
+Retrieval Evaluation
+Retrieval performance is measured through accuracy, recall, and precision metrics, with an F1 score combining recall and precision.
+Accuracy is typically defined as the ratio of correctly predicted instances to the total instances.
+In retrieval-based tasks, Top-K Accuracy is defined as:
 […]
-FID (Fréchet inception distance): FID is a metric used to assess the quality of images created by a generative model. The formula for FID is:
-FID $=\|\mu_{r}-\mu_{g}\|^{2}+tr(\Sigma_{r}+\Sigma_{g}-2\sqrt{\Sigma_{r}\Sigma_{g}})$ (4)
-where $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of real images’ feature representations, respectively.
-$\mu_{g}$ and $\Sigma_{g}$ are the mean and covariance of generated images’ feature representations, respectively.
-To extract features, InceptionV3 is typically used.
-ROUGE-N (N-gram Recall): The ROUGE metric is commonly used to evaluate text summarization and generation.
+Recall@K, which examines relevant items in top K results, is preferred over standard recall.
+Mean Reciprocal Rank (MRR) serves as another key metric for evaluation, which is utilized by Adjali et al.
+(2024); Nguyen et al.
+(2024).
+MRR measures the rank position of the first relevant result in the returned list.
+The formula for calculating MRR is:
+$\text{MRR}=\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{\text{rank}_{q}}$ (2)
+where $Q$ is the total number of queries. $rank_{q}$ is the rank of the first relevant result for query $q$.
+Modality Evaluation
+Modality-based evaluations primarily focus on text and image, assessing their alignment, text fluency, and image caption quality.
+For text evaluation, metrics include Exact Match (EM), BLEU Papineni et al.
+(2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005).
+The ROUGE metric is commonly used to evaluate text summarization and generation.
 ROUGE-N measures the overlap of N-grams between the generated and reference text.
@@
 ROUGE-L $=\frac{LCS(X,Y)}{|Y|}$ (4)
-BLEU: BLEU is another metric used for assessing text generation. The formula for calculating BLEU is:
+BLEU is another metric used for assessing text generation. The formula for calculating BLEU is:
 $\text{BLEU}(p_{n},\text{BP})=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$ (5)
@@
 Here, $rl$ represents the reference length and $cl$ represents the candidate length.
-BERTScore: BERTScore is a metric for evaluating the quality of text generation, based on the similarity between the contextual embeddings of words in the candidate and reference texts.
-The formula for calculating BERTScore is:
-$\text{BERTScore}(c,r)=\frac{1}{|c|}\sum_{i=1}^{|c|}\max_{j=1}^{|r|}\cos(\mathbf{e}_{i},\mathbf{e}_{j})$ (9)
-$c$ is the candidate sentence, and $r$ is the reference sentence.
-$e_{i}$ and $e_{j}$ are the embeddings (e.g., from BERT) for words $c_{i}$ and $r_{j}$ in the candidate and reference sentences, respectively.
-CLIPScore: CLIPScore is a metric that evaluates the similarity between the text and an image by using the CLIP model. The formula for calculating CLIPScore is:
-$\text{CLIPScore}=\frac{\cos(\mathbf{t},\mathbf{i})}{\|\mathbf{t}\|\cdot\|\mathbf{i}\|}$ (10)
-where t and i are text and image embedding respectively.
-Mean Reciprocal Rank (MRR): MRR is a metric used to evaluate the performance of systems that return a list of results, such as search engines or recommendation systems.
-MRR measures the rank position of the first relevant result in the returned list.
-The formula for calculating MRR is:
-$\text{MRR}=\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{\text{rank}_{q}}$ (11)
-where $Q$ is the total number of queries. $rank_{q}$ is the rank of the first relevant result for query $q$.
-Appendix D Loss Function
+MultiRAGen Shohan et al. (2024) uses Multilingual ROUGE for multilingual settings.
+For image captioning, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al.
+(2015) measures caption quality using TF-IDF and cosine similarity Yasunaga et al.
+(2023); Zhao et al.
+(2024); Luo et al.
+(2024a); Yuan et al.
+(2024); Sharifymoghaddam et al.
+(2024); Hu et al.
+(2023); Rao et al.
+(2024); Xu et al.
+(2024a); Kim et al.
+(2024); Zhang et al.
+(2024c), while SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al.
+(2016) focuses on semantics.
+SPIDEr Liu et al.
+(2017), used in Zhang et al.
+(2024c), combines both metrics.
+For semantic alignment, BERTScore Zhang et al.
+(2020) compares BERT embeddings Sun et al.
+(2024b); Shohan et al.
+(2024), and evaluates fluency Chen et al.
+(2022a); Zhi Lim et al.
+(2024); Ma et al.
+(2024d).
+CLIP Score Hessel et al.
+(2021), used in Sharifymoghaddam et al.
+(2024); Zhang et al.
+(2024c), measures image-text similarity using CLIP Radford et al.
+(2021).
+The formula for calculating CLIPScore is:
+$\text{CLIPScore}=\frac{\mathbf{t}.\mathbf{i}}{\|\mathbf{t}\|\|\mathbf{i}\|}$ (7)
+where t and i are text and image embedding, respectively.
+For image quality, FID (Fréchet Inception Distance) Heusel et al.
+(2017) compares feature distributions Yasunaga et al.
+(2023); Zhao et al.
+(2024); Sharifymoghaddam et al.
+(2024); Zhang et al.
+(2024c), while KID (Kernel Inception Distance) Bińkowski et al.
+(2018) provides an unbiased alternative.
+The formula for FID is:
+FID $=\|\mu_{r}-\mu_{g}\|^{2}+tr(\Sigma_{r}+\Sigma_{g}-2\sqrt{\Sigma_{r}\Sigma_{g}})$ (8)
+where $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of real images’ feature representations, respectively.
+$\mu_{g}$ and $\Sigma_{g}$ are the mean and covariance of generated images’ feature representations, respectively.
+To extract features, InceptionV3 Szegedy et al.
+(2016) is typically used.
+Inception Score (IS) evaluates image diversity and quality through classification probabilities Zhi Lim et al.
+(2024).
+For audio evaluation, Zhang et al.
+(2024c) use human annotators to assess sound quality (OVL) and text relevance (REL), while also employing Fréchet Audio Distance (FAD) Kilgour et al.
+(2019), an audio-specific variant of FID.
+System efficiency is measured through FLOPs, execution time, response time, and retrieval time per query Nguyen et al.
+(2024); Strand et al.
+(2024); Dang (2024); Zhou (2024).
+Domain-specific metrics include geodesic distance for geographical accuracy Zhou et al.
+(2024e), and Clinical Relevance for medical applications Lahiri and Hu (2024).
+Appendix D Robustness Advancements and Loss Functions
+D.1 Robustness and Noise Management
+Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024).
+Managing noisy retrieval inputs is critical for maintaining model performance.
+MORE Cui et al.
+(2024) injects irrelevant results during training to enhance focus on relevant inputs.
+AlzheimerRAG Lahiri and Hu (2024) uses progressive knowledge distillation to reduce noise while maintaining multimodal alignment.
+RAGTrans Cheng et al.
+(2024) leverages hypergraph-based knowledge aggregation to refine multimodal representations, ensuring more effective propagation of relevant information.
+RA-BLIP Ding et al.
+(2024b) introduces the Adaptive Selection Knowledge Generation (ASKG) strategy, which leverages the implicit capabilities of LLMs to filter relevant knowledge for generation through a denoising-enhanced loss term, eliminating the need for fine-tuning.
+This approach achieves strong performance compared to baselines while significantly reducing computational overhead by minimizing trainable parameters.
+RagVL Chen et al.
+(2024e) improves robustness through noise-injected training by adding hard negative samples at the data level and applying Gaussian noise with loss reweighting at the token level, enhancing the model’s resilience to multimodal noise.
+Finally, RA-CM3 Yasunaga et al.
+(2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.
+D.2 Loss Function
 InfoNCE (Information Noise Contrastive Estimation): The InfoNCE loss is commonly used in self-supervised learning, especially in contrastive learning methods. The formula for InfoNCE loss is:
@@
 $\scriptstyle\mathcal{L}=\sum_{i=1}^{N}\max(0,\|f(x_{a}^{i})-f(x_{p}^{i})\|^{2}-\|f(x_{a}^{i})-f(x_{n}^{i})\|^{2}+\alpha)$ (12)
-where $x_{a}^{i}$ is the anchor sample. $x_{p}^{i}$ and $x_{n}^{i}$ are the positive and negative samples respectively. $f(x)$ is the neural network.
+where $x_{a}^{i}$ is the anchor sample. $x_{p}^{i}$ and $x_{n}^{i}$ are the positive and negative samples, respectively. $f(x)$ is the neural network.
+Appendix E Applications and Relevant Tasks
+Multimodal RAG extends traditional RAG beyond unimodal settings to cross-modal tasks.
+In content generation, it enhances image captioning Zhi Lim et al.
+(2024); Hu et al.
+(2023); Rao et al.
+(2024) and text-to-image synthesis Yasunaga et al.
+(2023); Chen et al.
+(2022b) by retrieving relevant contextual information.
+It also improves coherence in visual storytelling and factual alignment in multimodal summarization Tonmoy et al.
+(2024).
+In knowledge-intensive applications, multimodal RAG supports open-domain QA Chen et al.
+(2024e); Ding et al.
+(2024b); Yuan et al.
+(2023), video-based QA Luo et al.
+(2024b), fact verification Khaliq et al.
+(2024), and zero-shot image–text retrieval Yang et al.
+(2024b), grounding responses in retrieved knowledge and thereby mitigating hallucinations.
+Additionally, the incorporation of chain-of-thought reasoning Zhai (2024); Khaliq et al.
+(2024) further enhances complex problem-solving and inference.
+Finally, their integration into AI assistants such as Gemini Team et al.
+(2024) enables natural language-driven visual search, document understanding, and multimodal reasoning.
+Multimodal RAGs are increasingly applied across diverse domains, including healthcare, software engineering, and creative industries (e.g., fashion and design automation).
+The taxonomy of application domains can be seen in  Figure 3.
+The following sections explore domain-specific adaptations of these techniques in greater depth.
+Healthcare and Medicine
+Multimodal RAG enhances clinical decision-making through integrated analysis of medical imaging, electronic health records, and biomedical literature.
+Systems like MMED-RAG Xia et al.
+(2024a) address diagnostic uncertainty in medical visual question answering by aligning radiology images with contextual patient data.
+In automated report generation, RULE Xia et al.
+(2024b) mitigates hallucinations through dynamic retrieval of clinically similar cases.
+Similarly, RA-RRG Choi et al.
+(2025) first leverages an LLM to extract key textual phrases from a report corpus, then employs a multimodal retriever to link the visual features to these relevant phrases.
+The coherent report is generated after being retrieved by another LLM without fine-tuning, thereby reducing hallucinations.
+FactMM-RAG Sun et al.
+(2024b) further automates radiology report drafting by retrieving biomarker correlations from medical ontologies, exemplifying the paradigm’s capacity to operationalize expert knowledge at scale.
+AsthmaBot Bahaj and Ghogho (2024) introduces a multimodal RAG-based approach for supporting asthma patients across multiple languages, enabling structured, language-specific semantic searches.
+Predictive frameworks such as Realm Zhu et al.
+(2024c) demonstrate robust risk assessment by fusing heterogeneous patient data streams, while Su et al.
+(2024a) advance privacy-preserving architectures for federated clinical data integration.
+Software Engineering
+Code generation systems leverage multimodal RAG to synthesize context-aware solutions from technical documentation and version histories.
+DocPrompting Zhou et al.
+(2023) improves semantic coherence in code completion by retrieving API specifications and debugging patterns.
+Commit message generation models like RACE Shi et al.
+(2022) contextualize code diffs against historical repository activity, while CEDAR Nashid et al.
+(2023) optimizes few-shot learning through retrieval-based prompt engineering.
+REDCODER Parvez et al.
+(2021) enhances code summarization via semantic search across open-source repositories, preserving syntactic conventions across programming paradigms.
+Fashion and E-Commerce
+Cross-modal alignment drives advancements in product discovery and design automation.
+UniFashion Zhao et al.
+(2024) enables style-aware retrieval by jointly embedding garment images and textual descriptors, while Dang (2024) reduces search friction through multimodal query expansion.
+For fashion image editing, Fashion-RAG Sanguigni et al.
+(2025) employs a retrieval-augmented approach, retrieving garments that match textual descriptions and integrating their attributes into image generation via textual inversion techniques within diffusion models, ensuring personalized and contextually relevant outputs.
+LLM4DESIGN Chen et al.
+(2024d) demonstrates architectural design automation by retrieving compliance constraints and environmental impact assessments, underscoring RAG’s adaptability to creative domains.
+Entertainment and Social Computing
+Multimedia analytics benefit from RAG’s capacity to correlate heterogeneous signals.
+SoccerRAG Strand et al.
+(2024) derives tactical insights by linking match footage with player statistics.
+MMRA Zhong et al.
+(2024) predicts content virality through joint modeling of visual aesthetics and linguistic engagement patterns.
+Emerging Applications
+Autonomous systems adopt multimodal RAG for explainable decision-making, as seen in RAG-Driver’s Yuan et al.
+(2024) real-time retrieval of traffic scenarios during navigation.
+ENWAR Nazar et al.
+(2024) enhances wireless network resilience through multi-sensor fusion, while Riedler and Langer (2024) streamline equipment maintenance by retrieving schematics during fault diagnosis.
+Geospatial systems such as Img2Loc Zhou et al.
+(2024e) advance image geolocalization through cross-modal landmark correlation.
+Appendix F Additional Future Directions
+High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing.
+Fixed extraction rates struggle to capture relevant frames, requiring adaptive selection based on content complexity and movement Kandhare and Gisselbrecht (2024).
+Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers emphasize the need for efficient, scalable architectures.
+Personalization mechanisms, like user-specific retrieval (e.g., adapting to medical history), remain underexplored.
+As these systems evolve, ensuring privacy and preventing sensitive data leakage in multimodal outputs is critical.
+Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples limits robust evaluation.